### v5

# onlineReferences

Abstract
Many studies exploit variation in policy adoption timing across units as an instrument for treatment. This paper formalizes the underlying identification strategy as an instrumented difference-in-differences (DID-IV). In this design, a Wald-DID estimand, which scales the DID estimand of the outcome by the DID estimand of the treatment, captures the local average treatment effect on the treated (LATET). We extend the canonical DID-IV design to multiple period settings with the staggered adoption of the instrument across units. Moreover, we propose a credible estimation method in this design that is robust to treatment effect heterogeneity. We illustrate the empirical relevance of our findings, estimating returns to schooling in the United Kingdom. In this application, the two-way fixed effects instrumental variable regression, the conventional approach to implement DID-IV designs, yields a negative estimate. By contrast, our estimation method indicates a substantial gain from schooling.
Abstract
[…]
Keywords: difference-in-differences, instrumental variable, local average treatment effect, returns to education
1   Introduction
To identify the effect of a treatment on an outcome, many studies have exploited variation in policy adoption timing across units as an instrument for treatment. For example, to estimate returns to schooling in Indonesia, Duflo (2001) leverages variation in the timing of introduction of school construction program across regions as an instrument for education attainment. Similarly, to estimate the causal link between parents’ and children’s education attainment, Black et al. (2005) leverage variation in the timing of implementation of school reforms across municipalities as an instrument for parents’ education attainment. Notably, the identification strategy underlying these studies is similar to difference-in-differences (DID) designs, in that the timing variation of a policy shock is used for identifying the treatment effects. Meanwhile, it also differs from DID designs, in that the variation is used to construct an instrument rather than a treatment.
In this paper, we formalize the underlying identification strategy as an instrumented difference-in-differences (DID-IV). We define the target parameter and identifying assumptions in this design, and develop a credible estimation method that is robust to treatment effect heterogeneity. We illustrate the empirical relevance of our findings with the setting of Oreopoulos (2006), estimating returns to schooling in the United Kingdom. In this empirical application, we demonstrate that the estimation method matters in this design in practice.
First, we consider a simple setting with two periods and two groups: some units are not exposed to the instrument during the two periods (unexposed group), while some units become exposed in the second period (exposed group). [^1] Note that this two-period/two-group ($2\times 2$) setting is already considered in de Chaisemartin (2010) and Hudson et al. (2017), and our $2\times 2$ DID-IV design is mainly based on these studies. In this paper, we revisit $2\times 2$ DID-IV designs for two reasons. First, we aim to complement de Chaisemartin (2010) and Hudson et al. (2017) along two dimensions. Specifically, we introduce the instrument path into $2\times 2$ DID-IV designs, and uncover the additional identifying assumption not give in the previous literature. In addition, we investigate the treatment adoption behavior across units, and clarify the interpretation of the parallel trends assumption in DID-IV designs. Second, we aim to compare $2\times 2$ DID-IV designs to Fuzzy DID designs considered in de Chaisemartin and D’Haultfœuille (2018). While both de Chaisemartin (2010) and de Chaisemartin and D’Haultfœuille (2018) treat the same setting, de Chaisemartin and D’Haultfœuille (2018) formalize $2\times 2$ DID-IV designs differently, calling them Fuzzy DID designs. In this setting, our DID-IV design mainly comprises a monotonicity assumption, and parallel trends assumptions in the treatment and the outcome between exposed and unexposed groups. Our target parameter is the local average treatment effect on the treated (LATET); this parameter measures the treatment effects, for those who belong to an exposed group and are induced to treatment by instrument in the second period. We show that in this design, a Wald-DID estimand, which scales the DID estimand of the outcome by the DID estimand of the treatment, captures the LATET.
In DID-IV designs, the parallel trends assumption in the outcome requires careful assessment because it does not depend on the potential outcomes, but depends on the outcome under no exposure to the instrument, which we call an unexposed outcome. Give this, we next clarify the interpretation of this assumption in DID-IV settings. Specifically, we first point out that, in DID-IV designs, time also plays the role of instrument; that is, it affects one’s treatment choices over time in the absence of a policy shock (instrument). Following the terminology in Imbens and Angrist (1994), we divide the population into time groups defined by the potential treatment choices without instrument during the two periods: time always-takers, time never-takers, time compliers, and time defiers. Subsequently, we show that the average time trend in an unexposed outcome equals a weighted average of the effect of time on an outcome through treatment in each time group. We call this an expected time gain, and show that the parallel trends assumption in the outcome requires the same expected time gain between exposed and unexposed groups.
de Chaisemartin and D’Haultfœuille (2018) (hereafter, “dCDH”) formalize $2\times 2$ DID-IV designs differently, calling them Fuzzy DID designs. The main difference between this article and dCDH lies in the definition of the target parameter: we focus on the LATET, while dCDH focus on the switcher local average treatment effect on the treated (SLATET); this parameter measures the treatment effects, for those who belong to an exposed group and start receiving the treatment in the second period. Because the target parameter differs from each other, the identifying assumptions in dCDH are also different from those used here.
Motivated by the differences between this article and dCDH, we next examine the detailed connections between DID-IV and Fuzzy DID, and point out the issues inherent in Fuzzy DID. First, we show that dCDH’s identifying assumptions ex ante impose strong restrictions on treatment adoption behavior across units, and these restrictions are asymmetric between unexposed and exposed groups. Second, under these restrictions, we show that dCDH’s target parameter, the SLATET, can be decomposed into a weighted average of two different causal parameters: one measures the treatment effects among the sub-population of the compliers and the other measures the treatment effects among the time compliers. This decomposition result has a concerning implication; the SLATET may not be policy-relevant even when the instrument is based on the policy change of interest to the researcher. This is because the SLATET is always contaminated by the treatment effects among the time compliers, who are affected by time but not affected by instrument.
Next, we extend the canonical DID-IV design to multiple period settings with the staggered adoption of the instrument across units. In most DID-IV applications, researchers exploit variation in policy adoption timing across units in more than two periods, instrumenting for the treatment with the natural variation. The instrument is constructed, for instance from the staggered adoption of school reforms across municipalities or across countries (e.g. Oreopoulos (2006), Lundborg et al. (2014), and Meghir et al. (2018)), phased-in introduction of head starts across states (e.g. Johnson and Jackson (2019)), or gradual adoption of broadband internet programs across municipalities (e.g. Akerman et al. (2015), Bhuller et al. (2013)). We refer to the underlying identification strategy as a staggered DID-IV design, and establish the target parameter and identifying assumptions. Specifically, we first partition units into mutually exclusive and exhaustive cohorts by the initial exposure date of the instrument. We then define our target parameter as the cohort specific local average treatment effect on the treated (CLATT); this parameter is a natural generalization of the LATET in $2\times 2$ DID-IV designs, and measures the treatment effects, for those who belong to cohort $e$ and are the compliers at a given relative period $l$ starting from the initial exposure to the instrument. Finally, we extend the identification assumptions in $2\times 2$ DID-IV designs to multiple period settings, and state the interpretation of each assumption.
We extend our DID-IV designs in several directions. We first show that our DID-IV framework can be applied when we have a non-binary, ordered treatment. We also consider extensions to repeated cross sections. Lastly, we consider the case where we introduce the treatment path in potential outcomes in $2\times 2$ DID-IV designs.
Finally, we propose a regression-based method to consistently estimate our target parameter in staggered DID-IV designs under heterogeneous treatment effects. In practice, when researchers implicitly rely on a staggered DID-IV design, they usually implement this design via two-way fixed effects instrumental variable (TWFEIV) regressions (e.g. Johnson and Jackson (2019), Lundborg et al. (2014), Black et al. (2005), Akerman et al. (2015), and Bhuller et al. (2013)). In companion paper (Miyaji (2024)), however, we show that in more than two periods, the TWFEIV estimand generally fails to summarize the treatment effects if the effect of the instrument on the treatment or the outcome is not stable over time. Our proposed method avoids this issue and is robust to treatment effect heterogeneity. The estimation procedure consists of two steps: we subset the data that contain only two cohorts and two periods and then, in each data set, we run the TWFEIV regression. We call this a stacked two stage least squares (STS) regression and ensure its validity. Following Callaway and Sant’Anna (2021), we propose the weighting scheme to summarize the treatment effects, in which the weight reflects the share of compliers in a given relative period $l$ in cohort $e$. We also discuss the procedure of pre-trends tests to check the validity of the parallel trends assumptions in DID-IV designs.
We illustrate our findings with the setting of Oreopoulos (2006), who estimates returns to schooling in the UK, exploiting variation in the timing of implementation of school reforms between Britain and North Ireland as an instrument for education attainment. In this application, we first assess the plausibility of the DID-IV identification strategy implicitly imposed by Oreopoulos (2006). We then estimate the TWFEIV regression in the author’s setting. We find that the TWFEIV estimate is strictly negative and not significantly different from zero. Finally, we use our estimation method to reassess returns to schooling in the UK. We find that our STS estimates are all positive in each relative period after school reform, and our weighting scheme yields a more plausible estimate than the TWFEIV estimate. Specifically, our weighting estimate indicates roughly a $20\%$ gain from schooling in the UK.
Overall, this paper provides a new econometric framework by combining the IV techniques with DID designs. Our DID-IV design can be applied to various empirical settings. First, it can cope with the general adoption process of the treatment, which the canonical DID designs cannot, when we are interested in the effect of a treatment on an outcome. For instance, suppose that some units already receive the treatment in period 00 before the policy shock that distributes the treatment in period $1$. In this setting, the canonical DID design is infeasible in practice unless these units are discarded ex ante from the data. By contrast, our DID-IV design can be directly applied to this setting, allowing us to identify the LATET. In other words, we can view that DID-IV is DID with noncompliance of the assigned treatment.
Second, our DID-IV design can yield valid counterfactuals to estimate the treatment effect when no control group exists under DID designs. For instance, suppose that the treatment is education attainment and we are interested in returns to schooling as in Duflo (2001). In this setting, the canonical DID design generally fails to construct the control group, as the data would not contain the units with zero education attainment in reality. By contrast, our DID-IV design can partition units into exposed and unexposed groups depending on whether the unit is exposed to the policy shock (instrument). Once these groups are constructed, one can identify the local average treatment effect by comparing the evolution of the treatment and the outcome between the two groups.
Finally, our DID-IV design allows a researcher to address the endogeneity issue in a given application. In many empirical settings, a researcher often fails to identify the treatment effects due to the endogenous adoption of the treatment or the potential violation of the parallel trends assumption under DID designs. In this situation, our DID-IV design offers an alternative identification strategy, constructing an instrument based on the timing variation of a policy shock. Notably, in this design, the identifying variation comes not from the random assignment of a policy shock across units, but mainly comes from the parallel trends assumption in the treatment and the outcome. This identification strategy would be less demanding than IV designs in certain applications, and enhance the internal and external validity of the IV estimate in practice (Imbens and Angrist (1994); Heckman and Vytlacil (2001); Heckman and Vytlacil (2005)).
The rest of the paper is organized as follows. The following subsection discusses the related literature. Section 2 establishes DID-IV designs in two periods and two groups settings. Section 3 formalizes the target parameter and identifying assumptions in staggered DID-IV designs. Section 4 contains extensions. Section 5 presents our estimation method. Section 6 presents our empirical application. Section 7 concludes. All proofs are given in the Appendix.
1.1   Related literature
Our paper is related to recent DID-IV literature (de Chaisemartin (2010); Hudson et al. (2017); de Chaisemartin and D’Haultfœuille (2018); de Chaisemartin et al. (2024)) and contributes to this literature in three important ways. [^2] While Ye et al. (2023) also consider the “instrumented difference-in-differences”, their target parameter is the average treatment effect (ATE), and they make strong assumptions for the Wald-DID estimand to capture this parameter. We therefore see that Ye et al. (2023) are distinct from the recent DID-IV literature.
First, this paper investigates the detailed connections between DID-IV and Fuzzy DID considered in de Chaisemartin and D’Haultfœuille (2018). In econometrics, a pioneering work formalizing $2\times 2$ DID-IV designs is de Chaisemartin (2010), who shows that if parallel trends assumptions in the treatment and the outcome, and a monotonicity assumption hold, the Wald-DID estimand identifies the local average treatment effect on the treated (LATET). Following de Chaisemartin (2010), Hudson et al. (2017) also consider $2\times 2$ DID-IV designs with non-binary, ordered treatment settings. Building on the work in de Chaisemartin (2010), however, de Chaisemartin and D’Haultfœuille (2018) formalize $2\times 2$ DID-IV designs differently and call them Fuzzy DID designs.
In this paper, we first formalize $2\times 2$ DID-IV designs, complementing de Chaisemartin (2010) and Hudson et al. (2017) in two ways. Specifically, the framework we consider is mainly based on de Chaisemartin (2010) and Hudson et al. (2017). Nevertheless, we introduce the instrument path into $2\times 2$ DID-IV designs and uncover the additional identifying assumption not given in the previous literature. We also investigate the treatment adoption behavior across units, and clarify the interpretation of the parallel trends assumption in the outcome in DID-IV designs.
Given the identification results and the terminology developed in this paper, we then compare DID-IV to Fuzzy DID, and point out the issues inherent in Fuzzy DID. The discussion organizes the relationship between the two designs, which essentially treat the same setting, in the previous literature, and indicates the relative advantages of DID-IV over Fuzzy DID.
Second, this paper extends $2\times 2$ DID-IV designs to multiple period settings with the staggered adoption of the instrument across units, calling it a staggered DID-IV design. In practice, empirical researchers often leverage variation in policy adoption timing across units as an instrument for treatment in more than two periods (e.g., Black et al. (2005), Lundborg et al. (2014), and Johnson and Jackson (2019)). No previous study, however, has extended $2\times 2$ DID-IV designs to such important settings. In this paper, we formalize the underlying identification strategy as staggered DID-IV designs, and establish the target parameter and the identifying assumptions. Our staggered DID-IV designs allow the practitioners to estimate the local average treatment effect even when the treatment adoption is endogenous over time.
Finally, this paper provides a credible estimation method in staggered DID-IV designs under heterogeneous treatment effects. When empirical researchers implicitly rely on the staggered DID-IV design in practice, they commonly implement this design via TWFEIV regressions (e.g. Johnson and Jackson (2019), Lundborg et al. (2014), Black et al. (2005), Akerman et al. (2015), and Bhuller et al. (2013)). In companion paper (Miyaji (2024)), however, we show that in more than two periods, the TWFEIV estimand generally fails to summarize the treatment effects if the effect of the instrument on the outcome or the treatment evolves over time. [^3] See also de Chaisemartin and D’Haultfœuille (2020), who decompose the numerator and denominator of the TWFEIV estimand separately, and point out the issue of interpreting this estimand causally in Section 3.4 of their Web Appendix. Our proposed estimation method would serve as an alternative to the TWFEIV estimator and make DID-IV designs more credible in a given application.
In work related to this paper, de Chaisemartin et al. (2024) also consider DID-IV designs following de Chaisemartin (2010) and Hudson et al. (2017). This paper is distinct from theirs in three ways. First, they consider continuous treatment with continuous but static instrument, while we consider binary or ordered treatment with binary but dynamic instrument. Second, they consider multiple period settings with potentially non-staggered instrument, while we consider multiple periods settings with staggered instrument. Third, they consider the Fuzzy DID as a special case of DID-IV, while we formally discuss the relationship between the two designs.
Our paper is also related to a growing DID literature in two ways. First, this paper provides an alternative identification strategy to DID designs, addressing cases where the parallel trends assumption is not satisfied or no control group exists in a given application. When empirical researchers rely on DID designs and run two-way fixed effects regressions, they often concern that the parallel trends assumption is not plausible in practice. To address this, they often turn to TWFEIV regressions, exploiting the timing variation of a policy shock as an instrument for treatment (e.g. Miller and Segal (2019)). In this paper, we formalize the underlying identification strategy as instrumented difference-in-differences. Our DID-IV design is a natural extension of DID design, in that the timing variation is used to construct an instrument rather than a treatment.
Second, in line with the DID literature, this paper develops a reliable estimation method in DID-IV designs in the presence of heterogeneous treatment effects. Recently, several studies have pointed out the issue of implementing DID designs via two-way fixed effects regressions, or its dynamic specifications under heterogeneous treatment effects. (Athey and Imbens (2022); Borusyak et al. (2024); Callaway and Sant’Anna (2021); de Chaisemartin and D’Haultfœuille (2020); Goodman-Bacon (2021); Imai and Kim (2021); Sun and Abraham (2021)). Some of them have proposed a credible estimation method that delivers a sensible estimand and is robust to treatment heterogeneity. Along with the recent DID literature, this paper proposes an alternative estimation method to the TWFEIV regression, and illustrate its usefulness through an empirical application.
2   DID-IV in two time periods
In this section, we formalize an instrumented difference-in-differences (DID-IV) in two-period/two-group settings. We first establish the target parameter and the identifying assumptions in this design. We then investigate the treatment adoption behavior across units, and clarify the interpretation of the parallel assumption in the outcome. Finally, we describe the connections between DID-IV and Fuzzy DID proposed by de Chaisemartin and D’Haultfœuille (2018).
2.1   Set up
We introduce the notation we use throughout Section 2. We consider a panel data setting with two periods and $N$ units. For any random variable $R$, we denote $\mathcal{S}(R)$ to be its support. For each $i\in\{1,\dots,N\}$ and $t\in\{0,1\}$, let $Y_{i,t}$ denote the outcome, and $D_{i,t}\in\{0,1\}$ denote the treatment status: $D_{i,t}=1$ if unit $i$ receives the treatment in period $t$ and $D_{i,t}=0$ if unit $i$ does not receive the treatment. Let $Z_{i,t}\in\{0,1\}$ denote the instrument status: $Z_{i,t}=1$ if unit $i$ is exposed to the instrument in period $t$ and $Z_{i,t}=0$ if unit $i$ is not exposed to the instrument in period $t$. Throughout Section 2, we assume that $\{Y_{i,0},Y_{i,1},D_{i,0},D_{i,1},Z_{i,0},Z_{i,1}\}_{i=1}^{N}$ are independent and identically distributed (i.i.d).
We introduce the path of the treatment and the instrument. Let $D_{i}=(D_{i,0},D_{i,1})$ be the treatment path and $Z_{i}=(Z_{i,0},Z_{i,1})$ the instrument path. We assume no one is exposed to the instrument in period $t=0$: $Z_{i,0}=0$ for all $i$; we refer to this as a sharp assignment of the instrument. We denote $E_{i}\in\{0,1\}$ as the group variable: $E_{i}=1$ if unit $i$ is exposed to the instrument in period $t=1$ (exposed group) and $E_{i}=0$ if unit $i$ is not exposed to the instrument in period $t=1$ (unexposed group). In contrast to the sharp assignment of the instrument, we allow the general adoption process of the treatment; that is, we assume the treatment path can take four values with non-zero probability: $\{(0,0),(0,1),(1,0),(1,1)\}\in\mathcal{S}(D)$.
In practice, researchers are interested in the effect of a treatment $D_{i,t}$ on an outcome $Y_{i,t}$, and the instrument $Z_{i,1}$ typically represents a program or policy shock that encourages people to adopt the treatment in period $t=1$. For instance, Duflo (2001) estimates returns to schooling in Indonesia, exploiting variation arising from a new school construction program across regions and cohorts as an instrument for education attainment.
Next, we introduce the potential outcomes framework. Let $Y_{i,t}(d,z)$ denote the potential outcome in period $t$ when unit $i$ receives the treatment path $d\in\mathcal{S}(D)$ and the instrument path $z\in\mathcal{S}(Z)$. Similarly, let $D_{i,t}(z)$ denote the potential treatment status in period $t$ when unit $i$ receives the instrument path $z\in\mathcal{S}(Z)$. We refer to $D_{i,t}((0,0))$ as unexposed treatment and $D_{i,t}((0,1))$ as exposed treatment. Since the treatment and the instrument takes only two values, one can write the observed treatments $D_{i,t}$ and outcomes $Y_{i,t}$ as follows.
$D_{i,t}=\sum_{z\in\mathcal{S}(Z)}\mathbf{1}\{Z_{i}=z\}D_{i,t}(z),\hskip 8.53581ptY_{i,t}=\sum_{z\in\mathcal{S}(Z)}\sum_{d\in\mathcal{S}(D)}\mathbf{1}\{Z_{i}=z,D_{i}=d\}Y_{i,t}(d,z).$
We make a no carryover assumption on the potential outcomes $Y_{i,t}(d,z)$.
Assumption 1 (No carryover assumption).
$\forall z\in\mathcal{S}(Z),\forall d\in\mathcal{S}(D),Y_{i,0}(d,z)=Y_{i,0}(d_{0},z),Y_{i,1}(d,z)=Y_{i,1}(d_{1},z),$
where $d=(d_{0},d_{1})$ is the generic element of the treatment path $D_{i}$.
Assumption 1 states that potential outcomes $Y_{i,t}(d,z)$ depend only on the current treatment status $d_{t}$ and the instrument path $z$. In the DID literature, several studies adopt this assumption in non-staggered treatment settings (e.g. de Chaisemartin and D’Haultfœuille (2020); Imai and Kim (2021)). [^4] When the treatment is binary and only two periods exist, one can explicitly introduce the treatment path in potential outcomes, as it takes only four values. We therefore consider this case in online Appendix Section A.3 as an extension. One can also weaken Assumption 1 in other cases, such as non-binary, ordered treatment, or multiple time periods, but it requires cumbersome notations and complicates the definition of the target parameter, being not practical for empirical applications.
Next, we introduce the group variable $G_{i}^{Z}\equiv(D_{i,1}((0,0)),D_{i,1}((0,1)))$ that describes the type of unit $i$ according to the response of $D_{i,1}$ on the instrument path $z$. Following the terminology in Imbens and Angrist (1994), we define $G_{i}^{Z}=(0,0)\equiv NT^{Z}$ as the never-takers, $G_{i}^{Z}=(1,1)\equiv AT^{Z}$ as the always-takers, $G_{i}^{Z}=(0,1)\equiv CM^{Z}$ as the compliers, and $G_{i}^{Z}=(1,0)\equiv DF^{Z}$ as the defiers.
Henceforth, we keep Assumption 1. In the next subsection, we define the target parameter in $2\times 2$ DID-IV designs.
2.2   Target parameter in $2\times 2$ DID-IV designs
In $2\times 2$ DID-IV designs, our target parameter is the local average treatment effect on the treated (LATET) in period $t=1$ defined below.
Definition.
The local average treatment effect on the treated (LATET) in period $t=1$ is
$LATET$ $\equiv E[Y_{i,1}(1)-Y_{i,1}(0)|E_{i}=1,D_{i,1}((0,1))>D_{i,1}((0,0))]$
$=E[Y_{i,1}(1)-Y_{i,1}(0)|E_{i}=1,CM^{Z}].$
This parameter measures the treatment effects in period $1$, for those who belong to an exposed group ($E_{i}=1$), and are induced to treatment by instrument in period $1$. In the DID-IV literature, de Chaisemartin (2010) considers the same target parameter, and Hudson et al. (2017) define the local treatment effect (LATE) in period $t=1$ as their target parameter, which is unconditional on $E_{i}$. The LATET has been also proposed in heterogeneous effects IV models with binary instrument (e.g. Słoczyński (2020); Słoczyński et al. (2022)).
We define the LATET as the target parameter for two reasons. First, this parameter would be particularly of interest if the instrument reflects a policy change of interest to the researcher (Heckman and Vytlacil (2001), Heckman and Vytlacil (2005)). Second, the LATET is a natural extension of the target parameter in DID designs, the so-called average treatment effects on the treated (ATT); both causal parameters measure the treatment effects among the units affected by a policy shock (the treatment in DID designs) and belonging to an exposed group (treatment group).
2.3   Identification assumptions in $2\times 2$ DID-IV designs
This subsection formalizes the identification assumptions in $2\times 2$ DID-IV designs. In two periods and two groups settings, a popular estimand is the ratio between the DID estimand of the outcome and the DID estimand of the treatment (Duflo (2001), Field (2007)):
$w_{DID}$ $=\frac{E[Y_{i,1}-Y_{i,0}|E_{i}=1]-E[Y_{i,1}-Y_{i,0}|E_{i}=0]}{E[D_{i,1}-D_{i,0}|E_{i}=1]-E[D_{i,1}-D_{i,0}|E_{i}=0]}.$
Following the terminology in de Chaisemartin and D’Haultfœuille (2018), we call this the Wald-DID estimand.
We consider the following identifying assumptions for the Wald-DID estimand to capture the LATET.
Assumption 2 (Exclusion restriction for potential outcomes).
$\forall z\in\mathcal{S}(Z),\forall d\in\mathcal{S}(D),\forall t\in\{0,1\},Y_{i,t}(d,z)=Y_{i,t}(d).$
This assumption requires that the instrument path does not directly affect potential outcomes other than through treatment. This assumption is common in the IV literature; see e.g., Imbens and Angrist (1994) and Abadie (2003). In the DID-IV literature, de Chaisemartin (2010) and Hudson et al. (2017) impose a similar assumption without introducing the instrument path.
Given Assumption 1 and Assumption 2, we can write the observed outcomes $Y_{i,t}$ as
$Y_{i,t}=D_{i,t}Y_{i,t}(1)+(1-D_{i,t})Y_{i,t}(0).$
Assumption 1 and Assumption 2 also allow us to introduce the notion of exposed and unexposed outcomes. For any $z\in\mathcal{S}(Z)$, let $Y_{i,t}(D_{i,t}(z))$ denote the outcome if the instrument path is $z$:
$Y_{i,t}(D_{i,t}(z))\equiv D_{i,t}(z)Y_{i,t}(1)+(1-D_{i,t}(z))Y_{i,t}(0).$
Hereafter, we refer to $Y_{i,t}(D_{i,t}((0,0)))$ as unexposed outcomes and $Y_{i,t}(D_{i,t}((0,1)))$ as exposed outcomes. [^5] Exposed and unexposed outcomes are not new concepts in econometrics. For example, under IV designs, we compare the expectation of the outcome with instrument and the one without instrument in the numerator of the wald estimand: $E[Y|Z=1]-E[Y|Z=0]$. This comparison corresponds to the one between the expectation of the exposed outcome and the expectation of the unexposed outcome; that is, $E[Y|Z=1]-E[Y|Z=0]=E[Y(D(1))|Z=1]-E[Y(D(0))|Z=0]$.
Next, we make the following monotonicity assumption as in Imbens and Angrist (1994).
Assumption 3 (Monotonicity assumption at period $t=1$).
$Pr(D_{i,1}((0,1))\geq D_{i,1}((0,0)))=1\hskip 8.53581pt\text{or}\hskip 5.69054ptPr(D_{i,1}((0,1))\leq D_{i,1}((0,0)))=1.$
This assumption requires that the instrument path affects the treatment choice at period $t=1$ in a monotone (uniform) way. It implies that the group variable $G_{i}^{Z}$ can take three values with non-zero probability. In the DID-IV literature, de Chaisemartin (2010) and Hudson et al. (2017) make the same assumption. Hereafter, we consider the type of monotonicity assumption that rules out the existence of the defiers $DF^{Z}$.
Assumption 4 (No anticipation in the first stage).
[…]
Assumption 4 requires that the potential treatment choice before the exposure to instrument is equal to the baseline treatment choice $D_{i,0}((0,0))$ in an exposed group. [^6] In the DID literature, recent studies impose the no anticipation assumption on untreated potential outcomes in several ways. Callaway and Sant’Anna (2021) and Sun and Abraham (2021) assume the average version of the no anticipation assumption, whereas Athey and Imbens (2022) assume it for all units $i$. Roth et al. (2023) take the intermediate approach: they adopt the no anticipation assumption for the treated units. The no anticipation assumption on potential treatment choices in Assumption 4 is in line with that of Roth et al. (2023). [^7] This assumption is not provided in the previous DID-IV literature: de Chaisemartin (2010) and Hudson et al. (2017) implicitly impose this assumption by writing observed treatment choice in period $t=0$ as $D_{i,0}(0)$. This assumption restricts the anticipatory behavior and would be plausible if the instrument path is ex ante not known for all the units in an exposed group.
Next, we impose the parallel trends assumptions in the treatment and the outcome. de Chaisemartin (2010) and Hudson et al. (2017) make the similar assumptions.
Assumption 5 (Parallel Trends Assumption in the treatment).
$E[D_{i,1}((0,0))-D_{i,0}((0,0))|E_{i}=0]=E[D_{i,1}((0,0))-D_{i,0}((0,0))|E_{i}=1].$
Assumption 5 is a parallel trends assumption in the treatment. This assumption requires that the expectation of the treatment between exposed and unexposed groups would have followed the same path if the assignment of the instrument had not occurred. For instance, in Duflo (2001), this assumption requires that the evolution of mean education attainment would have been the same between exposed and unexposed groups if the policy shock had not occurred during the two periods.
Assumption 6 (Parallel Trends Assumption in the outcome).
$E[Y_{i,1}(D_{i,1}((0,0)))-Y_{i,0}(D_{i,0}((0,0)))|E_{i}=0]$
$=$ $E[Y_{i,1}(D_{i,1}((0,0)))-Y_{i,0}(D_{i,0}((0,0)))|E_{i}=1].$
Assumption 6 is a parallel trends assumption in the outcome, requiring that the evolution of the unexposed outcome is, on average, the same between exposed and unexposed groups. For instance, in Duflo (2001), this assumption requires that the expectation of log annual earnings would have followed the same path from period 00 to period $1$ between exposed and unexposed groups in the absence of a policy shock.
Note that Assumption 6 depends on the unexposed outcome, and does not depend on the potential outcomes directly. In Subsection 2.4, however, we show that the average time trends of the unexposed outcome in Assumption 6 can be written as the weighted average of the expectation of the potential outcomes, and this assumption can be interpreted as the restriction on the time effect between exposed and unexposed groups.
When empirical researchers exploit variation arising from a policy shock as an instrument for treatment and use the Wald-DID estimand, they often refer to Assumption 5 and 6. For instance, Duflo (2001) estimates returns to schooling in Indonesia, relying on “the identification assumption that the evolution of wages and education across cohorts would not have varied systematically from one region to another in the absence of the program”.
Finally, we assume a relevance condition. This assumption guarantees that the Wald-DID estimand $w_{DID}$ is well defined. [^8] Note that Assumption 3, 4, 5, and 7 imply that the share of compliers in period $1$ is non-zero. $E[D_{i,1}-D_{i,0}|E_{i}=1]-E[D_{i,1}-D_{i,0}|E_{i}=0]=Pr(CM^{Z}|E_{i}=1)>0.$ This ensures that the LATET in period $t=1$ is also well defined.
Assumption 7 (Relevance condition).
$E[D_{i,1}-D_{i,0}|E_{i}=1]-E[D_{i,1}-D_{i,0}|E_{i}=0]>0.$
The theorem below shows that if Assumptions 1-7 hold, the Wald-DID estimand captures the LATET in period $1$.
Theorem 1.
If Assumptions 1-7 hold, the Wald-DID estimand $w_{DID}$ is equal to the LATET in period $t=1$; that is,
$w_{DID}=E[Y_{i,1}(1)-Y_{i,1}(0)|E_{i}=1,CM^{Z}].$
holds.
Proof.
See Appendix. ∎
2.4   Interpreting the parallel trends assumption in the outcome
In this subsection, we clarify the interpretation of the parallel trends assumption in the outcome in DID-IV designs. It is less clear than that in DID designs because this assumption does not depend on the untreated outcome, but depends on the unexposed outcome.
First, we introduce the additional notation. Let $G_{i}^{T}=(D_{i,0}((0,0)),D_{i,1}((0,0)))$ denote the group variable for unit $i$ that represents the treatment adoption process during the two periods if the instrument path is $z=(0,0)$. Recall that in DID-IV designs, the units are allowed to take the treatment both period 00 and $1$ without instrument. In this setting, one can see that time also plays the role of instrument; that is, time also affects one’s treatment choices during the two periods. Similar to the group variable $G_{i}^{Z}$ introduced in Subsection 2.1, we define $G_{i}^{T}=(0,0)\equiv NT^{T}$ as the time never-takers, $G_{i}^{T}=(1,1)\equiv AT^{T}$ as the time always-takers, $G_{i}^{T}=(0,1)\equiv CM^{T}$ as the time compliers, and $G_{i}^{T}=(1,0)\equiv DF^{T}$ as the time defiers.
Table 1: Exposed group (assigned to $z=(0,1)$)
observed counterfactual
$(D_{0}((0,0)),D_{1}((0,1)))$ $D_{1}((0,0))=1$ $D_{1}((0,0))=0$
$D_{0}((0,0))=1,D_{1}((0,1))=1$ $AT^{Z}\land AT^{T}$ $CM^{Z}\land DF^{T}$
$D_{0}((0,0))=1,D_{1}((0,1))=0$ $DF^{Z}\land AT^{T}$ $NT^{Z}\land DF^{T}$
$D_{0}((0,0))=0,D_{1}((0,1))=1$ $AT^{Z}\land CM^{T}$ $CM^{Z}\land NT^{T}$
$D_{0}((0,0))=0,D_{1}((0,1))=0$ $DF^{Z}\land CM^{T}$ $NT^{Z}\land NT^{T}$
Table 2: Unexposed group (assigned to $z=(0,0)$)
observed counterfactual
$(D_{0}((0,0)),D_{1}((0,0)))$ $D_{1}((0,1))=1$ $D_{1}((0,1))=0$
$D_{0}((0,0))=1,D_{1}((0,0))=1$ $AT^{Z}\land AT^{T}$ $DF^{Z}\land AT^{T}$
$D_{0}((0,0))=1,D_{1}((0,0))=0$ $CM^{Z}\land DF^{T}$ $NT^{Z}\land DF^{T}$
$D_{0}((0,0))=0,D_{1}((0,0))=1$ $AT^{Z}\land CM^{T}$ $DF^{Z}\land CM^{T}$
$D_{0}((0,0))=0,D_{1}((0,0))=0$ $CM^{Z}\land NT^{T}$ $NT^{Z}\land NT^{T}$
Notes: These tables represent mutually exclusive and exhaustive types under DID-IV designs. The types painted in gray color are excluded by Assumption 3 (monotonicity assumption).
By exploiting the group variables $G_{i}^{Z}$ and $G_{i}^{T}$, we can describe how the identifying assumptions in $2\times 2$ DID-IV designs restrict the treatment adoption behavior across units in exposed and unexposed groups, respectively. Table 1 and Table 2 show that we can partition units in exposed and unexposed groups into six mutually exclusive and exhaustive types under $2\times 2$ DID-IV designs. [^9] In an exposed group, we can observe $(D_{0}((0,0)),D_{1}((0,1)))$, but not $D_{1}((0,0))$ for each unit (see Table 1). We use Assumption 4 (no anticipation assumption in the first stage), replacing $D_{0}((0,1))$ with $D_{0}((0,0))$. In an unexposed group, we can observe $(D_{0}((0,0)),D_{1}((0,0)))$, but not $D_{1}((0,1))$ for each unit (see Table 2). If, for instance, the observed treatment choices of unit $i$ in an exposed group are $D_{i,0}((0,0))=0$ and $D_{i,1}((0,1))=1$, and the counterfactual treatment choice is $D_{i,1}((0,0))=0$, that unit belongs to the type $CM^{Z}\land NT^{T}$: unit $i$ is induced to the treatment by instrument but not by time. If the observed treatment choices of unit $i$ in an unexposed group are $D_{i,0}((0,0))=0$ and $D_{i,1}((0,0))=1$, and the counterfactual treatment choice is $D_{i,1}((0,1))=1$, that unit belongs to the type $AT^{Z}\land CM^{T}$: unit $i$ is induced to the treatment by time but not by instrument. Here, Assumption 3 (monotonicity assumption) excludes the defiers $DF^{Z}$ in exposed and unexposed groups respectively, written in gray color in both tables.
We now consider the interpretation of the parallel trends assumption in the outcome. Let $\Delta_{0}$ and $\Delta_{1}$ denote the average time trends of the unexposed outcome in exposed and unexposed groups, respectively.
$\Delta_{0}\equiv E[Y_{i,1}(D_{i,1}((0,0)))-Y_{i,0}(D_{i,0}((0,0)))|E_{i}=0],$
$\Delta_{1}\equiv E[Y_{i,1}(D_{i,1}((0,0)))-Y_{i,0}(D_{i,0}((0,0)))|E_{i}=1].$
Note that $\Delta_{0}$ is estimable by the observed change in the outcome in an unexposed group during the two periods. By contrast, $\Delta_{1}$ is fundamentally unobservable because exposed group is assigned to the instrument path $z=(0,1)$.
To make the interpretations of $\Delta_{0}$ and $\Delta_{1}$ clear, we introduce the additional notation. For each time group $g^{t}\in\{AT^{T},NT^{T},CM^{T},DF^{T}\}$, let $w_{g^{t},e}\equiv Pr[G_{i}^{T}=g^{t}|E_{i}=e]$ denote the population share in group $E_{i}=e\in\{0,1\}$. Let $\Delta_{g^{t},e}$ denote the expectation of the potential outcome path in the absence of a policy shock for time group $g^{t}$ in group $E_{i}=e$:
$\Delta_{AT^{T},e}\equiv E[Y_{i,1}(1)-Y_{i,0}(1)|AT^{T},E_{i}=e],$ $\Delta_{NT^{T},e}\equiv E[Y_{i,1}(0)-Y_{i,0}(0)|NT^{T},E_{i}=e],$
$\Delta_{CM^{T},e}\equiv E[Y_{i,1}(1)-Y_{i,0}(0)|CM^{T},E_{i}=e],$ $\Delta_{DF^{T},e}\equiv E[Y_{i,1}(0)-Y_{i,0}(1)|DF^{T},E_{i}=e].$
We can interpret each $\Delta_{g^{t},e}$ as the effect of time on the outcome through treatment for time group $g^{t}$ in group $E_{i}=e$. First, $\Delta_{AT^{T},e}$ and $\Delta_{NT^{T},e}$ only reflect the expectation of the time effect for treated and untreated potential outcomes, respectively.
$\Delta_{AT^{T},e}$ $=\underbrace{E[Y_{i,1}(1)-Y_{i,0}(1)|AT^{T},E_{i}=e]}_{\text{Timeeffect among $AT^{T}$}},\hskip 5.69054pt\Delta_{NT^{T},e}$ $=\underbrace{E[Y_{i,1}(0)-Y_{i,1}(0)|NT^{T},E_{i}=e]}_{\text{Timeeffect among $NT^{T}$}}.$
This is because the time always-takers $AT^{T}$ and the time never-takers $NT^{T}$ do not change their treatment status during the two periods. Next, $\Delta_{CM^{T},e}$ and $\Delta_{DF^{T},e}$ can be decomposed into two terms.
$\Delta_{CM^{T},e}$ $=E[Y_{i,1}(0)-Y_{i,0}(0)|CM^{T},E_{i}=e]+\underbrace{E[Y_{i,1}(1)-Y_{i,1}(0)|CM^{T},E_{i}=e]}_{\text{Selection gain among $CM^{T}$}},$
$\Delta_{DF^{T},e}$ $=E[Y_{i,1}(1)-Y_{i,0}(1)|DF^{T},E_{i}=e]+\underbrace{E[Y_{i,1}(0)-Y_{i,1}(1)|DF^{T},E_{i}=e]}_{\text{Selection gain among $DF^{T}$}}.$
The first terms in $\Delta_{CM^{T},0}$ and $\Delta_{DF^{T},0}$ represent the average time effect for untreated and treated potential outcomes in each time group in group $E_{i}=e$, respectively. The second term represents the selection gain in each time group in group $E_{i}=e$, respectively. The latter terms arise because the time compliers $CM^{T}$ and the time defiers $DF^{T}$ are affected by time and change their treatment status from period 00 to period $1$.
The following theorem establishes the relationship between $\Delta_{e}$ and $\Delta_{g^{t},e}$.
Theorem 2.
Suppose Assumptions 1-4 hold, then
$\Delta_{e}=\sum_{g^{t}}w_{g^{t},e}\Delta_{g^{t},e},$
where the weight $w_{g^{t},e}$ is non-negative and sums to one; that is, $\sum_{g^{t}}w_{g^{t},e}=1$ holds.
Proof.
This directly follows from the Law of Iterated Expectation:
$\Delta_{e}$ $=E[Y_{1}(1)-Y_{0}(1)|AT^{T},E=e]Pr(AT^{T}|E=e)$
$+E[Y_{1}(0)-Y_{0}(0)|NT^{T},E=e]Pr(NT^{T}|E=e)$
$+E[Y_{1}(1)-Y_{0}(0)|CM^{T},E=e]Pr(CM^{T}|E=e)$
$+E[Y_{1}(0)-Y_{0}(1)|DF^{T},E=e]Pr(DF^{T}|E=e)$
$=\sum_{g^{t}}w_{g^{t},e}\Delta_{g^{t},e}.$
Clearly, the weight $w_{g^{t},e}$ sums to one by construction. ∎
Theorem 2 shows that the average time trend of the unexposed outcome in group $e$ can be written as the weighted average of the expectation of the potential outcomes, and can be interpreted as measuring the convex average of the time effect for each time group $g^{t}$ in group $e$. The weight assigned to each $\Delta_{g^{t},e}$ is natural in that it reflects the population share in each time group $g^{t}$ in group $e$. Hereafter, we call $\Delta_{e}$ an expected time gain in group $e$.
From Theorem 2, we can interpret that the parallel trends assumption in the outcome (Assumption 6), $\Delta_{0}=\Delta_{1}$, requires the same expected time gain between exposed and unexposed groups; that is, the effects of time on the outcome through treatment are the same between the two groups.
Assumption 6 is required for DID-IV identification strategy because in the numerator of the Wald-DID estimand, we should isolate the effect of the instrument on the outcome through treatment from that of time in an exposed group. The underlying idea to capture the former effect is parallel to that in DID designs: we purge the latter effect by exploiting the observed time trend of the outcome in the group that did not receive the policy shock (instrument) in the second period.
Finally, we provide an empirical example to illustrate the interpretation of each notation and Assumption 6 in practice.
Example.
Suppose that the treatment $D_{i,t}\in\{0,1\}$ represents participation in a job training program, the outcome $Y_{i,t}$ represents real earnings, and the instrument $Z_{i,1}\in\{0,1\}$ represents a policy shock that encourages people to receive the treatment in the second period. [^10] Assume that the treatment is non-staggered and the treatment and the instrument are not randomized across units. In this setting, the instrumental variable strategy nor the canonical DID designs are directly applicable. In this setting, the time compliers $CM^{T}$ are the units who start adopting the program in the second period in the absence of the policy shock, and $w_{CM^{T},0}$ measures its share in an unexposed group. $\Delta_{CM^{T},0}$ represents the average time trends of real earnings among the time compliers in an unexposed group, and can be decomposed into the time effects on real earnings without policy shock, and the causal effects of program participation on real earnings in the second period, as their treatment adoption behavior is affected by time. Assumption 6 requires that the average time trends of real earnings are the same without policy shock between exposed and unexposed groups. This assumption can be interpreted as the restriction on the effect of time on real earnings through program participation between the two groups.
2.5   Comparing DID-IV to Fuzzy DID
de Chaisemartin and D’Haultfœuille (2018) also investigate the identification assumptions for the Wald-DID estimand to capture the causal effects under the same setting as in this article. [^11] The Wald-DID estimand considered in dCDH corresponds to ours in this article. In their empirical application, dCDH revisit Duflo (2001), in which the group variable $E_{i}$ ($G_{i}$ in dCDH) takes one if unit $i$ belongs to the region that are highly exposed to the policy shock. However, they formalize $2\times 2$ DID-IV designs differently, calling them Fuzzy DID designs. In this section, we compare DID-IV to Fuzzy DID and point out the issues embedded in Fuzzy DID in an overview of the detailed discussion provided in online Appendix Section E.
The main difference between dCDH and this article lies in the definition of the target parameter: we focus on the LATET, while dCDH focus on the switcher local average treatment effect on the treated (SLATET) defined below.121212de Chaisemartin and D’Haultfœuille (2018) consider the same situation in repeated cross section settings. Therefore, we rewrite their target parameter by the notation used in this article.
Definition.
The switcher local average treatment effect on the treated (SLATET) is
$SLATET$ $\equiv E[Y_{i,1}(1)-Y_{i,1}(0)|E_{i}=1,D_{i,1}((0,1))>D_{i,0}((0,1))]$
$=E[Y_{i,1}(1)-Y_{i,1}(0)|E_{i}=1,SW],$
where the switcher (SW) is the units who become treated in period $1$.
This parameter measures the treatment effects among the units who belong to an exposed group ($E_{i}=1$) and switch to receive the treatment in period $1$ (SW). Because the target parameter is different from each other, the identifying assumptions in Fuzzy DID designs are also different from those in $2\times 2$ DID-IV designs. Furthermore, dCDH argue that the Wald-DID estimand identifies the SLATET only if the treatment effect is stable over time, and propose the alternative estimands that do not require the strong restriction on treatment effect heterogeneity.
In online Appendix Section E, we first show that the identifying assumptions in Fuzzy DID designs ex ante impose strong restrictions on treatment adoption behavior across units and these restrictions are asymmetric between exposed and unexposed groups. Specifically, we examine how dCDH’s identifying assumptions exclude the types in the two groups and construct Tables 3-4, which are in line with Tables 1-2.131313Because dCDH consider repeated cross section settings, we can observe either $D_{0}(z)$ or $D_{1}(z)$ for unit $i$ in group $z$ in each table.
Comparing Tables 1-2 with Tables 3-4, one can figure out the difference between dCDH’s restrictions and ours. In our DID-IV design, we exclude only the defiers $DF^{Z}$ in exposed and unexposed groups by a monotonicity assumption. By contrast, dCDH exclude the defiers $DF^{Z}$ and the type $NT^{Z}\land DF^{T}$ in an exposed group, and the defiers $DF^{Z}$, the time defiers $CM^{T}$ and the time compliers $DF^{T}$ in an unexposed group.
Table 3: Exposed group ($z=1$)
observed counterfactual
$D_{0}(0)$  or  $D_{1}(1)$ $D_{1}(0)=1$ $D_{1}(0)=0$
$D_{0}(0)=1,D_{1}(1)=1$ $AT^{Z}\land AT^{T}$ $CM^{Z}\land DF^{T}$
$D_{0}(0)=1,D_{1}(1)=0$ $DF^{Z}\land AT^{T}$ $NT^{Z}\land DF^{T}$
$D_{0}(0)=0,D_{1}(1)=1$ $AT^{Z}\land CM^{T}$ $CM^{Z}\land NT^{T}$
$D_{0}(0)=0,D_{1}(1)=0$ $DF^{Z}\land CM^{T}$ $NT^{Z}\land NT^{T}$
Table 4: Unexposed group ($z=0$)
observed counterfactual
$D_{0}(0)$  or  $D_{1}(0)$ $D_{1}(1)=1$ $D_{1}(1)=0$
$D_{0}(0)=1,D_{1}(0)=1$ $AT^{Z}\land AT^{T}$ $DF^{Z}\land AT^{T}$
$D_{0}(0)=1,D_{1}(0)=0$ $CM^{Z}\land DF^{T}$ $NT^{Z}\land DF^{T}$
$D_{0}(0)=0,D_{1}(0)=1$ $AT^{Z}\land CM^{T}$ $DF^{Z}\land CM^{T}$
$D_{0}(0)=0,D_{1}(0)=0$ $CM^{Z}\land NT^{T}$ $NT^{Z}\land NT^{T}$
Notes: These tables represent mutually exclusive and exhaustive types under Fuzzy DID designs. The types painted in gray color are excluded by dCDH’s identifying assumptions.
Next, under Tables 3-4, we show that dCDH’s target parameter, the SLATET, can be decomposed into a weighted average of two different causal parameters. The one parameter measures the treatment effects among the type $CM^{Z}\land AT^{T}$, who are the sub-population of the compliers. The other causal parameter measures the treatment effects among the time compliers $CM^{T}$. [^14] Intuitively, we cannot identify this parameter in DID-IV settings because the untreated potential outcomes in period $t=1$ among the time compliers are not observed in both groups (see Tables 3-4). This decomposition result has a concerning implication; the SLATET may not be policy-relevant even when the instrument is based on the policy change of interest to the researcher, as this parameter is always contaminated by the treatment effects among the time compliers, who are affected by time but not affected by the instrument.
Finally, we revisit dCDH’s argument regarding the use of the Wald-DID estimand, and show that this problem is less serious than one might expect, and does not arise once a researcher imposes Assumption 6 instead of the parallel trends assumption in the untreated outcome.
First, we show that their argument is based on the parallel trends assumption in DID designs. Recall that in our DID-IV design, we do not make the parallel trends assumption in the untreated outcome, but make the parallel trends assumption in the unexposed outcome. Because some units are allowed to adopt the treatment without instrument in DID-IV settings, the former type of parallel trends assumption is generally not sufficient to capture the average time trend of the outcome in an unexposed group. We show that this fact leads dCDH to impose the stable treatment effect assumption for the Wald-DID estimand to identify the SLATET.
Next, we argue that the different type of the parallel trends assumption, Assumption 6, is suitable for DID-IV settings, and the parallel trends assumption in the untreated outcome would be difficult to justify in practice for three reasons. First, as already noted, this assumption is not sufficient to identify the average time trends of the outcome even in an unexposed group. Second, in most applications of the DID-IV method, we cannot impute the untreated potential outcomes in general (e.g., Duflo (2001), Black et al. (2005)). For instance, in Duflo (2001), this assumption requires the data to contain units with zero education attainment, which is unrealistic in practice. Finally, in most of the DID-IV settings, we cannot test this assumption using pre-exposed period data, as some units can potentially adopt the treatment before period 00.
3   DID-IV in multiple time periods
We now extend the $2\times 2$ DID-IV design to multiple period settings with the staggered adoption of the instrument across units (Black et al. (2005); Bhuller et al. (2013) Lundborg et al. (2014); and Meghir et al. (2018)). We call it a staggered DID-IV design, and establish the target parameter and identifying assumptions.
3.1   Set up
We introduce the notation we use throughout Section 3 to Section 5. We consider a panel data setting with $T$ periods and $N$ units. For each $i\in\{1,\dots N\}$ and $t\in\{1,\dots,T\}$, let $Y_{i,t}$ denote the outcome, $D_{i,t}\in\{0,1\}$ denote the treatment status, and $Z_{i,t}\in\{0,1\}$ denote the instrument status. Let $D_{i}=(D_{i,1},\dots,D_{i,T})$ and $Z_{i}=(Z_{i,1},\dots,Z_{i,T})$ denote the path of the treatment and the instrument for unit $i$, respectively. Throughout Section 3 to Section 5, we assume that $\{Y_{i,t},D_{i,t},Z_{i,t}\}_{t=1}^{T}$ are i.i.d.
We make the following assumption about the assignment process of the instrument.
Assumption 8 (Staggered adoption for $Z_{i,t}$).
For $s<t$, $Z_{i,s}\leq Z_{i,t}$ where $s,t\in\{1,\dots T\}$.
Assumption 8 requires that once units start getting exposed to the instrument, units remain exposed to that instrument. In the DID literature, several studies make a similar assumption for the treatment and call it the “staggered treatment adoption” (e.g. Athey and Imbens (2022); Callaway and Sant’Anna (2021); and Sun and Abraham (2021)).
Given Assumption 8, one can uniquely characterize one’s instrument path by the initial exposure date of the instrument, denoted as $E_{i}=\min\{t:Z_{i,t}=1\}$. If unit $i$ is not exposed to the instrument for all time periods, we define $E_{i}=\infty$. Based on the initial exposure period $E_{i}$, one can uniquely partition units into mutually exclusive and exhaustive cohorts $e$ for $e\in\{1,2,\dots,T,\infty\}$.
Similar to the two periods setting in Subsection 2.1, we allow the general adoption process for the treatment: the treatment can potentially turn on/off repeatedly over time. de Chaisemartin and D’Haultfœuille (2020) and Imai and Kim (2021) also consider the same setting in the recent DID literature.
Next, we introduce the potential outcomes framework in multiple time periods. Let $Y_{i,t}(d,z)$ denote the potential outcome in period $t$ when unit $i$ receives the treatment path $d\in\mathcal{S}(D)$ and the instrument path $z\in\mathcal{S}(Z)$. Similarly, let $D_{i,t}(z)$ denote the potential treatment status in period $t$ when unit $i$ receives the instrument path $z\in\mathcal{S}(Z)$.
Assumption 8 allows us to rewrite $D_{i,t}(z)$ by the initial exposure date $E_{i}=e$. Let $D_{i,t}^{e}$ denote the potential treatment status in period $t$ if unit $i$ is first exposed to the instrument in period $e$. Let $D_{i,t}^{\infty}$ denote the potential treatment status in period $t$ if unit $i$ is never exposed to the instrument. We call $D_{i,t}^{\infty}$ the “never exposed treatment”. Since the adoption date of the instrument uniquely pins down one’s instrument path, we can write the observed treatment status $D_{i,t}$ for unit $i$ in period $t$ as
$D_{i,t}=D_{i,t}^{\infty}+\sum_{1\leq e\leq T}(D_{i,t}^{e}-D_{i,t}^{\infty})\cdot\mathbf{1}\{E_{i}=e\}.$
We define the effect of an instrument on treatment for unit $i$ in period $t$ as the difference between the observed treatment status to the never exposed treatment status: $D_{i,t}-D_{i,t}^{\infty}$. We refer to $D_{i,t}-D_{i,t}^{\infty}$ as the individual exposed effect in the first stage. [^15] In the DID literature, Callaway and Sant’Anna (2021) and Sun and Abraham (2021) define the effect of a treatment on an outcome in the same fashion.
Next, we introduce the group variable that describes the type of unit $i$ in period $t$, based on the reaction of potential treatment choices in period $t$ to the instrument path $z$. Let $G_{i,e,t}\equiv(D_{i,t}^{\infty},D_{i,t}^{e})$ ($t\geq e$) be the group variable in period $t$ for unit $i$ and the initial exposure date $e$. Following the terminology in section 2.1, we define $G_{i,e,t}=(0,0)\equiv NT_{e,t}$ as the never-takers, $G_{i,e,t}=(1,1)\equiv AT_{e,t}$ as the always-takers, $G_{i,e,t}=(0,1)\equiv CM_{e,t}$ as the compliers and $G_{i,e,t}=(1,0)\equiv DF_{e,t}$ as the defiers in period $t$ and the initial exposure date $e$.
Finally, we make a no carryover assumption on potential outcomes $Y_{i,t}(d,z)$.
Assumption 9 (No carryover assumption in multiple time periods).
$\forall z\in\mathcal{S}(Z),\forall d\in\mathcal{S}(D),\forall t\in\{1,\dots,T\},Y_{i,t}(d,z)=Y_{i,t}(d_{t},z).$
Henceforth, we keep Assumption 8 and Assumption 9. In the next section, we define the target parameter in staggered DID-IV designs.
3.2   Target parameter in staggered DID-IV designs
In staggered DID-IV designs, our target parameter is the cohort specific local average treatment effect on the treated (CLATT) defined below.
Definition.
The cohort specific local average treatment effect on the treated (CLATT) at a given relative period $l$ from the initial adoption of the instrument is
$CLATT_{e,e+l}$ $=E[Y_{i,e+l}(1)-Y_{i,e+l}(0)|E_{i}=e,D_{i,e+l}^{e}>D_{i,e+l}^{\infty}]$
$=E[Y_{i,e+l}(1)-Y_{i,e+l}(0)|E_{i}=e,CM_{e,e+l}].$
Each CLATT is a natural generalization of the LATET in Subsection 2.2 and suitable for the setting of the staggered instrument adoption. This parameter measures the treatment effects at a given relative period $l$ from the initial exposure date $E_{i}=e$, for those who belong to cohort $e$, and are the compliers $CM_{e,e+l}$ who are induced to treatment by instrument in period $e+l$. Each CLATT can potentially vary across cohorts and over time because it depends on cohort $e$, relative period $l$, and the compliers $CM_{e,e+l}$. [^16] Note that the composition of the compliers can overlap over the post-exposure periods in cohort $e$. For instance, some units in cohort $e$ may belong to $CM_{e,e+l}$ for all the relative periods ($l\geq e$): they are affected by the instrument for all post-exposure periods.
This parameter would be policy-relevant if the instrument is based on the policy change of interest to the researcher. For example, in a job training program setting considered in Subsection 2.5, each CLATT would be of particular interest if the instrument $Z_{i,t}$ represents an important policy lever for the researchers (e.g., nudge); the share of compliers in each relative period after the policy shock represents the dynamic effect of that policy change on the take-up rate of the program, and the corresponding CLATT can be interpreted as estimating the effect of participation in job training on real earnings among those affected.
3.3   Identification assumptions in staggered DID-IV designs
In this subsection, we establish the identification assumptions in staggered DID-IV designs. These assumptions are the natural generalization of Assumptions 1-6 in $2\times 2$ DID-IV designs.
Assumption 10 (Exclusion restriction in multiple time periods).
$\forall z\in\mathcal{S}(Z),\forall d\in\mathcal{S}(D),\forall t\in\{1,\dots,T\},Y_{i,t}(d,z)=Y_{i,t}(d)\hskip 8.53581pta.s.$
Assumption 10 extends the exclusion restriction in two time periods (Assumption 2) to multiple period settings. This assumption requires that the instrument path does not directly affect the potential outcome for all time periods and its effects are only through treatment.
Given Assumption 9 and Assumption 10, we can write the potential outcome $Y_{i,t}(d,z)$ as $Y_{i,t}(d_{t})=D_{i,t}Y_{i,t}(1)+(1-D_{i,t})Y_{i,t}(0)$. Following Subsection 2.3, we introduce the potential outcomes in period $t$ if unit $i$ is assigned to the instrument path $z\in\mathcal{S}(Z)$.
$Y_{i,t}(D_{i,t}(z))\equiv D_{i,t}(z)Y_{i,t}(1)+(1-D_{i,t}(z))Y_{i,t}(0).$
Since the initial exposure date $E_{i}$ completely characterizes the instrument path, we can write the potential outcomes for cohort $e$ and cohort $\infty$ as $Y_{i,t}(D_{i,t}^{e})$ and $Y_{i,t}(D_{i,t}^{\infty})$, respectively. The potential outcome $Y_{i,t}(D_{i,t}^{e})$ represents the outcome status in period $t$ if unit $i$ is first exposed to the instrument in period $e$ and the potential outcome $Y_{i,t}(D_{i,t}^{\infty})$ represents the outcome status in period $t$ if unit $i$ is never exposed. We refer to $Y_{i,t}(D_{i,t}^{\infty})$ as the “never exposed outcome”.
Assumption 11 (Monotonicity assumption in multiple time periods).
$Pr(D_{i,e+l}^{e}\geq D_{i,e+l}^{\infty})=1\hskip 5.69054pt\text{or}\hskip 5.69054ptPr(D_{i,e+l}^{e}\leq D_{i,e+l}^{\infty})=1\hskip 5.69054pt\text{for all}\hskip 5.69054pte\in\mathcal{S}(E_{i})\hskip 5.69054pt\text{and for all}\hskip 5.69054ptl\geq 0.$
This assumption requires that the instrument path affects the treatment adoption behavior in a monotone way for all relative periods after the initial exposure date $E_{i}=e$. Recall that we define $D_{i,t}-D_{i,t}^{\infty}$ to be the effect of an instrument on treatment for unit $i$ in period $t$. Assumption 11 requires that the individual exposed effect in the first stage should be non-negative (or non-positive) during the periods after the initial exposure to the instrument for all $i$. This assumption implies that the group variable $G_{i,e,t}\equiv(D_{i,t}^{\infty},D_{i,t}^{e})$ can take three values with non-zero probability for all $e$ and all $t\geq e$. Hereafter, we consider the type of the monotonicity assumption that rules out the existence of the defiers $DF_{e,t}$ for all $t\geq e$ in any cohort $e$.
Assumption 12 (No anticipation in the first stage).
[…]
Assumption 12 requires that potential treatment choices in any $l$ period before the initial exposure to the instrument is equal to the never exposed treatment. This assumption is a natural generalization of Assumption 4 to multiple period settings and restricts the anticipatory behavior before the initial exposure to the instrument.
Assumption 13 (Parallel trends assumption in the treatment in multiple time periods).
[…]
Assumption 13 is a parallel trends assumption in the treatment in multiple periods and cohorts. It requires that the trends of the treatment across cohorts would have followed the same path, on average, if there was no exposure to the instrument. Assumption 13 is analogous to that of Callaway and Sant’Anna (2021) and Sun and Abraham (2021) in DID designs; these studies impose the same type of parallel trends assumption on untreated potential outcomes.
Assumption 14 (Parallel trends assumption in the outcome in multiple time periods).
[…]
Assumption 14 is a parallel trends assumption in the outcome with settings of multiple periods and cohorts. It requires that the expectation of the never exposed outcome across cohorts would have followed the same evolution if the instrument had not been assigned. From the discussion in Subsection 2.4, this assumption can be interpreted as requiring the same expected time gain across cohorts and over time; that is, the effects of time on the outcome through treatment are the same on average across cohorts and over time.
Assumption 15 (Relevance condition in multiple time periods).
$\forall e\in\mathcal{S}(E_{i}),\hskip 5.69054pt\forall l\geq 0,\hskip 5.69054ptE[D_{i,e+l}-D^{\infty}_{i,e+l}|E_{i}=e]>0.$
Assumption 15 is a relevance condition in multiple period settings. With a binary treatment, Assumption 11 and 15 implies that the fraction of the compliers should be non-negative for all $e$ and $l\geq 0$:
$E[D_{i,e+l}-D^{\infty}_{i,e+l}|E_{i}=e]=Pr(CM_{e,l}|E_{i}=e)>0.$
This condition ensures that each $CLATT_{e,l}$ is well defined.
4   Extensions
This section contains extensions to non-binary, ordered treatments and repeated cross sections. It also includes the case that the treatment path is introduced in potential outcomes in $2\times 2$ DID-IV designs. For more details and proofs, see online Appendix Section A.
Non-binary, ordered treatment
Up to now, we have considered only the case of a binary treatment. However the same idea can be applied when treatment takes a finite number of ordered values: $D_{i,t}\in\{0,1,\dots,J\}$. When only two periods exist and treatment is non-binary, our target parameter is the average causal response on the treated (ACRT) defined below.
Definition.
The average causal response on the treated (ACRT) is
$ACRT\equiv\sum_{j=1}^{J}w_{j}\cdot E[Y_{i,1}(j)-Y_{i,1}(j-1)|D_{i,1}((0,1))\geq j>D_{i,1}((0,0)),E_{i}=1].$
where the weight $w_{j}$ is:
$w_{j}=\frac{Pr(D_{i,1}((0,1))\geq j>D_{i,1}((0,0))|E_{i}=1)}{\sum_{j=1}^{J}Pr(D_{i,1}((0,1))\geq j>D_{i,1}((0,0))|E_{i}=1)}.$
The ACRT is a weighted average of the effect of one unit increase in the treatment on the outcome, for those who belong to an exposed group and are induced to increase the treatment in period $t=1$ by instrument. The ACRT is similar to the average causal response (ACR) considered in Angrist and Imbens (1995), with the difference that each weight $w_{j}$ and the associated causal parameter in the ACRT are conditional on $E_{i}=1$.
Online Appendix Section A.1 shows that if we have a non-binary, ordered treatment, the Wald-DID estimand captures the ACRT under the same assumptions in Theorem 1.
With a non-binary, ordered treatment in staggered DID-IV designs, our target parameter is the cohort specific average causal response on the treated (CACRT), a natural generalization of the ACRT. The identifying assumptions are the same as in Subsection 3.3.
Definition.
The cohort specific average causal response on the treated (CACRT) at a given relative period $l$ from the initial adoption of the instrument is
$CACRT_{e,e+l}\equiv\sum_{j=1}^{J}w^{e}_{e+l,j}\cdot E[Y_{i,e+l}(j)-Y_{i,e+l}(j-1)|E_{i}=e,D_{i,e+l}^{e}\geq j>D_{i,e+l}^{\infty}]$
where the weight $w^{e}_{e+l,j}$ is:
$w^{e}_{e+l,j}=\frac{Pr(D_{i,e+l}^{e}\geq j>D_{i,e+l}^{\infty}|E_{i}=e)}{\sum_{j=1}^{J}Pr(D_{i,e+l}^{e}\geq j>D_{i,e+l}^{\infty}|E_{i}=e)}.$
Repeated cross sections
In some applications, researchers have only access to repeated cross section data. [^17] It also includes the case that researchers use the cross section data, and exploit a policy shock across cohorts as an instrument for treatment as in Duflo (2001). Online Appendix Section A.2 presents the identification assumptions in DID-IV designs under repeated cross section settings.
Introducing the treatment path in $2\times 2$ DID-IV designs
In Subsection 2.1, we impose Assumption 1 (no carryover assumption) on potential outcomes $Y_{i,t}(d,z)$, as we consider the non-staggered adoption of the treatment. We can extend our results to the case that we introduce the treatment path in potential outcomes with binary treatment. Online Appendix Section A.3 shows that the similar assumptions to Assumptions 1-6 are sufficient for the Wald-DID to capture the LATET-like parameter.
5   Estimation and inference
In this section, we propose a credible estimation method in staggered DID-IV designs that is robust to treatment effect heterogeneity. First, we propose a simple regression-based method for estimating each CLATT. Following Callaway and Sant’Anna (2021), we then propose a weighting scheme to construct the summary causal parameters from each CLATT. We also discuss the pre-trends tests for checking the plausibility of parallel trends assumptions in DID-IV designs.
Remark 1.
In practice, when researchers implicitly rely on a staggered DID-IV design, they commonly implement this design via two-way fixed instrumental variable (TWFEIV) regressions (e.g., Johnson and Jackson (2019), Lundborg et al. (2014), Black et al. (2005), Akerman et al. (2015), and Bhuller et al. (2013)):
$Y_{i,t}=\phi_{i.}+\lambda_{t.}+\beta_{IV}D_{i,t}+v_{i,t},$
$D_{i,t}=\gamma_{i.}+\zeta_{t.}+\pi Z_{i,t}+\eta_{i,t}.$
In companion paper (Miyaji (2024)), however, we show that in more than two periods, the TWFEIV estimand potentially fails to summarize the treatment effects in the presence of heterogeneous treatment effects. Specifically, we show that the TWFEIV estimand is a weighted average of all possible CLATTs under staggered DID-IV designs, but some weights can be negative if the effect of the instrument on the treatment or the outcome is not stable over time.
5.1   Stacked two stage least squares regression
We propose a regression-based method to consistently estimate our target parameter in staggered DID-IV designs. Our estimation method consists of two steps.
Step 1.
We create the data sets for each $CLATT_{e,e+l}$. Each data set includes the units of time $t=e-1$ and $t=e+l$, who are either in cohort $e\in\{2,\dots,T\}$ or in the set of some unexposed cohorts, $U$ ($e\notin U$). Researchers should carefully specify the choice set $U$ in practice. If a never exposed cohort $E_{i}=\infty$ exists, we can set $U=\{\infty\}$ and estimate $CLATT_{e,e+l}$ for all cohorts ($e\in\{2,\dots,T\}$) and all relative time periods ($l\geq 0$). If no never exposed cohorts exist, we can set $U=\max\{E_{i}\}$, the last exposed cohort, and can estimate $CLATT_{e,e+l}$ for cohort $e\in\{2,\dots,\max\{E_{i}\}-1\}$ and relative time period $l\in\{0,\dots,\max\{E_{i}\}-1-e\}$ because every unit will be exposed to the instrument after time $t=\max\{E_{i}\}$.
Step 2.
In each data set, we run the following IV regression and obtain the IV estimator $\hat{\beta}^{e,l}_{IV}$:
$Y_{i,t}=\beta^{e,l}_{0}+\beta^{e,l}_{i,.}\mathbf{1}\{E_{i}=e\}+\beta^{e,l}_{,.t}\mathbf{1}\{T_{i}=e+l\}+\beta^{e,l}_{IV}D_{i,t}+\epsilon^{e,l}_{i,t},$
where the group indicator $\mathbf{1}\{E_{i}=e\}$ and the post-period indicator $\mathbf{1}\{T_{i}=e+l\}$ are the included instruments and the interaction of the two is the excluded instrument.
Formally, our proposed estimator $\widehat{CLATT}_{e,e+l}\equiv\hat{\beta}^{e,l}_{IV}$ takes the following form:
$\widehat{CLATT}_{e,e+l}\equiv\hat{\beta}^{e,l}_{IV}=\frac{\hat{\alpha}^{e,l}}{\hat{\pi}^{e,l}},$
where $\hat{\alpha}^{e,l}$ and $\hat{\pi}^{e,l}$ are:
$\hat{\alpha}^{e,l}$ $=\frac{E_{N}[(Y_{i,e+l}-Y_{i,e-1})\cdot\mathbf{1}\{E_{i}=e\}]}{E_{N}[\mathbf{1}\{E_{i}=e\}]}-\frac{E_{N}[(Y_{i,e+l}-Y_{i,e-1})\cdot\mathbf{1}\{E_{i}\in U\}]}{E_{N}[\mathbf{1}\{E_{i}\in U\}]}$
$\equiv\hat{\alpha}_{e,l}^{1}-\hat{\alpha}_{e,l}^{2}.$
$\hat{\pi}^{e,l}$ $=\frac{E_{N}[(D_{i,e+l}-D_{i,e-1})\cdot\mathbf{1}\{E_{i}=e\}]}{E_{N}[\mathbf{1}\{E_{i}=e\}]}-\frac{E_{N}[(D_{i,e+l}-D_{i,e-1})\cdot\mathbf{1}\{E_{i}\in U\}]}{E_{N}[\mathbf{1}\{E_{i}\in U\}]}$
$\equiv\hat{\pi}_{e,l}^{1}-\hat{\pi}_{e,l}^{2}.$
Here, $E_{N}[\cdot]$ is the sample analog of the conditional expectation. Note that the IV estimator $\widehat{CLATT}_{e,e+l}$ is the Wald-DID estimator for each $CLATT_{e,e+l}$, where the choice set $U$ is an unexposed group and time $t=e-1$ is a pre-exposed period. From the estimation procedure, we call this a stacked two-stage least squares (STS) estimator. [^18] Our STS estimator is related to the DID estimators in staggered DID designs proposed by Callaway and Sant’Anna (2021) and Sun and Abraham (2021). The DID estimator of the treatment and the outcome in our STS estimator corresponds to that of Sun and Abraham (2021), and coincides with that of Callaway and Sant’Anna (2021) for the case when no covariates exist and never treated units are used as a control group. Their proposed methods avoid the issue of TWFE estimators in staggered DID designs, while our STS estimator avoids the issue of TWFEIV estimators in staggered DID-IV designs.
Theorem 3 below guarantees the validity of our STS estimator.
Theorem 3.
Suppose Assumptions 8-14 hold. Then, the STS estimator $\widehat{CLATT}_{e,e+l}$ is consistent and asymptotically normal.
[…]
Proof.
See Appendix. ∎
From Theorem 3, we can also construct the standard error of the STS estimator, using the sample analogue of the asymptotic variance $V(\psi_{i,e,l})$.
Remark 2.
Our estimation procedure is the same if we have a non-binary, ordered treatment or use repeated cross section data. Online Appendix Section B presents the influence function of the STS estimator in repeated cross section settings.
5.2   Weighting scheme
In this subsection, we explain how one can construct the summary causal parameters from each CLATT in staggered DID-IV designs, based on the weighting scheme proposed by Callaway and Sant’Anna (2021).
We consider the following weighting scheme as in Callaway and Sant’Anna (2021):
$\theta^{IV}=\sum_{e}\sum_{t=1}^{T}w(e,t)\cdot CLATT_{e,t},$
where $w(e,t)$ are some reasonable weighting functions assigned to each $CLATT_{e,t}$.
To propose the weighting functions for a variety of summary causal parameters in staggered DID-IV designs, we define the average effect of the instrument on the treatment at a given relative period $l$ from the initial exposure to the instrument in cohort $e$, called the cohort specific average exposed effect on the treated in the first stage ($CAET_{e,e+l}^{1}$).
Definition.
The cohort specific average exposed effect on the treated in the first stage ($CAET_{e,e+l}^{1}$) at a given relative period $l$ from the initial adoption of the instrument is
$CAET_{e,e+l}^{1}=E[D_{i,e+l}-D_{i,e+l}^{\infty}|E_{i}=e].$
If treatment is binary, each $CAET_{e,l}^{1}$ is equal to the share of the compliers in cohort $e$ in period $e+l$:
$CAET^{1}_{e,e+l}$ $=E[D_{i,e+l}^{e}-D_{i,e+l}^{\infty}|E_{i}=e]$
$=Pr(CM_{e,e+l}|E_{i}=e).$
In staggered DID designs, Callaway and Sant’Anna (2021) propose various aggregated measures along with different dimensions of treatment effect heterogeneity. We can employ their framework directly, but in staggered DID-IV designs, we should more carefully specify the weighting functions assigned to each summary measure.
For instance, to aggregate dynamic treatment effects in cohort $e$ over time in staggered DID designs, Callaway and Sant’Anna (2021) consider the following summary measure:
$\theta_{sel}(\tilde{e})=\frac{1}{T-\tilde{e}+1}\sum_{t=\tilde{e}}^{T}CATT(\tilde{e},t),$
where $CATT(\tilde{e},t)$ is the cohort specific average treatment effect on the treated at time $t$ in cohort $\tilde{e}$ (see, e.g., Callaway and Sant’Anna (2021), Sun and Abraham (2021)).
Table 5: Weights for a variety of summary causal parameters
Target Parameter $w(e,t)$
$\theta_{es(l)}^{IV}$ $\mathbf{1}\{e+l\leq T\}\mathbf{1}\{t=e+l\}P(E=e|E+l\leq T)\displaystyle\frac{CAET_{e,e+l}^{1}}{\sum_{e\in\mathcal{S}(E)}CAET_{e,e+l}^{1}}$
$\theta_{es(l,l^{\prime})}^{bal,IV}$ $\mathbf{1}\{e+l^{\prime}\leq T\}\mathbf{1}\{t=e+l\}P(E=e|E+l^{\prime}\leq T)\displaystyle\frac{CAET_{e,e+l}^{1}}{\sum_{e\in\mathcal{S}(E)}CAET_{e,e+l}^{1}}$
$\theta_{sel(\tilde{e})}^{IV}$ $\mathbf{1}\{t\geq e\}\mathbf{1}\{e=\tilde{e}\}\displaystyle\frac{CAET_{\tilde{e},t}^{1}}{\sum_{t=\tilde{e}}^{T}CAET_{\tilde{e},t}^{1}}$
$\theta_{c(\tilde{t})}^{IV}$ $\mathbf{1}\{t\geq e\}\mathbf{1}\{t=\tilde{t}\}P(E=e|E\leq t)\displaystyle\frac{CAET_{e,t}^{1}}{\sum_{e\in\mathcal{S}(E)}CAET_{e,t}^{1}}$
$\theta_{c(\tilde{t})}^{cumm,IV}$ $\mathbf{1}\{t\geq e\}\mathbf{1}\{t\leq\tilde{t}\}P(E=e|E\leq t)\displaystyle\frac{CAET_{e,t}^{1}}{\sum_{e\in\mathcal{S}(E)}CAET_{e,t}^{1}}$
$\theta_{W}^{o,IV}$ $\mathbf{1}\{t\geq e\}P(E=e|E\leq T)/\sum_{e\in\mathcal{S}(E)}\sum_{t=1}^{T}\mathbf{1}\{t\geq e\}P(E=e|E\leq T)$
$\theta_{sel}^{o,IV}$ $\mathbf{1}\{t\geq e\}P(E=e|E\leq T)\displaystyle\frac{CAET_{e,t}^{1}}{\sum_{t=e}^{T}CAET_{e,t}^{1}}$
Notes: This table represents the specific expressions for the weights on each $CLATT(e,t)$ or each $CACRT(e,t)$ in a variety of summary causal parameters. Each target parameter without the superscript IV is defined in Callaway and Sant’Anna (2021). The superscript IV is added to associate each parameter with staggered DID-IV designs.
In staggered DID-IV designs, we propose the following summary measure $\theta_{sel}(\tilde{e})^{IV}$ that corresponds with $\theta_{sel}(\tilde{e})$:
$\theta_{sel}(\tilde{e})^{IV}=\sum_{t=\tilde{e}}^{T}\frac{CAET^{1}_{\tilde{e},t}}{\sum_{t=\tilde{e}}^{T}CAET^{1}_{\tilde{e},t}}CLATT_{\tilde{e},t}.$
This parameter summarizes each $CLATT_{\tilde{e},t}$ in cohort $\tilde{e}$ across all post-exposed periods and the weight assigned to each $CLATT_{\tilde{e},t}$ reflects the relative share of the compliers $CM_{\tilde{e},t}$ in period $t$ during the periods after the initial exposure to the instrument in cohort $\tilde{e}$. This weighting scheme would be reasonable in that it is designed to be larger in the period when the proportion of the compliers is relatively higher in cohort $\tilde{e}$.
By similar arguments, we can specify the weighting functions for various summary measures, which correspond with those in Callaway and Sant’Anna (2021). Table 5 summarizes the specific expressions for the weights assigned to each $CLATT_{\tilde{e},t}$ in each summary causal parameter. Note that our proposed weighting functions are the same under non-binary, ordered treatment settings, in which our target parameter is the $CACRT_{\tilde{e},t}$.
Estimation and inference
We can construct the consistent estimator for the summary causal parameter $\theta^{IV}$ as follows.
$\hat{\theta}^{IV}=\sum_{e}\sum_{t=1}^{T}\hat{w}(e,t)\cdot\widehat{CLATT}_{e,t},$
where $\hat{w}(e,t)$ is the sample analog of each ${w}(e,t)$ and $\widehat{CLATT}_{e,t}$ is the consistent estimator for each $CLATT_{e,t}$, obtained from the STS regression in Subsection 5.1. Each $\hat{w}(e,t)$ is the regular asymptotically linear estimator:
$\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\zeta_{i,e,t}^{w}+o_{p}(1),$
where $\zeta_{i,e,t}^{w}$ represents the influence function, and satisfies $E[\zeta_{i,e,t}^{w}]=0$ and $E[\zeta_{i,e,t}^{w}{\zeta_{i,e,t}^{w}}^{\top}]<\infty$.
Corollary 1 below represents the asymptotic distribution of the plug-in estimator $\hat{\theta}_{IV}$ and ensures its validity.
Corollary 1.
If the assumptions of Theorem 3 hold,
[…]
where the influence function $l_{i}^{\theta_{IV}}$ takes the following form:
$l_{i}^{\theta_{IV}}=\sum_{e}\sum_{t=1}^{T}\left(w(e,t)\cdot\psi_{i,e,t}+\zeta_{i,e,t}^{w}\cdot{CLATT}_{e,t}\right).$
5.3   Pre-trends test
In most DID applications, researchers assess the plausibility of the parallel trends assumption by testing for pre-treatment differences in the outcome between treatment and control groups (pre-trends test). In this section, we describe the procedure of the pre-trends test in DID-IV designs to check the validity of the parallel trends assumption in the treatment and the outcome.
Suppose that data are available for period $t=-1$. Then, one can assess the plausibility of Assumption 5 between period $t=-1$ and $t=0$ by testing the following null hypothesis
$E[D_{i,0}-D_{i,-1}|E_{i}=0]=E[D_{i,0}-D_{i,-1}|E_{i}=1]$
$\iff$ $E[D_{i,0}((0,0))-D_{i,-1}((0,0))|E_{i}=0]=E[D_{i,0}((0,0))-D_{i,-1}((0,0))|E_{i}=1].$ (1)
Similarly, one can assess the plausibility of Assumption 6 between period $t=-1$ and $t=0$ by testing the following null hypothesis
$E[Y_{i,0}-Y_{i,-1}|E_{i}=0]=E[Y_{i,0}-Y_{i,-1}|E_{i}=1]$
$\iff$ $E[Y_{i,0}(D_{i,0}((0,0)))-Y_{i,-1}(D_{i,-1}((0,0)))|E_{i}=0]$
$=$ $E[Y_{i,0}(D_{i,0}((0,0)))-Y_{i,-1}(D_{i,-1}((0,0)))|E_{i}=1].$ (2)
These tests can be generalized to multiple pre-exposed periods settings by using the pre-trends tests recently developed in the context of DID designs (Callaway and Sant’Anna (2021), Sun and Abraham (2021), Borusyak et al. (2024)); that is, one can apply these tools to the first stage and the reduced form ($D_{i,t}$ or $Y_{i,t}$ is outcome, $Z_{i,t}$ is treatment) respectively to confirm the plausibility of Assumption 13 and 14 in pre-exposed periods.
6   Application
We illustrate the empirical relevance of our findings in the setting of Oreopoulos (2006), estimating returns to schooling in the United Kingdom. We first assess the identifying assumptions in staggered DID-IV designs implicitly imposed by Oreopoulos (2006). We then estimate the TWFEIV regression in the author’s setting. Finally, we estimate the target parameter and its summary measure by employing our proposed method and weighting scheme.
6.1   Setting
Oreopoulos (2006) estimates returns to schooling using a major education reform in the UK that increased the years of compulsory schooling from 14 to 15. Specifically, Oreopoulos (2006) exploits variation resulting from the different timing of implementation of school reforms between Britain (England, Scotland, and Wales) and Northern Ireland as an instrument for education attainment: the school-leaving age increased in Britain in 1947, but was not implemented until 1957 in Northern Ireland. [^19] In the first part of his analysis, Oreopoulos (2006) adopts regression discontinuity designs (RDD) and analyzes the data sets in Britain and Northern Ireland separately. Due to the imprecision of their standard errors, Oreopoulos (2006) then moves to “a difference-in-differences and instrumental-variables analysis by combining the two sets of U.K. data”. The data are a sample of individuals in Britain and Northern Ireland, who were aged $14$ between $1936$ and $1965$, constructed from combining the series of U.K. General Household Surveys between $1984$ and $2006$; see Oreopoulos (2006), Oreopoulos (2008) for details.
Oreopoulos (2006) (more precisely, Oreopoulos (2008)) runs the following two-way fixed effects instrumental variable regression with the education reform as an excluded instrument for education attainment.
$Y_{i,t}=\mu_{n.}+\delta_{.t}+\beta_{IV}D_{i,t}+\epsilon_{i,t},$ (3)
$D_{i,t}=\gamma_{n.}+\zeta_{.t}+\pi Z_{i,t}+\eta_{i,t}.$ (4)
Here, a cohort (a year when aged $14$) plays a role of time as it determines the exposure to the policy change. The dependent variables $Y_{i,t}$ and $D_{i,t}$ are log annual earnings and education attainment for unit $i$ and cohort $t$, respectively. Both the first stage and the reduced form regressions include a birth cohort fixed effect and a North Ireland fixed effect. The binary instrument $Z_{i,t}\in\{0,1\}$ takes one if unit $i$ in cohort $t$ is exposed to the policy change.
The staggered introduction of the school reform can be viewed as a natural experiment, but is not randomized across regions in reality; Oreopoulos (2006) notes that the reform was implemented with political support, taking into account costs and benefit. This indicates that Oreopoulos (2006) implicitly relies on a staggered DID-IV identification strategy instead of exploiting the random variation of the policy change. Indeed, Oreopoulos (2006) presents the corresponding plots of British and Northern Irish average education attainment and average log earnings by cohort to illustrate the evolution of these variables before and after the policy shock.
6.2   Assessing the identifying assumptions in staggered DID-IV designs
We first discuss the validity of the staggered DID-IV identification strategy in Oreopoulos (2006). In the author’s setting, our target parameter is the cohort specific average causal response on the treated (CACRT), as education attainment is a non-binary, ordered treatment.
Exclusion restriction.
It would be plausible, given that the policy reform did not affect log annual earnings other than by increasing education attainment. This assumption may be violated for instance if the reform affected both the quality and quantity of education.
Monotonicity assumption.
It would be automatically satisfied in the author’s setting: the policy change (instrument) increased the minimum schooling-leaving age from $14$ to $15$, ensuring that there are no defiers during the periods after the policy shock.
No anticipation in the first stage.
It would be plausible that there is no anticipation, if the treatment adoption behavior is the same as the one in the absence of the policy change before its implementation in England. This assumption may be violated if units have private knowledge about the probability of extended education and manipulate their education attainment before the policy shock.
Next, we assess the validity of the parallel trends assumptions in the treatment and the outcome, using the interacted two-way fixed effects regressions proposed by Sun and Abraham (2021) in the first stage and the reduced form, respectively. The results are shown in Figure 1.
Figure 1: The effect of the instrument in the first stage and reduced form in the setting of Oreopoulos (2006). Notes: This figure presents the results of the effect of the school reform on education attainment (Panel (a)) and on log annual earnings (Panel (b)) under the staggered DID-IV identification strategy. The unexposed group $U$ is a last-exposed cohort, North Ireland and the reference period is $t=1946$. The blue lines represent the estimates with pointwise $95\%$ confidence intervals for pre-exposed periods in both panels. These should be equal to zero under the null hypothesis that the parallel trends assumptions in the treatment and the outcome hold. The red lines represent the estimates with pointwise $95\%$ confidence intervals for post-exposed periods in both panels.
Parallel trends assumption in the treatment.
It requires that the expectation of education attainment would have followed the same path between England and North Ireland across cohorts in the absence of the school reform. Panel (a) in Figure 1 plots the results of the interacted two-way fixed effects regression in the first stage along with a $95\%$ pointwise confidence interval. The pre-exposed estimates are not significantly different from zero and indicate the validity of Assumption 13.
Parallel trends assumption in the outcome.
It requires that the expectation of log annual earnings would have followed the same evolution between England and North Ireland across cohorts in the absence of the school reform. Panel (b) in Figure 1 plots the results of the interacted two-way fixed effects regression in the reduced form along with a $95\%$ pointwise confidence interval. The pre-exposed estimates seems consistent with Assumption 14: though an upward pre-trends exists, all the estimates before the initial exposure to the policy change are not significantly different from zero.
Figure 1 also sheds light on the dynamic effects of the school reform on education attainment and log annual earnings in post-exposed periods. In Panel (a), the estimated increase in years of schooling after the reform ranges from $0.56$ to $0.87$, and all estimates are statistically significant. In Panel (b), the estimated increase in log annual earnings after the reform varies from $8\%$ to $25\%$, and $5$ out of $10$ estimates are statistically significant. [^20] Note that the post-exposed estimates in Panel (b) do not capture each CACRT in post-exposed periods: each estimate in the reduced form is not scaled by the corresponding estimate in the first stage.
6.3   Illustrating our estimation method
We start by estimating two-way fixed effects instrumental variable regression in the author’s setting. To clearly illustrate the pitfalls of TWFEIV regression, in our estimation, we slightly modify the author’s specification. Oreopoulos (2006) (more precisely Oreopoulos (2008)) includes some covariates (survey year, sex, and a quartic in age) and runs the weighted regression in the main specification (see Oreopoulos (2006), Oreopoulos (2008) for details), whereas we exclude such covariates and do not apply their weights to our regression.
The result is shown in Table 6. The TWFEIV estimate is $-0.009$ and not significantly different from 00. This indicates that, on the whole, the returns to schooling in the UK is nearly zero. However, this may be a misleading conclusion. We cannot interpret that the TWFEIV estimand captures the properly weighted average of all possible CACRTs if the effect of the school reform on education attainment or log annual earnings is not stable across cohorts (Miyaji (2024)).
In online Appendix Section C, we quantify the bias terms of the TWFEIV estimand by using Lemma 7 in Miyaji (2024), and show that the TWFEIV estimand is negatively biased in Oreopoulos (2006). Specifically, we show that all the bias terms are positive, while all the assigned weights are negative, which yields the downward bias for the TWFEIV estimand. [^21] When we run the TWFEIV regression in the author’s setting, we treat the set of already exposed units in North Ireland as controls between $1957$ and $1965$, as both regions are already exposed to the policy shock during these periods. This procedure performed by the TWFEIV regression is called the “bad comparisons” in the recent DID literature (c.f. Goodman-Bacon (2021)), and yields the bias in the TWFEIV estimand from the properly weighted average of all possible CACRTs.
Figure 2: The cohort specific average causal response on the treated in each relative period in the setting of Oreopoulos (2006). Notes: This figure shows the results for returns to schooling under the staggered DID-IV identification strategy. The unexposed group $U$ is North Ireland and the reference period is period $t=1946$. The red lines represent the stacked two stage least squares estimates with pointwise $95\%$ confidence intervals for post-exposed periods.
We now employ the proposed method to estimate each CACRT. First, we create data sets by cohort $t$ ($t=1947,\dots,1956$). Each data set only contains units of cohort $t$ and cohort $t=1946$ in England and North Ireland. We define North Ireland as an unexposed group $U$ and discard the units aged at $14$ between $1957$ and $1965$ to avoid the bad comparisons. We then run TSLS regression in each data set. The standard error is calculated by using the influence function shown in equation (B) in online Appendix Section LABEL:ApeC.1.
Table 6: Returns to schooling in Oreopoulos (2006)
Estimate Standard Error 95% CI
TSLS with fixed effects -0.009 0.039 [-0.085, 0.068]
$\theta_{sel}^{IV}(e)$ 0.240 0.098 [0.047, 0.433]
Notes: Sample size $82790$ observations.
Figure 2 plots the point estimates and the corresponding $95\%$ confidence intervals in each relative period after the school reform. The estimates for each $CACRT_{e,t}$ range from $14\%$ to $38\%$ with wide confidence intervals, and $4$ out of $10$ estimates are statistically significant.
Finally, we estimate the summary causal measure by aggregating each $CACRT_{e,t}$. Specifically, we estimate the weighted average of each $CACRT_{e,t}$ during post-exposed periods in England ($e=1947$).
$\theta_{sel}^{IV}(e)=\sum_{t=1947}^{1956}\frac{CAET^{1}_{e,t}}{\sum_{t=1947}^{1956}CAET^{1}_{e,t}}CACRT_{e,t}.$
Here, each weight assigned to $CACRT_{e,t}$ represents the relative amount of the effect of the policy reform on education attainment during post-exposed periods in England.
The result is shown in Table 6. The estimate is $0.24$ and it is significantly different from zero. The estimated returns to schooling are substantial, perhaps because each $CACRT_{e,t}$ captures the returns to schooling among the compliers: such units may belong to relatively low-skilled labor or low-income family that potentially have much to gain from school reform.
The estimates obtained from our proposed method and weighting scheme are significantly different from the TWFEIV estimate: our STS estimates and its summary measure are all positive, whereas the TWFEIV estimate is strictly negative. Overall, our results indicate that the economic returns of education are substantial in the UK, and the estimation method matters in staggered DID-IV designs in practice.
7   Conclusion
In this paper, we formalize an instrumented difference-in-differences (DID-IV). First, we consider a simple setting with two periods and two groups. In this setting, our DID-IV design mainly comprises a monotonicity assumption, and parallel trends assumptions in the treatment and the outcome between the two groups. We show that in $2\times 2$ DID-IV designs, the Wald-DID estimand captures the local average treatment effect on the treated (LATET). After establishing $2\times 2$ DID-IV designs, we compare DID-IV to Fuzzy DID designs considered in de Chaisemartin and D’Haultfœuille (2018), and point out the issues inherent in Fuzzy DID designs.
Next, we consider the DID-IV design in more than two periods with units being exposed to the instrument at different times. We call this a staggered DID-IV design, and formalize the target parameter and identifying assumptions. Specifically, in this design, our target parameter is the cohort specific local average treatment effects on the treated (CLATT). The identifying assumptions are the natural generalization of those in $2\times 2$ DID-IV designs.
We also provide the estimation method in staggered DID-IV designs that does not require strong restrictions on treatment effect heterogeneity. Our estimation method carefully chooses the comparison groups and does not suffer from the bias arising from the time-varying exposed effects. We also propose the weighting scheme in staggered DID-IV designs, and explain how one can conduct pre-trends tests to assess the validity of the parallel trends assumptions in DID-IV designs.
Finally, we illustrate the empirical relevance of our findings with the setting of Oreopoulos (2006), who estimates returns to schooling in the UK, exploiting the timing variation of the introduction of school reforms between British and North Ireland. In this application, the TWFEV regression, the conventional approach to implement a staggered DID-IV design, yields a negative estimate. By contrast, our STS regression and weighting scheme indicate the substantial gain from schooling. This empirical application illustrates that the estimation method matters in staggered DID-IV designs in practice.
Overall, this paper provides a new econometric framework for estimating the causal effects when the treatment adoption is potentially endogenous over time, but researchers can exploit variation in policy adoption timing as an instrument for treatment. To avoid the issue of using the TWFEIV estimator, we also provide a reliable estimation method that is free from strong restrictions on treatment effect heterogeneity. Further developing alternative estimation methods and diagnostic tools will be a promising area for future research, facilitating the credibility of DID-IV design in practice.
Appendix
Proof of Theorem 1.
Proof.
First we consider the numerator of the Wald-DID estimand. For the numerator, we have
$E[Y_{1}-Y_{0}|E=1]-E[Y_{1}-Y_{0}|E=0]$
$=$ $E[Y_{1}(D_{1}((0,1)))-Y_{0}(D_{0}((0,1)))|E=1]-E[Y_{1}(D_{1}((0,0)))-Y_{0}(D_{0}((0,0)))|E=0]$
$=$ $E[Y_{1}(D_{1}((0,1)))-Y_{1}(D_{1}((0,0)))|E=1]$
+$+\{E[Y_{1}(D_{1}((0,0)))-Y_{0}(D_{0}((0,1)))|E=1]-E[Y_{1}(D_{1}((0,0)))-Y_{0}(D_{0}((0,0)))|E=0]\}$
$=$ $E[Y_{1}(D_{1}((0,1)))-Y_{1}(D_{1}((0,0)))|E=1]$
+$+\{E[Y_{1}(D_{1}((0,0)))-Y_{0}(D_{0}((0,0)))|E=1]-E[Y_{1}(D_{1}((0,0)))-Y_{0}(D_{0}((0,0)))|E=0]\}$
$=$ $E[(D_{1}((0,1))-D_{1}((0,0)))\cdot(Y_{1}(1)-Y_{1}(0))|E=1]$
$=$ $LATET\cdot Pr(CM^{Z}|E=1).$ (5)
The first equality follows from Assumption 1 and Assumption 2. The second equality follows from the simple manipulation. The third equality follows from Assumption 4. The fourth equality follows from Assumption 6 and simple calculation. The fifth equality follows from the Law of Iterated Expectations and Assumption 3.
For the denominator, we have
$E[D_{1}-D_{0}|E=1]-E[D_{1}-D_{0}|E=0]$
$=$ $E[D_{1}((0,1))-D_{0}((0,1))|E=1]-E[D_{1}((0,0))-D_{0}((0,0))|E=0]\hskip 8.53581pt$
$=$ $E[D_{1}((0,1))-D_{1}((0,0))|E=1]$
+$+\{E[D_{1}((0,0))-D_{0}((0,1))|E=1]-E[D_{1}((0,0))-D_{0}((0,0))|E=0]\}$
$=$ $E[D_{1}((0,1))-D_{1}((0,0))|E=1]$
+$+\{E[D_{1}((0,0))-D_{0}((0,0))|E=1]-E[D_{1}((0,0))-D_{0}((0,0))|E=0]\}$
$=$ $Pr(CM^{Z}|E=1).$ (6)
where the second equality follows from the simple manipulation. The third equality follows from Assumption 4 and the forth equality follows from Assumption 3 and Assumption 5.
Combining the result (5) with (6), we obtain the desirable result. ∎
Proof of Theorem 2.
Proof.
First, we prove the consistency of our SLS estimator. We assume that $\hat{\alpha}^{e,l}$ and $\hat{\pi}^{e,l}$ are well defined. More precisely, we assume that cohort $e$ is non-empty; that is, $\sum_{i=1}^{n}\mathbf{1}\{E_{i}=e\}>0$ and there exist some sets of the unexposed cohort $U\subseteq\{u:e-1<u\leq T\}$ that are non-empty; that is, $\sum_{i=1}^{n}\mathbf{1}\{E_{i}\in U\}>0$. This assumption guarantees that each Wald-DID estimator $\widehat{CLATT}_{e,e+l}$ is also well defined.
First we derive the probability limit of $\widehat{CLATT}_{e,e+l}$. By the Law of Large Numbers and continuous mapping theorem, we have
[…]
Next, we show that the right hand side of (7) is equal to $CLATT_{e,e+l}$. For the numerator, we have
$E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=e]-E[Y_{i,e+l}-Y_{i,e-1}|E_{i}\in U]$
$=$ $E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=e]-\sum_{e\in U}E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=e]Pr(E_{i}=e|E_{i}\in U)$
$=$ $E[Y_{i,e+l}(D_{i,e+l}^{e})-Y_{i,e-1}(D_{i,e-1}^{\infty})|E_{i}=e]$ (8)
$-\sum_{e\in U}E[Y_{i,e+l}(D_{i,e+l}^{\infty})-Y_{i,e-1}(D_{i,e-1}^{\infty})|E_{i}\in U]Pr(E_{i}=e|E_{i}\in U)$
$=$ $E[Y_{i,e+l}(D_{i,e+l}^{e})-Y_{i,e+l}(D_{i,e+l}^{\infty})|E_{i}=e]+E[Y_{i,e+l}(D_{i,e+l}^{\infty})-Y_{i,e-1}(D_{i,e-1}^{\infty})|E_{i}=e]$
$-\sum_{e\in U}E[Y_{i,e+l}(D_{i,e+l}^{\infty})-Y_{i,e-1}(D_{i,e-1}^{\infty})|E_{i}\in U]Pr(E_{i}=e|E_{i}\in U)$
$=$ $E[Y_{i,e+l}(D_{i,e+l}^{e})-Y_{i,e+l}(D_{i,e+l}^{\infty})|E_{i}=e]$
$+E[Y_{i,e+l}(D_{i,e+l}^{\infty})-Y_{i,e-1}(D_{i,e-1}^{\infty})]-E[Y_{i,e+l}(D_{i,e+l}^{\infty})-Y_{i,e-1}(D_{i,e-1}^{\infty})]$
$=$ $E[Y_{i,e+l}(1)-Y_{i,e+l}(0)|E_{i}=e,CM_{e,e+l}]Pr(CM_{e,e+l}|E_{i}=e).$ (9)
The third equality follows from Assumption 9, 10 and 12. The fourth equality follows from Assumption 14 and the fifth follows from Assumption 11. For the denominator, by similar calculations,
$E[D_{i,e+l}-D_{i,e-1}|E_{i}=e]-E[D_{i,e+l}-D_{i,e-1}|E_{i}\in U]$
$=E[D_{i,e+l}^{e}-D_{i,e+l}^{\infty}|E_{i}=e]+E[D_{i,e+l}^{\infty}-D_{i,e-1}^{\infty}|E_{i}=e]$ (10)
$-\sum_{e\in U}E[D_{i,e+l}^{\infty}-D_{i,e-1}^{\infty}|E_{i}=e]Pr(E_{i}=e|E_{i}\in U)$
$=Pr(CM_{e,e+l}|E_{i}=e),$ (11)
where the final equality follows from Assumption 11 and Assumption 13. Combining (9) with (11), we obtain the desirable result.
Next, we prove that our STS estimator is asymptotically normal, deriving its influence function. We define $\theta_{e,l}$ to be
$\theta_{e,l}$ $\equiv\frac{E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=e]-E[Y_{i,e+l}-Y_{i,e-1}|E_{i}\in U]}{E[D_{i,e+l}-D_{i,e-1}|E_{i}=e]-E[D_{i,e+l}-D_{i,e-1}|E_{i}\in U]}(=CLATT_{e,e+l})$
$\equiv\frac{\alpha_{e,l}^{1}-\alpha_{e,l}^{2}}{\pi_{e,l}^{1}-\pi_{e,l}^{2}}.$
The following fact, also found in de Chaisemartin and D’Haultfœuille (2018), is repeatedly used in the derivation of the influence function.
Fact 1.
If
$\sqrt{n}(\hat{A}-A)=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}a_{i}+o_{p}(1),\hskip 8.53581pt\sqrt{n}(\hat{B}-B)=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}b_{i}+o_{p}(1),$
we have
$\sqrt{n}\left(\frac{\hat{A}}{\hat{B}}-\frac{A}{B}\right)=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\frac{a_{i}-(A/B)b_{i}}{B}+o_{p}(1).$
Then, by some calculations, we obtain
$\sqrt{n}(\widehat{CLATT}_{e,e+l}-CLATT_{e,e+l})$ $=\frac{1}{\sqrt{n}}\sum_{i}^{n}\psi_{i,e,l}+o_{p}(1).$
The influence function $\psi_{i,e,l}$ is:
$\psi_{i,e,l}$
$=\frac{1}{\pi_{e,l}^{1}-\pi_{e,l}^{2}}\Bigg{(}\frac{\mathbf{1}\{E_{i}=e\}\left[\delta_{i,e,l}^{p}-E[\delta_{i,e,l}^{p}|E_{i}=e]\right]}{E[\mathbf{1}\{E_{i}=e\}]}-\frac{\mathbf{1}\{E_{i}=U\}\left[\delta_{i,e,l}^{p}-(E[\delta_{i,e,l}^{p}|E_{i}=U])\right]}{E[\mathbf{1}\{E_{i}=U\}]}\Bigg{)},$ (12)
where we define $\delta_{i,e,l}^{p}=(Y_{i,e+l}-Y_{i,e-1})-\theta_{e,l}\cdot(D_{i,e+l}-D_{i,e-1})$.
Therefore, the asymptotic variance $V(\psi_{i,e,l})$ is
$V(\psi_{i,e,l})$ $=\frac{1}{(\pi_{e,l}^{1}-\pi_{e,l}^{2})^{2}}\Bigg{(}\frac{E\big{[}\mathbf{1}\{E_{i}=e\}\left[\delta_{i,e,l}^{p}-E[\delta_{i,e,l}^{p}|E_{i}=e]\right]^{2}\big{]}}{E[\mathbf{1}\{E_{i}=e\}]^{2}}$
$+\frac{E\big{[}\mathbf{1}\{E_{i}=U\}\left[\delta_{i,e,l}^{p}-(E[\delta_{i,e,l}^{p}|E_{i}=U]\right]^{2}\big{]}}{E[\mathbf{1}\{E_{i}=U\}]^{2}}\Bigg{)}.$
Appendix A Extensions
This section considers the non-binary, ordered treatment and repeated cross section settings in DID-IV designs. It also includes the case where we introduce the treatment path in $2\times 2$ DID-IV designs.
A.1   Non-binary, ordered treatment
In this subsection, we present the identification results in DID-IV designs with non-binary, ordered treatment. Theorem 4 below shows that if we have a non-binary, ordered treatment, the Wald-DID estimand is equal to the ACRT under Assumptions 1-7, which replace the binary treatment with a non-binary one.
Theorem 4.
If Assumptions 1-7 hold, the Wald-DID estimand $W_{DID}$ is equal to the ACRT.
Proof.
Let $\lambda_{Z,j}^{t}=\mathbf{1}(D_{t}((0,z))\geq j)$ be the indicator function for $t\in\{0,1\},z\in\{0,1\}$ and $j\in\{0,\dots,J+1\}$. Note that $\lambda_{Z,0}^{t}=1$ and $\lambda_{Z,J+1}^{t}=0$ hold for all $t$ and $z$ by construction. Here, we can rewrite the observed outcome $Y_{t}$ as follows:
$Y_{t}$ $=Z_{1}\cdot Y_{t}(D_{t}((0,1)))+(1-Z_{1})\cdot Y_{t}(D_{t}((0,0)))$
=$=\left\{Z_{1}\cdot\sum_{j=0}^{J}Y_{t}(j)\cdot(\lambda_{1,j}^{t}-\lambda_{1,j+1}^{t})\right\}+\left\{(1-Z_{1})\cdot\sum_{j=0}^{J}Y_{t}(j)\cdot(\lambda_{0,j}^{t}-\lambda_{0,j+1}^{t})\right\}.$
where we use Assumption 1 and 2.
For the numerator of the Wald DID estimand, we have
$E[Y_{1}-Y_{0}|E=1]-E[Y_{1}-Y_{0}|E=0]$
$=$ $E[Y_{1}(D_{1}((0,1)))-Y_{1}(D_{1}((0,0)))|E=1].$
This follows from Assumption 4 and Assumption 6.
In terms of the $\lambda_{Z,j}^{t}$, we write the previous expression as:
$E[Y_{1}(D_{1}((0,1)))-Y_{1}(D_{1}((0,0)))|E=1]$
$=E\left[\sum_{j=0}^{J}Y_{1}(j)\cdot[\lambda_{1,j}^{1}-\lambda_{1,j+1}^{1}-\lambda_{0,j}^{1}+\lambda_{0,j+1}^{1}]|E=1\right]$
$=E\left[\sum_{j=1}^{J}(Y_{1}(j)-Y_{1}(j-1))\cdot[\lambda_{1,j}^{1}-\lambda_{0,j}^{1}]+Y_{1}(0)\cdot(\lambda_{1,0}^{1}-\lambda_{0,0}^{1})|E=1\right]$
$=E\left[\sum_{j=1}^{J}(Y_{1}(j)-Y_{1}(j-1))\cdot[\lambda_{1,j}^{1}-\lambda_{0,j}^{1}]|E=1\right].$
where the second equality follows from $\lambda_{1,0}^{1}-\lambda_{0,0}^{1}=1$.
Note that $\lambda_{1,j}^{1}\geq\lambda_{0,j}^{1}$ from Assumption 3 and that $\lambda_{1,j}^{1}$ and $\lambda_{0,j}^{1}$ can take only two values, zero or one. Therefore, we obtain:
$E\left[\sum_{j=1}^{J}(Y_{1}(j)-Y_{1}(j-1))\cdot[\lambda_{1,j}^{1}-\lambda_{0,j}^{1}]|E=1\right]$
$=\sum_{j=1}^{J}E[(Y_{1}(j)-Y_{1}(j-1)|\lambda_{1,j}^{1}-\lambda_{0,j}^{1}=1,E=1]Pr(\lambda_{1,j}^{1}-\lambda_{0,j}^{1}=1|E=1)$
$=\sum_{j=1}^{J}E[(Y_{1}(j)-Y_{1}(j-1)|D_{1}((0,1))\geq j>D_{1}((0,0)),E=1]$
×$\times Pr(D_{1}((0,1))\geq j>D_{1}((0,0))|E=1).$ (13)
For the denominator of the Wald-DID estimand, by similar argument, we obtain:
$E[D_{1}-D_{0}|E=1]-E[D_{1}-D_{0}|E=0]$
$=$ $E[D_{1}((0,1))-D_{1}((0,0))|E=1].$
This follows from Assumption 4 and 5.
Then, we have:
$E[D_{1}((0,1))-D_{1}((0,0))|E=1]$
$=E\left[\sum_{j=0}^{J}j\cdot[\lambda_{1,j}^{1}-\lambda_{1,j+1}^{1}-\lambda_{0,j}^{1}+\lambda_{0,j+1}^{1}]|E=1\right]$
$=E\left[\sum_{j=1}^{J}[\lambda_{1,j}^{1}-\lambda_{0,j}^{1}]|E=1\right]$
$=\sum_{j=1}^{J}Pr(D_{1}((0,1))\geq j>D_{1}((0,0))|E=1).$ (14)
Combining the result (14) with (13), we obtain the desired result. ∎
A.2   Repeated cross sections
Here, we present our identification results under repeated cross section settings. First, we consider two groups and two periods settings. Next, we consider multiple period settings with staggered instrument.
A.2.1 Two time periods
Let $Y_{i}$ and $D_{i}$ denote the outcome and the treatment for unit $i$. Let $Z_{i}$ denote the instrument path for unit $i$: $Z_{i}=(0,0)$ if unit $i$ is not exposed to the instrument and $Z_{i}=(0,0)$ if unit $i$ is exposed to the instrument. Let $E_{i}\in\{0,1\}$ denote the group indicator for unit $i$: $E_{i}=1$ if $Z_{i}=(0,1)$ and $E_{i}=0$ if $Z_{i}=(0,0)$. Let $T_{i}\in\{0,1\}$ denote the binary indicator for time. For all $z\in\mathcal{S}(Z)$, let $Y_{i}(0,z),Y_{i}(1,z)$, and $D_{i}(z)$ denote potential outcomes and potential treatment choices for unit $i$. Let $D_{i,t}(z)$ denote potential treatment choices under $T_{i}=t$. We assume that $\{Y_{i},D_{i},Z_{i},E_{i},T_{i}\}_{i=1}^{n}$ are i.i.d.
In two periods and two groups settings, our target parameter is the local average treatment effect on the treated in period $1$:
$LATET$ $\equiv E[Y_{i}(1)-Y_{i}(0)|T_{i}=1,E_{i}=1,D_{i,1}((0,1))>D_{i,1}((0,0))]$
$=E[Y_{i}(1)-Y_{i}(0)|T_{i}=1,E_{i}=1,CM^{Z}].$
We make the following identification assumptions for the Wald-DID estimand to capture the LATET. These assumptions are suitable for repeated cross section settings.
Assumption 14 (Exclusion restriction).
$\forall z\in\mathcal{S}(Z),Y_{i}(0,z)=Y_{i}(0),Y_{i}(1,z)=Y_{i}(1).$
Assumption 15 (Monotonicity).
$Pr(D_{i,1}((0,1))\geq D_{i,1}((0,0)))=1.$
Assumption 16 (No anticipation in the first stage).
[…]
Assumption 17 (Parallel trends assumption in the treatment).
$E[D_{i}((0,0))|E_{i}=1,T_{i}=1]-E[D_{i}((0,0))|E_{i}=1,T_{i}=0]$
$=$ $E[D_{i}((0,0))|E_{i}=0,T_{i}=1]-E[D_{i}((0,0))|E_{i}=0,T_{i}=0].$
Assumption 18 (Parallel trends assumption in the outcome).
$E[Y_{i}(D_{i}(0,0))|E_{i}=1,T_{i}=1]-E[Y_{i}(D_{i}(0,0))|E_{i}=1,T_{i}=0]$
$=$ $E[Y_{i}(D_{i}(0,0))|E_{i}=0,T_{i}=1]-E[Y_{i}(D_{i}(0,0))|E_{i}=0,T_{i}=0].$
Assumption 19 (Relevance condition).
$E[D_{i}|E_{i}=1,T_{i}=1]-E[D_{i}|E_{i}=1,T_{i}=0]$
-- $(E[D_{i}|E_{i}=0,T_{i}=1]-E[D_{i}|E_{i}=0,T_{i}=0])>0.$
By similar arguments in the proof of Theorem 1 (Theorem 4), one can show that the Wald-DID estimand is equal to the LATET (ACRT) in period $1$ under Assumptions 14-19 (thus, we omit the proof for brevity).
A.2.2 Multiple time periods
Let $T_{i}\in\{1,\dots,T\}$ denote the time period when unit $i$ is observed and let $Z_{i,t}$ denote the instrument status under $T_{i}=t$. We adopt Assumption 8 in Section 3; that is, we assume the staggered assignment of the instrument across units. In addition, we adopt Assumption 14 (Exclusion restriction), which allows us to write $Y_{i}(D_{i},z)=Y_{i}(D_{i})$. Let $E_{i}\in\{1,\dots,T,\infty\}$ denote the cohort which unit $i$ belongs to. Let $Y_{i}(D_{i}^{e})$ and $D_{i}^{e}$ denote potential outcomes and potential treatment choices when unit $i$ belongs to cohort $e$. Let $D_{i,t}^{e}$ denote potential treatment choices under $T_{i}=t$ when unit $i$ belongs to cohort $e$. We assume that $\{Y_{i},D_{i},Z_{i},E_{i},T_{i}\}_{i=1}^{n}$ are i.i.d.
In multiple periods and cohorts settings, our target parameter is the cohort specific local average treatment effect on the treated in period $T=e+l$:
$CLATT_{e,e+l}$ $\equiv E[Y_{i}(1)-Y_{i}(0)|T_{i}=e+l,E_{i}=e,D_{i,e+l}^{e}>D_{i,e+l}^{\infty}]$
$=E[Y_{i}(1)-Y_{i}(0)|T_{i}=e+l,E_{i}=e,CM_{e,e+l}].$
When repeated cross section data are available, the staggered DID-IV designs consist of Assumptions 8, 11, 12 (in Section 3), Assumption 14, and the following parallel trends assumptions and relevance condition.
Assumption 20 (Parallel trends assumption in the treatment in multiple time periods).
$\text{For all}\hskip 2.84526pts<t,\hskip 2.84526ptE[D_{i}^{\infty}|E_{i}=e,T_{i}=t]-E[D_{i}^{\infty}|E_{i}=e,T_{i}=s]\hskip 2.84526pt\text{is identical for all}\hskip 2.84526pte\in\mathcal{S}(E_{i}).$
Assumption 21 (Parallel trends assumption in the outcome in multiple time periods).
For $\text{For all}\hskip 5.69054pts<t,\hskip 5.69054ptE[Y_{i}(D_{i}^{\infty})|E_{i}=e,T_{i}=t]-E[Y_{i}(D_{i}^{\infty})|E_{i}=e,T_{i}=s]$
is identical for $\text{is identical for all}\hskip 5.69054pte\in\mathcal{S}(E_{i}).$
Assumption 22 (Relevance condition in multiple time periods).
$\forall e\in\mathcal{S}(E_{i}),\hskip 5.69054pt\forall l\geq 0,\hskip 5.69054ptE[D_{i}-D^{\infty}_{i}|E_{i}=e,T_{i}=e+l]>0.$
A.3   Introducing the treatment path in $2\times 2$ DID-IV designs
We discuss the DID-IV identification strategy in two periods and two groups with a binary treatment, introducing the treatment path in potential outcomes $Y_{i,t}(d,z)$. In this setting, our target parameter is the LATET’ defined below.
Definition.
The local average treatment effect on the treated with the treatment path (LATET’) is
$LATET^{\prime}\equiv w_{0}\cdot\Delta_{0}+w_{1}\cdot\Delta_{1}$
where $\Delta_{0}$ and $\Delta_{1}$ are:
$\Delta_{0}\equiv E[Y_{1}((1,1))-Y_{1}((1,0))|E_{i}=1,(CM^{Z}\land DF^{T})],$
$\Delta_{1}\equiv E[Y_{1}((0,1))-Y_{1}((0,0))|E_{i}=1,(CM^{Z}\land NT^{T})].$
The weights $w_{0}$ and $w_{1}$ are:
$w_{0}=\frac{Pr((CM^{Z}\land DF^{T})|E_{i}=1)}{Pr((CM^{Z}\land DF^{T})|E_{i}=1)+Pr((CM^{Z}\land NT^{T})|E_{i}=1)},\hskip 5.69054ptw_{1}=1-w_{0}.$
We consider the identifying assumptions for the Wald-DID estimand to identify the LATET’. First, we make the following no anticipation assumptions on the outcome instead of imposing Assumption 1 (No carryover assumption).
Assumption 23 (No anticipation of the outcome among the $CM^{Z}\land NT^{T}$ in an exposed group).
$\forall z\in\mathcal{S}(Z),Y_{i,0}((0,1),z)=Y_{i,0}((0,0),z)\hskip 5.69054pt\text{for all}\hskip 5.69054pti\hskip 5.69054pt\text{with}\hskip 5.69054ptE_{i}=1\hskip 5.69054pt\text{and}\hskip 5.69054pt(CM^{Z}\land NT^{T}).$
This assumption requires that the potential outcome at period 00 is the same as the baseline outcome $Y_{i,0}((0,0),z)$, given the instrument path $z$, for those assigned to the instrument $z=(0,1)$ and belonging to the type $CM^{Z}\land NT^{T}$.
Assumption 24 (No anticipation of the outcome among the $CM^{Z}\land DF^{T}$ in an exposed group).
$\forall z\in\mathcal{S}(Z),Y_{i,0}((1,1),z)=Y_{i,0}((1,0),z)\hskip 5.69054pt\text{for all}\hskip 5.69054pti\hskip 5.69054pt\text{with}\hskip 5.69054ptE_{i}=1\hskip 5.69054pt\text{and}\hskip 5.69054pt(CM^{Z}\land DF^{T}).$
This assumption requires that the potential outcome at period 00 is the same as the baseline outcome $Y_{i,0}((1,0),z)$, given the instrument path $z$, for those assigned to the instrument $z=(0,1)$ and belonging to the type $CM^{Z}\land DF^{T}$.
We can rationalize Assumptions 23-24 by considering the similarity between these assumptions and the no anticipation assumption made in the recent DID literature; all of them restrict the anticipatory behavior for those who belong to an exposed group (treatment group) and are affected by a policy shock (the treatment in DID designs).
Next, we make the parallel trends assumption in the outcome suitable for the set up considered here. We first define the similar notation as in Subsection 2.4. Let $\bar{\Delta}_{g^{t},e}$ denote the expectation of the time trends in each groups in group $E_{i}=e$:
$\bar{\Delta}_{AT^{T},e}\equiv E[Y_{i,1}((1,1))-Y_{i,0}((1,1))|AT^{T},E_{i}=e],$
$\bar{\Delta}_{NT^{T},e}\equiv E[Y_{i,1}((0,0))-Y_{i,0}((0,0))|NT^{T},E_{i}=e],$
$\bar{\Delta}_{CM^{T},e}\equiv E[Y_{i,1}((0,1))-Y_{i,0}((0,1))|CM^{T},E_{i}=e],$
$\bar{\Delta}_{DF^{T},e}\equiv E[Y_{i,1}((1,0))-Y_{i,0}((1,0))|DF^{T},E_{i}=e].$
Let $\bar{\Delta}_{0}$ and $\bar{\Delta}_{1}$ denote the expected time gain in unexposed and exposed groups, respectively:
$\bar{\Delta}_{0}\equiv\sum_{g^{t}}w_{g^{t},0}\bar{\Delta}_{g^{t},0},\hskip 5.69054pt\bar{\Delta}_{1}\equiv\sum_{g^{t}}w_{g^{t},1}\bar{\Delta}_{g^{t},1},$
where the definition of the weight $w_{g^{t},e}$ is same as in Subsection 2.4.
As noted in Subsection 2.4, the parallel trends assumption in the outcome can be interpreted as requiring the same expected time gain between exposed and unexposed groups. The following assumption makes such a restriction in the current set up.
Assumption 25 (Parallel trend assumption in the outcome).
The expected time gain between exposed and unexposed groups is the same: $\bar{\Delta}_{0}=\bar{\Delta}_{1}$.
The theorem below shows that the Wald-DID estimand is equal to the LATET’ under Assumptions 2-5, 7 and Assumptions 23-25.
Theorem 5.
Suppose Assumptions 2-5, 7 and Assumptions 23-25 hold. Then, the Wald-DID estimand corresponds to the LATET’.
Proof.
First, we consider the denominator of the Wald-DID estimand. From the argument in the proof of Theorem 1, one can show the following result:
$E[D_{1}-D_{0}|E=1]-E[D_{1}-D_{0}|E=0]=Pr(CM^{Z}|E=1)$
$=Pr(CM^{Z}\land DF^{T}|E=1)+Pr(CM^{Z}\land NT^{T}|E=1).$ (15)
Next, for the numerator, we first decompose $E[Y_{1}-Y_{0}|E=0]$ as follows.
$E[Y_{1}-Y_{0}|E=0]=\bar{\Delta}_{0}.$ (16)
This follows from Table 2 and the definition of the expected time gain $\bar{\Delta}_{0}$.
Finally, we decompose $E[Y_{1}-Y_{0}|E=1]$ as follows.
$E[Y_{1}-Y_{0}|E=1]=\bar{\Delta}_{CM^{T},1}w_{CM^{T},1}+\bar{\Delta}_{AT^{T},1}w_{AT^{T},1}$
$+E[Y_{1}((0,1))-Y_{0}((0,1))|E=1,(CM^{Z}\land NT^{T})]Pr((CM^{Z}\land NT^{T})|E=1)$
$+E[Y_{1}((1,1))-Y_{0}((1,1))|E=1,(CM^{Z}\land DF^{T})]Pr((CM^{Z}\land DF^{T})|E=1)$
$+E[Y_{1}((1,0))-Y_{0}((1,0))|E=1,(NT^{Z}\land DF^{T})]Pr((NT^{Z}\land DF^{T})|E=1)$
$+E[Y_{1}((0,0))-Y_{0}((0,0))|E=1,(NT^{Z}\land NT^{T})]Pr((NT^{Z}\land NT^{T})|E=1)$
$=\bar{\Delta}_{CM^{T},1}w_{CM^{T},1}+\bar{\Delta}_{AT^{T},1}w_{AT^{T},1}$
$+E[Y_{1}((0,1))-Y_{1}((0,0))|E=1,(CM^{Z}\land NT^{T})]Pr((CM^{Z}\land NT^{T})|E=1)$
$+E[Y_{1}((0,0))-Y_{0}((0,0))|E=1,(CM^{Z}\land NT^{T})]Pr((CM^{Z}\land NT^{T})|E=1)$
$+E[Y_{1}((1,1))-Y_{1}((1,0))|E=1,(CM^{Z}\land DF^{T})]Pr((CM^{Z}\land DF^{T})|E=1)$
$+E[Y_{1}((1,0))-Y_{0}((1,0))|E=1,(CM^{Z}\land DF^{T})]Pr((CM^{Z}\land DF^{T})|E=1)$
$+E[Y_{1}((1,0))-Y_{0}((1,0))|E=1,(NT^{Z}\land DF^{T})]Pr((NT^{Z}\land DF^{T})|E=1)$
$+E[Y_{1}((0,0))-Y_{0}((0,0))|E=1,(NT^{Z}\land NT^{T})]Pr((NT^{Z}\land NT^{T})|E=1)$
$=\bar{\Delta}_{CM^{T},1}w_{CM^{T},1}+\bar{\Delta}_{AT^{T},1}w_{AT^{T},1}$
$+\bar{\Delta}_{NT^{T},1}w_{NT^{T},1}+\bar{\Delta}_{DF^{T},1}w_{DF^{T},1}$
$+E[Y_{1}((0,1))-Y_{1}((0,0))|E=1,(CM^{Z}\land NT^{T})]Pr((CM^{Z}\land NT^{T})|E=1)$
$+E[Y_{1}((1,1))-Y_{1}((1,0))|E=1,(CM^{Z}\land DF^{T})]Pr((CM^{Z}\land DF^{T})|E=1)$
$=\bar{\Delta}_{1}$
$+E[Y_{1}((0,1))-Y_{1}((0,0))|E=1,(CM^{Z}\land NT^{T})]Pr((CM^{Z}\land NT^{T})|E=1)$
$+E[Y_{1}((1,1))-Y_{1}((1,0))|E=1,(CM^{Z}\land DF^{T})]Pr((CM^{Z}\land DF^{T})|E=1)$ (17)
The first equality follows from Table 1 and Assumption 2. The second equality follows from Assumption 23, Assumption 24 and simple manipulation. The third equality follows from Table 1. The final equality follows from the definition of the expected time gain $\bar{\Delta}_{1}$.
From the result (16) with (17) and Assumption 25, we obtain
$E[Y_{1}-Y_{0}|E=1]-E[Y_{1}-Y_{0}|E=0]$
$=E[Y_{1}((0,1))-Y_{1}((0,0))|E=1,(CM^{Z}\land NT^{T})]Pr((CM^{Z}\land NT^{T})|E=1)$
$+E[Y_{1}((1,1))-Y_{1}((1,0))|E=1,(CM^{Z}\land DF^{T})]Pr((CM^{Z}\land DF^{T})|E=1).$ (18)
Combining the result (18) with (15), we obtain the desired result. ∎
Appendix B Deriving the influence function in repeated cross section settings
[…]
$\theta_{e,l}^{r}\equiv\frac{\alpha_{e,e+l}-\alpha_{e,e-1}-[\beta_{U,e+l}-\beta_{U,e-1}]}{\pi_{e,e+l}-\pi_{e,e-1}-[\gamma_{U,e+l}-\gamma_{U,e-1}]},$
where
$\alpha_{e,t}=E[Y_{i}|E_{i}=e,T_{i}=t],\hskip 5.69054pt\beta_{U,t}=E[Y_{i}|E_{i}\in U,T_{i}=t],$
$\pi_{e,t}=E[D_{i}|E_{i}=e,T_{i}=t],\hskip 5.69054pt\gamma_{U,t}=E[D_{i}|E_{i}\in U,T_{i}=t]\hskip 5.69054pt(t=e+l,e-1).$
In repeated cross section settings, our STS estimator is the sample analog of $\theta_{e,l}^{r}$, which we denote $\hat{\theta}_{e,l}^{r}$. Using similar arguments in the proof of Theorem 3, one can show that $\hat{\theta}_{e,l}^{r}$ is consistent for $CLATT_{e,e+l}$ under Assumptions 8, 11, 12, 14, 20-22, and have the following influence function:
$\psi_{i,e,l}^{r}=\frac{1}{\pi_{e,e+l}-\pi_{e,e-1}-[\gamma_{U,e+l}-\gamma_{U,e-1}]}$
×$\times\Bigg{(}\frac{\mathbf{1}_{e,e+l}\cdot\left[\delta_{i,e,l}-E[\delta_{i,e,l}|E_{i}=e,T_{i}=e+l]\right]}{E[\mathbf{1}_{e,e+l}]}-\frac{\mathbf{1}_{e,e-1}\cdot\left[\delta_{i,e,l}-E[\delta_{i,e,l}|E_{i}=e,T_{i}=e-1]\right]}{E[\mathbf{1}_{e,e-1}]}$
$-\frac{\mathbf{1}_{U,e+l}\cdot\left[\delta_{i,e,l}-E[\delta_{i,e,l}|E_{i}\in U,T_{i}=e+l]\right]}{E[\mathbf{1}_{U,e+l}]}+\frac{\mathbf{1}_{U,e-1}\cdot\left[\delta_{i,e,l}-E[\delta_{i,e,l}|E_{i}\in U,T_{i}=e-1]\right]}{E[\mathbf{1}_{U,e-1}]}\Bigg{)},$ (19)
where $\mathbf{1}_{A}$ is the indicator function and takes one if $A$ is true, and $\delta_{i,e,l}^{r}=Y_{i}-\theta_{e,l}^{r}\cdot D_{i}$.
Appendix C Decomposing the TWFEIV estimand in Oreopoulos (2006)
[…]
$\hat{\beta}_{IV}$ $=\frac{\sum_{t}N_{e,t}\hat{Z}_{e,t}\left[Y_{e,t}-Y_{e,1946}-(Y_{U,t}-Y_{U,1946})\right]}{\sum_{t}N_{e,t}\hat{Z}_{e,t}\left[D_{e,t}-D_{e,1946}-(D_{U,t}-D_{U,1946})\right]}$
$=\frac{\sum_{t}N_{e,t}\hat{Z}_{e,t}\left[D_{e,t}-D_{e,1946}-(D_{U,t}-D_{U,1946})\right]\cdot\widehat{WDID}_{e,t}}{\sum_{t}N_{e,t}\hat{Z}_{e,t}\left[D_{e,t}-D_{e,1946}-(D_{U,t}-D_{U,1946})\right]},$
where we define:
$\widehat{WDID}_{e,t}\equiv\frac{\left[Y_{e,t}-Y_{e,1946}-(Y_{U,t}-Y_{U,1946})\right]}{\left[D_{e,t}-D_{e,1946}-(D_{U,t}-D_{U,1946})\right]},$
and $\hat{Z}_{e,t}$ is the residuals from regressing $Z_{i,t}$ on the cohort and North Ireland fixed effects [^24] We can write $\hat{Z}_{i,t}=\hat{Z}_{e,t}$ because $Z_{i,t}$ only varies across cohorts and regions..
[…]
$\beta_{IV}=$ $\sum\limits_{U-1\geq t\geq e}w_{e,t}^{1}\cdot CLATT_{e,t}+\sum\limits_{t\geq U}w_{e,t}^{2}\cdot\Delta_{e,t},$ (20)
where we define:
$CAET^{1}_{e,t}=E[D_{i}-D_{i}^{\infty}|T_{i}=t,E_{i}=e],$
$\Delta_{e,t}=\displaystyle\frac{CAET^{1}_{e,t}\cdot CLATT_{e,t}-CAET^{1}_{U,t}\cdot CLATT_{U,t}}{CAET^{1}_{e,t}-CAET^{1}_{U,t}},$
and the weights $w^{1}_{e,t}$ and $w^{2}_{e,t}$ are:
$w_{e,t}^{1}$ (21)
$=\frac{E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot CAET^{1}_{e,t}}{\sum\limits_{U-1\geq t\geq e}E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot CAET^{1}_{e,t}+\sum\limits_{t\geq U}E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot(CAET^{1}_{e,t}-CAET^{1}_{U,t})},$ (22)
$w_{e,t}^{2}$ (23)
$=\frac{E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot(CAET^{1}_{e,t}-CATT^{1}_{U,t})}{\sum\limits_{U-1\geq t\geq e}E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot CAET^{1}_{e,t}+\sum\limits_{t\geq U}E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot(CAET^{1}_{e,t}-CAET^{1}_{U,t})},$ (24)
where $n_{e,t}$ and $E[\hat{Z}_{i,t}|E_{i}=e]$ represent the population share and residuals for England in cohort $t$ respectively.
[…]
Figure 3 plots the weight and the corresponding estimate for each $CLATT_{e,t}$ and $\Delta_{e,t}$ (bias term). The consistent estimators for each $w^{1}_{e,t}$, $w^{2}_{e,t}$, $CLATT_{e,t}$ and $\Delta_{e,t}$ are constructed from the sample analogue, exploiting $\widehat{WDID}_{e,t}$ and $(D_{e,t}-D_{e,1946})-(D_{U,t}-D_{U,1946})$.
[…]
Figure 3: Decomposition result for the TWFEIV estimand in Oreopoulos (2006). Notes: This figure plots the estimated weights and the corresponding estimates for each $CLATT_{e,t}$ and $\Delta_{e,t}$ (bias term). The closed squares are the weights and the corresponding estimates for each $CLATT_{e,t}$. The closed triangles are the weights and the corresponding estimates for each $\Delta_{e,t}$ (bias term).
Appendix D Relationship between DID and DID-IV
In this appendix, we describe the relationship between DID and DID-IV designs, and establish that DID-IV is DID with noncompliance of the assigned treatment.
Following Subsection 2.1, we introduce the notation in DID designs. Let $Z_{i}=(Z_{i,0},Z_{i,1})$ denote the path of a program that distributes the treatment $D_{i,t}$, and let $D_{i}=(D_{i,0},D_{i,1})$ denote the path of the treatment for unit $i$. In the canonical DID set up, $Z_{i}=(Z_{i,0},Z_{i,1})$ takes only two values: $Z_{i}=(0,1)$ if unit $i$ is assigned to the program in period $1$ (treatment group) and $Z_{i}=(0,0)$ if unit $i$ is not assigned to the program in period $1$ (control group). Let $E_{i}\in\{0,1\}$ denote the group variable: $E_{i}=1$ if unit $i$ belongs to the treatment group and $E_{i}=0$ if unit $i$ belongs to the control group. We refer to treatment and control groups as exposed and unexposed groups, respectively.
Note that in the common DID set up, the treatment status $D_{i,t}$ is equal to the exposure to the program $Z_{i,t}$ because we ex ante impose two restrictions on the treatment adoption behavior across units. First, we assume that all units are not allowed to receive the treatment in period 00.
$D_{i,0}((0,1))=D_{i,0}((0,0))=0\hskip 8.53581pt\text{for all $i$}.$ (25)
Second, we assume that unit $i$ should receive the treatment in period $1$ if that unit belongs to an exposed group, and that it should not receive the treatment in period $1$ if that unit belongs to an unexposed group.
$1=D_{i,1}((0,1))>D_{i,1}((0,0))=0\hskip 8.53581pt\text{for all $i$}.$ (26)
These restrictions imply that we have Tables 5-6 under DID designs: there exists only the type $CM^{Z}\land NT^{T}$ in exposed and unexposed groups, respectively.
Then, under these restrictions, we can show that DID-IV designs nest DID designs as a special case along with three dimensions. First, Assumptions 3-5 in DID-IV designs are automatically satisfied in DID designs. Assumption 3 follows from (26) and Assumption 4 follows from (25). Assumption 5 follows from combining (25) with (26).
Second, the target parameter in DID designs, the ATT is equal to the LATET in the canonical DID set up; this follows from Tables 5-6:
Table 5: Exposed group (assigned to $z=(0,1)$)
observed counterfactual
$(D_{0}((0,0)),D_{1}((0,1)))$ $D_{1}((0,0))=0$
$D_{0}((0,0))=0,D_{1}((0,1))=1$ $CM^{Z}\land NT^{T}$
Table 6: Treatment group (assigned to $z=(0,0)$)
observed counterfactual
$(D_{0}((0,0)),D_{1}((0,0)))$ $D_{1}((0,1))=1$
$D_{0}((0,0))=0,D_{1}((0,0))=0$ $CM^{Z}\land NT^{T}$
$ATT$ $=E[Y_{i,1}(1)-Y_{i,1}(0)|E_{i}=1]$
$=E[Y_{i,1}(1)-Y_{i,1}(0)|E_{i}=1,CM^{Z}]$
$=LATET.$
Note that both ATT and LATET measure the treatment effects, for those who are in an exposed group (assigned to $z=(0,1)$) and are induced to treatment by instrument ($CM^{Z}$). Therefore, we define the LATET as our target parameter in DID-IV designs in Subsection 2.2.
Finally, a well-known parallel trends assumption in DID designs can also be interpreted as requiring the same expected time gain between exposed and unexposed groups. From Tables 5-6, we can rewrite the parallel trends assumption in the untreated outcome as:
$E[Y_{i,1}(0)-Y_{i,0}(0)|E_{i}=0]=E[Y_{i,1}(0)-Y_{i,0}(0)|E_{i}=1]$
$\iff$ $E[Y_{i,1}(0)-Y_{i,0}(0)|E_{i}=0,NT^{T}]=E[Y_{i,1}(0)-Y_{i,0}(0)|E_{i}=1,NT^{T}]$
$\iff$ $\Delta_{NT^{T},0}=\Delta_{NT^{T},1}$
$\iff$ $\Delta_{0}=\Delta_{1},$
where the last equivalence holds because we have only the time never takers $NT^{T}$ in exposed and unexposed groups respectively under DID designs.
$w_{NT^{T},0}=Pr(NT^{T}|E_{i}=e)=1,\hskip 8.53581pt\text{for all}\hskip 8.53581pte=0,1.$
[…]
Appendix E Comparing DID-IV to Fuzzy DID
[…]
E.1   Comparing the identifying assumptions in dCDH and this article
In this subsection, we investigate the detailed connections between dCDH’s identifying assumptions and ours. Because dCDH consider the repeated cross section settings, when we compare the identifying assumptions between the two papers, we refer to Assumptions 14-19. For now, we ignore the difference in the definition of target parameter.
dCDH consider the following identification assumptions for the Wald-DID estimand to capture the SLATET. [^25] In repeated cross section settings, the formal definition is provided in Assumption 31. Note that in dCDH, a group variable $G$ plays the role of instrument $Z$ in our notation. [^26] This follows from two observations in dCDH. First, dCDH call the groups $G=1$ and $G=0$ treatment and control groups, respectively. Second, the group variable $G$ is included in Assumption 28 (Treatment participation equation). We therefore replace $G$ with $Z$ and call the groups $Z=1$ and $Z=0$ exposed and unexposed groups, respectively. [^27] Note that in Appendix A.2, we distinguish the group variable $E$ from the instrument path $Z$ to uncover Assumption 16 (No anticipation in the first stage). For the detailed explanations for each assumption, see dCDH.
Assumption 26 (Fuzzy design).
$E[D|Z=1,T=1]>E[D|Z=1,T=0],$ (27)
$E[D|Z=1,T=1]-E[D|Z=1,T=0]>E[D|Z=0,T=1]-E[D|Z=0,T=0].$ (28)
Assumption 27 (Stable percentage of treated units in an unexposed group).
$1\displaystyle 0<E[D|Z=0,T=1]=E[D|Z=0,T=0]<$ 1 .
Assumption 28 (Treatment participation equation).
[…]
Assumption 29 (Parallel trends assumption in the untreated outcome).
$E[Y(0)|Z=1,T=1]-E[Y(0)|Z=1,T=0]$
$=$ $E[Y(0)|Z=0,T=1]-E[Y(0)|Z=0,T=0].$
Assumption 30 (Stable treatment effect over time).
$E[Y(1)-Y(0)|Z,T=1,D_{0}(Z)=1]=E[Y(1)-Y(0)|Z,T=0,D_{0}(Z)=1].$
In Assumption 30, we modify the dCDH’s notation $D(t)$ to ease its interpretation: we define $D_{t}(Z)=\mathbf{1}\{V\geq v_{Zt}\}$ to be the treatment status under time $T=t$ in group $Z$ instead of defining it as $D(t)$.
[…]
Finally, dCDH implicitly impose Assumption 14 and Assumption 16 when they formalize potential outcomes and treatment participation equations. Specifically, dCDH do not introduce the instrument path in their framework and define potential outcomes as $Y(D)$, which implies Assumption 14. As one can see in Assumption 28, dCDH assume that the treatment participation equation only depends on the current instrument status, which implies Assumption 16 since dCDH assume the sharp assignment of the instrument: $v_{10}=v_{00}$.
The differences between dCDH’s assumptions and ours are as follows. First, dCDH assume the parallel trends assumption in the untreated outcome (Assumption 29), whereas we assume the parallel trends assumption in the unexposed outcome. Second, dCDH assume the stable treatment rate assumption in an unexposed group (Assumption 27), whereas we assume the parallel trends assumption in the treatment between exposed and unexposed groups. Third, dCDH impose the independence assumption between unobserved heterogeneity $V$ and time variable $T$ given $Z$ in Assumption 28, whereas we do not impose such restriction. Forth, dCDH argue that the Wald-DID estimand requires the stable treatment effect assumption during the two periods (Assumption 30) to capture the SLATET, whereas we do not require such restriction for this estimand to identify the LATET. Finally, dCDH assume the condition (27) in Assumption 26, which requires that the treatment rate in an exposed group increases between the two periods, whereas we do not assume this condition.
Remark E.1.1.
When Assumption 27 is violated, dCDH consider the following assumption for the Wald-DID estimand to capture the SLATET.
Assumption 31 (Homogeneous treatment effect between exposed and unexposed groups).
$SLATET=SLATET^{\prime},$
where we define282828dCDH denote SLATET and SLATET’ as $\Delta$ and $\Delta^{\prime}$ respectively. When we define the SLATET, we use $D_{0}(0)=D_{0}(1)$, which follows from the sharp assignment of the instrument: $v_{10}=v_{00}$.:
$SLATET$ $=E[Y(1)-Y(0)|Z=1,T=1,D_{1}(1)>D_{0}(0)],$
$SLATET^{\prime}$ $=E[Y(1)-Y(0)|Z=0,T=1,D_{1}(0)\neq D_{0}(0)].$
This assumption requires that the treatment effects among the switchers should be the same between exposed and unexposed groups, which dCDH call the homogeneous treatment effect assumption. In Appendix E.4, we present the decomposition result for the SLATET, and show that we cannot straightforwardly interpret this assumption as requiring the homogeneous treatment effect between the two groups. In the following discussion, we therefore treat Assumptions 26-30 as dCDH’s main identifying assumptions.
E.2   Strong restrictions under Fuzzy DID designs
In this subsection, we show that dCDH’s identifying assumptions ex ante impose strong restrictions on treatment adoption behavior across units. First, we prove some lemmas implied by dCDH’s assumptions. Based on these lemmas, we then create the tables that present all the types in exposed and unexposed groups under Fuzzy DID designs. These tables show that dCDH’s assumptions ex ante exclude many types in both groups and these restrictions are heavily assigned on an unexposed group. At the end, we provide the empirical implication of the distinction between dCDH’s restrictions and ours.
First, we show that Assumption 28 implies the following assumption.
Assumption 32 (Monotone change in treatment in exposed and unexposed groups).
For all $z\in\{0,1\}$,
[…]
holds.
Here, recall that we define $D_{t}(Z)=\mathbf{1}\{V\geq v_{Zt}\}$ to be the treatment status under $T=t$ in group $Z$.
Lemma 1.
Assumption 28 implies Assumption 32.
Proof.
Fix $z\in\{0,1\}$. Then we have $v_{z1}\leq v_{z0}$ or $v_{z1}\geq v_{z0}$.
Therefore, conditioning $Z=z$, we have
[…]
because we assume that the unobserved heterogeneity $V$ is independent of time variable $T$ conditional on the instrument $Z$. ∎
Lemma 1 shows that under Fuzzy DID designs, the treatment status changes between period 00 and period $1$ in a monotone (uniform) way in exposed and unexposed groups, respectively. Note that the treatment participation equation itself does not imply Assumption 32: the independence between the unobserved heterogeneity $V$ and time variable $T$ given $Z$ yields this assumption. Indeed, in this article, we do not assume this restriction.
As we have seen in Subsection 2.4, time also plays the role of instrument in DID-IV settings. From this observation, we can view that in an unexposed group ($z=0$), Assumption 32 corresponds to a monotonicity assumption with respect to time, formally defined below.
Assumption 33 (Monotonicity assumption with respect to time in an unexposed group).
[…]
Assumption 33 requires that time affects the treatment adoption process from period 00 to period $1$ in a monotone (uniform) way in an unexposed group $z=0$. In terms of the notation introduced in Subsection 2.4, this assumption implies that the group variable $G^{T}$ can take three values with non-zero probability. Hereafter, we consider the type of Assumption 33 which rules out the existence of the time-defiers $DF^{T}$ in an unexposed group.
Next, we show that dCDH’s assumptions imply that there are no time compliers $CM^{T}$ and time defiers $DF^{T}$ in an unexposed group ($z=0$).
Lemma 2.
Suppose Assumption 27, Assumption 28 and Assumption 33 hold. Then, there are no time compliers $CM^{T}$ and time defiers $DF^{T}$ in an unexposed group ($z=0$).
Proof.
Without loss of generality, suppose that Assumption 33 excludes the time defiers $DF^{T}$ in an unexposed group ($z=0$).
Then, Assumption 27 implies:
$E[D|T=1,Z=0]=E[D|T=0,Z=0]$
$\iff$ $Pr(D_{1}(0)>D_{0}(0)|Z=0)=0$
$\iff$ $Pr(CM^{T}|Z=0)=0$ (29)
The first equivalence follows from $V\mathop{\perp\!\!\!\!\perp}T|Z$ in Assumption 28. Equation (29) implies that there are no time compliers $CM^{T}$ in an unexposed group ($z=0$). ∎
Finally, we show that the condition (27) in Assumption 26 (Fuzzy design) and Assumption 28 imply that in an exposed group ($z=1$), there are no units switching from treatment to non-treatment between period 00 and period $1$. In other words, in Assumption 32, we should have:
$Pr(D_{0}(1)\leq D_{1}(1)|Z=1)=1.$
Lemma 3.
The condition (27) in Assumption 26 and Assumption 28 imply $Pr(D_{0}(1)\leq D_{1}(1)|Z=1)$=1.
Proof.
From the condition (27) in Assumption 26, we have
$E[D|Z=1,T=1]>E[D|Z=1,T=0]$
$\iff$ $E[D_{1}(1)-D_{0}(1)|Z=1]>0$
$\iff$ $Pr(D_{1}(1)>D_{0}(1)|Z=1)-Pr(D_{0}(1)>D_{1}(1)|Z=1)>0,$ (30)
where the first equivalence follows from $V\mathop{\perp\!\!\!\!\perp}T|Z$ in Assumption 28. Recall that Assumption 28 implies Assumption 32 from Lemma 1. Suppose that $Pr(D_{0}(1)\geq D_{1}(1)|Z=1)=1$ holds. Then, we have $Pr(D_{1}(1)>D_{0}(1)|Z=1)=0$. This contradicts with the condition (30). ∎
We now clarify how dCDH’s identifying assumptions restrict the treatment adoption behavior across units in exposed and unexposed groups, respectively. Tables 3-4 in Subsection 2.5 show that their identifying assumptions exclude many types in both groups and the restrictions are heavily imposed on an unexposed group: five and three types exist in exposed and unexposed groups respectively, where the types painted in gray color are excluded in both groups.
Comparing Tables 3-4 to Tables 1-2, one can see the difference between dCDH’s restrictions and ours [^29] Tables 1-2 are constructed under panel data settings, but the types represented in these tables remain consistent even when considering repeated cross section settings. The only modification is that as in Tables 3-4, one can only observe either $D_{0}(z)$ or $D_{1}(z)$ for unit $i$ in group $z$ in each table.: in both groups, we only exclude the defiers $DF^{Z}$ by Assumption 15 (Monotonicity), whereas dCDH additionally exclude the type $NT^{Z}\land DF^{T}$ in an exposed group by Lemma 3, and additionally exclude the time defiers $DF^{T}$ and the time compliers $CM^{T}$ in an unexposed group by Lemma 2.
The distinction between dCDH’s restrictions and ours appears to be important in many empirical settings. Unlike the monotonicity assumption, which eliminate the compliers $CM^{Z}$ in both groups, dCDH’s additional restrictions might be difficult to defend in some applications, and the practitioners may want to avoid imposing these restrictions in practice. For instance, in some cases, we may have no valid reason for excluding the specific type $NT^{Z}\land DF^{T}$ in an exposed group, who are not affected by instrument but affected by time. Similarly, in other cases, we may have no justification to exclude the time defiers $DF^{T}$ and time compliers $CM^{T}$ in an unexposed group because of the steady growth of the treatment rate.
E.3   Decomposing the SLATET
In this subsection, we show that dCDH’s target parameter, the SLATET, can be decomposed into a weighted average of two different causal parameters under Fuzzy DID designs.
The following theorem presents the decomposition result for the SLATET, and clarifies the difference between dCDH’s target parameter (SLATET) and ours (LATET).
Theorem 6.
Suppose Assumptions 26-30 hold. Then, we can decompose the SLATET into two different causal parameters:
$SLATET\equiv E[Y(1)-Y(0)|Z=1,T=1,D_{1}(1)>D_{0}(0)]$
$=E[Y(1)-Y(0)|Z=1,T=1,CM^{T}]Pr(CM^{T}|Z=1,SW)$
$+E[Y(1)-Y(0)|Z=1,T=1,CM^{Z}\land NT^{T}]Pr(CM^{Z}\land NT^{T}|Z=1,SW),$ (31)
where the switcher (SW) in an exposed group ($Z=1$) is the units who become treated in time $T=1$.
Proof.
The proof directly follows from Table 3 in an exposed group. ∎
Theorem 6 shows that the SLATET can be decomposed into a weighted average of two different causal parameters under Fuzzy DID designs: one parameter measures the treatment effects among the time compliers $CM^{T}$ in an exposed group and the other measures the treatment effects among the type $(CM^{Z}\land NT^{T})$ in an exposed group.
This theorem has a concerning implications; the SLATET may not be policy-relevant even when the instrument is based on the policy change of interest to the researcher. Suppose that the instrument represents an important policy lever for researchers, and they are intrinsically interested in the treatment effects among the compliers $CM^{Z}$ in an exposed group. In this setting, our target parameter, the LATET, is a policy-relevant by construction. However, the SLATET may not be policy-relevant because this parameter captures the treatment effects among the sub-population of the compliers and the time compliers, who are affected by time but not affected by the instrument. As the share of the time compliers $CM^{T}$ increases, the SLATET puts more weight on the treatment effects among the time compliers, which contaminates the interpretation of the magnitude of this paramter even in such policy-relevant contexts.
Using the decomposition result for the SLATET, we finally clarify the interpretation of Assumption 31. We show that this assumption cannot be interpreted as requiring the homogeneous treatment effect among the same sub-population between exposed and unexposed groups.
If we do not assume Assumption 27, we have Table 5 in an unexposed group. [^30] Here, we assume that Assumption 28 and 33 exclude the defiers $DF^{Z}$ and the time defiers $DF^{T}$ in an unexposed group, respectively.
Table 5: Unexposed group ($z=0$)
observed counterfactual
$D_{0}(0)$  or  $D_{1}(0)$ $D_{1}(1)=1$ $D_{1}(1)=0$
$D_{0}(0)=1,D_{1}(0)=1$ $AT^{Z}\land AT^{T}$ $DF^{Z}\land AT^{T}$
$D_{0}(0)=1,D_{1}(0)=0$ $CM^{Z}\land DF^{T}$ $NT^{Z}\land DF^{T}$
$D_{0}(0)=0,D_{1}(0)=1$ $AT^{Z}\land CM^{T}$ $DF^{Z}\land CM^{T}$
$D_{0}(0)=0,D_{1}(0)=0$ $CM^{Z}\land NT^{T}$ $NT^{Z}\land NT^{T}$
Then, from Table 5, we can interpret the SLATET’ (see Assumption 31) as the treatment effects among the time compliers in an unexposed group in period $1$:
$SLATET^{\prime}$ $\equiv E[Y(1)-Y(0)|Z=0,T=1,D_{1}(0)\neq D_{0}(0)]$
$=E[Y(1)-Y(0)|Z=0,T=1,CM^{T}].$ (32)
By combining equation (31) with (32), one can see that Assumption 31 ($STATET=SLATET^{\prime}$) does not require the homogeneous treatment effects in the same sub-population between the two groups. Rather, this assumption requires that a weighted average of the treatment effects among the time compliers $CM^{T}$ and the type $(CM^{Z}\land NT^{T})$ in an exposed group should be equal to the treatment effect among the time compliers $CM^{T}$ in an unexposed group. This interpretation implies that the requirement of the homogeneous treatment effect in Assumption 31 is less clear than what one might expect.
E.4   Revisiting the issue regarding the use of the Wald-DID estimand
Finally, we revisit the issue pointed by dCDH regarding the use of the Wald-DID estimand, and show that this problem is not serious than one might expect, and does not arise under our DID-IV designs. We also clarify why dCDH exclude so many types in exposed and unexposed groups as seen in Tables 3-4. Hereafter, we assume Assumptions 26-30 in Fuzzy DID designs.
dCDH argue that the Wald-DID estimand requires Assumption 30 to identify the SLATET, and propose alternative estimands that do not rely on this assumption. We first show that their argument depends on the parallel trends assumption in DID designs. Next, we argue that this assumption would be difficult to justify in DID-IV settings.
First, we clarify the interpretation of Assumption 30 by introducing the additional notation. This assumption is actually conditional on $D_{0}(Z)=1$ in addition to $Z$ and $T$, but it is not immediately clear from dCDH why this strong restriction is required only for the sub-population. For each time group $g^{t}$, let $w_{g^{t},z}=Pr(G^{T}=g^{t}|Z=z)$ denote the population share conditional on group $Z=z$, and let $w_{g^{t},z,D_{0}(0)=1}=Pr(G^{T}=g^{t}|Z=z,D_{0}(0)=1)$ denote the population share conditional on $D_{0}(0)=1$ in addition to group $Z=z$. Additionally, let $\delta_{z,t,g^{t}}=E[Y(1)-Y(0)|Z=z,T=t,G^{T}=g^{t}]$ denote the average treatment effect conditional on group $Z=z$, time $T=t$, and time group $G^{T}=g^{t}$.
Then, from Tables 3-4, we can decompose Assumption 30 as follows:
$E[Y(1)-Y(0)|Z=0,T=1,D_{0}(0)=1]$
$=$ $E[Y(1)-Y(0)|Z=0,T=0,D_{0}(0)=1]$
$\iff$ $\delta_{0,1,AT^{T}}=\delta_{0,0,AT^{T}},$ (33)
$E[Y(1)-Y(0)|Z=1,T=1,D_{0}(1)=1]$
$=$ $E[Y(1)-Y(0)|Z=1,T=0,D_{0}(1)=1]$
$\iff$ $\delta_{1,1,AT^{T}}\cdot w_{AT^{T},1,D_{0}(0)=1}+\delta_{1,1,DF^{T}}\cdot w_{DF^{T},1,D_{0}(0)=1}$
$=$ $\delta_{1,0,AT^{T}}\cdot w_{AT^{T},1,D_{0}(0)=1}+\delta_{1,0,DF^{T}}\cdot w_{DF^{T},1,D_{0}(0)=1},$ (34)
where we use the Law of Iterated Expectation, $D_{0}(0)=D_{0}(1)$ (implied by the sharp assignment of the instrument), and $V\mathop{\perp\!\!\!\!\perp}T|Z$ in Assumption 28. Equation (33) shows that in an unexposed group ($z=0$), Assumption 30 requires the stable treatment effect for the time always-takers $AT^{T}$. Equation (34) shows that in an exposed group ($z=1$), this assumption requires that a weighted average of the treatment effects among the time defiers $DF^{T}$ and the time always takers $AT^{T}$ is stable during the two periods.
Conditions (33) and (34) are required for Fuzzy DID designs because dCDH rely on the parallel trends assumption in the untreated outcome. In DID-IV settings, we can not observe the average time trends of untreated potential outcomes for the time always takers $AT^{T}$ in an unexposed group and the time always takers $AT^{T}$ and the time defiers $DF^{T}$ in an exposed group because they always adopt the treatment during the two periods. Therefore, if we assume the parallel trends assumption in the untreated outcomes, we should recover them by using conditions (33) and (34).
To see this formally, we use our notation to briefly sketch dCDH’s proof about how the Wald-DID estimand identifies the SLATET under fuzzy DID designs. First, consider the conditional expectation $E[Y|Z=0,T=1]-E[Y|Z=0,T=0]$, which appears in the numerator of the Wald-DID estimand. We can decompose this quantity into two terms.
$E[Y|Z=0,T=1]-E[Y|Z=0,T=0]$
$=$ $(\delta_{0,1,AT^{T}}-\delta_{0,0,AT^{T}})\cdot w_{AT^{T},0}+\{E[Y(0)|Z=0,T=1]-E[Y(0)|Z=0,T=0]\}.$
The first term arises because we can not observe the average time trends of the untreated potential outcome for the time always-takers $AT^{T}$ in an unexposed group, and we should add and subtract this quantity to construct the second term. The first term disappears if we assume condition (33).
Similarly, one can decompose the conditional expectation $E[Y|Z=1,T=1]-E[Y|Z=1,T=0]$ as follows.
$E[Y|Z=1,T=1]-E[Y|Z=1,T=0]$
$=$ $\Big{\{}\delta_{1,1,AT^{T}}\cdot w_{AT^{T},1,D_{0}(0)=1}+\delta_{1,1,DF^{T}}\cdot w_{DF^{T},1,D_{0}(0)=1}$
$-\delta_{1,0,AT^{T}}\cdot w_{AT^{T},1,D_{0}(0)=1}-\delta_{1,0,DF^{T}}\cdot w_{DF^{T},1,D_{0}(0)=1}\Big{\}}$
$+$ $\{E[Y(0)|Z=1,T=1]-E[Y(0)|Z=1,T=0]\}$
$+$ $E[Y(1)-Y(0)|SW,Z=1,T=1]Pr(SW|Z=1),$
where the type SW (switcher) includes the units who switch from non-treatment to treatment during the two periods. The first term arises because we can not observe the average time trends of the untreated potential outcome for the time always-takers $AT^{T}$ and the time never-takers $NT^{T}$ in an exposed group, and we should add and subtract this quantity to construct the second term. The first term disappears if we assume condition (34). Here, we do not require the stable treatment assumption for the time compliers $CM^{T}$ and the type $CM^{Z}\land NT^{T}$ in an exposed group: by adding and subtracting the expectation of the untreated potential outcome path for these types, we have the third term.
Then, by noticing that the denominator of the Wald-DID estimand identifies $Pr(SW|Z=1)$ under Assumption 27 and Assumption 28, the Wald-DID estimand captures the SLATET under Assumption 29 (parallel trends assumption in the untreated outcome) and conditions (33) and (34). From the above discussion, one can see that imposing the parallel trends assumption in the untreated outcome leads dCDH to assume conditions (33) and (34).
Here, we clarify why dCDH ex ante impose strong restrictions on treatment adoption behavior across units for the Wald-DID estimand to identify the SLATET. This also stems from their reliance on the parallel trends assumption in the untreated outcome.
Remark E.4.1.
dCDH exclude the type $NT^{Z}\land DF^{T}$ in an exposed group and the time defiers $DF^{T}$ in an unexposed group because if we add and subtract the average time trends of the untreated outcome for these types, we additionally identify the treatment effects among them in period 00 (not in period $1$), as they adopt the treatment in period 00, but leave it in period $1$ (see Tables 3-4). This identification result, however, is not intuitive because the identification variation comes from a policy shock (instrument) in period $1$.
Remark E.4.2.
dCDH exclude the time compliers $CM^{T}$ in an unexposed group in order to avoid the SLATET in an exposed group being offset by the SLATET’ in an unexposed group. If we add and subtract the average time trends of the untreated outcome for the time compliers, we additionally capture the SLATET’ in an unexposed group from the conditional expectation $E[Y|Z=0,T=1]-E[Y|Z=0,T=0]$.
Thus far, we have confirmed that dCDH’s argument rests on the parallel trends assumption in the untreated outcome. Finally, we argue that this assumption would be difficult to rely on in DID-IV settings for three reasons.
[…]
Finally, this assumption is generally not testable using the pre-exposed period data, as some units can already adopt the treatment before period 00, and we cannot compare the pre-trend of the untreated outcome between exposed and unexposed groups.
Although dCDH propose the placebo test in their appendix, it has shortcomings for checking the plausibility of the parallel trends assumption in the untreated outcome. dCDH’s test has two steps. [^31] Although dCDH ”assume” the conditions (35) and (36), they note that when one of the two conditions fails, the second test can ”no longer be used to test” Assumption 29. We, therefore, describe the dCDH’s test in two steps. First, we test whether the share of treated units is zero in both groups from $T=-1$ to $T=0$.
$E[D|Z=1,T=0]-E[D|Z=1,T=-1]=0,$ (35)
$E[D|Z=0,T=0]-E[D|Z=0,T=-1]=0.$ (36)
Second, we test the following null hypothesis:
$E[Y|Z=1,T=0]-E[Y|Z=1,T=-1]$ (37)
$=$ $E[Y|Z=0,T=0]-E[Y|Z=0,T=-1].$ (38)
dCDH’s test has some weaknesses. First, the conditions (35) and (36), which should be checked in the first step, are strong, but no guidance is provided when these conditions are violated in practice. Second, in this test, the second step does not directly test Assumption 29 in the pre-exposed periods, as it does not compare the evolution of the mean untreated outcome between exposed and unexposed groups before period $T=0$. Indeed, dCDH describe the equality (37) as a necessary condition for Assumption 29 and Assumption 30 to be satisfied under the conditions (35) and (36). Finally, it is not clear how this test should be conducted with multiple pre-exposed period data. Although dCDH assume “for instance that data is available for the period $T=-1$”, it is not indicated how this test can be performed in multiple pre-exposed period settings.
Surprisingly, dCDH’s test corresponds to our pre-trend test in Subsection 5.3 if we replace the conditions (35)-(36) in the first step with the weaker one.
$E[D|Z=1,T=0]-E[D|Z=1,T=-1]$ (39)
$=$ $E[D|Z=0,T=0]-E[D|Z=0,T=-1].$ (40)
Under repeated cross section settings, the conditions (37)-(39) coincide with the conditions (1)-(2). This indicates that the rationale behind dCDH’s test is more suitable for the assessment of parallel trends assumptions in DID-IV designs.
E.5   Other connections
In this subsection, we document the relationship between DID-IV and Fuzzy DID under non-binary, ordered treatment, and multiple groups settings. We also compare the Wald-DID estimand to the time-corrected Wald estimand considered in dCDH.
E.5.1 Non-binary, ordered treatment
In this subsection, we investigate the relationship between DID-IV to Fuzzy DID under non-binary, ordered treatment settings.
First, we compare the target parameter between the two designs. With a non-binary, ordered treatment $D\in\{0,1,\dots,J\}$, dCDH consider the following target parameter, which we call the switcher average causal response (SACR).
$SACR\equiv\sum_{j=1}^{J}w_{j}\cdot E(Y(j)-Y(j-1)|Z=1,T=1,D_{0}(0)<j<D_{1}(1)),$
where $w_{j}$ is
$w_{j}=\frac{P(D\geq j|Z=1,T=1)-P(D\geq j|Z=1,T=0)}{E[D|Z=1,T=1]-E[D|Z=1,T=0]}.$
[…]
Next, we compare the identifying assumptions between the two designs. In order for the Wald-DID estimand to identify the SACR, dCDH replace Assumption 28 with the following one.
Assumption 34 (Ordered treatment equation).
$D=\sum_{j=1}^{J}\mathbf{1}\{V\geq v^{j}_{ZT}\},\hskip 5.69054pt\text{with}\hskip 5.69054pt-\infty=v^{0}_{ZT}<v^{1}_{ZT}\dots<v^{J+1}_{ZT}=\infty\hskip 5.69054pt\text{and}\hskip 5.69054ptV\mathop{\perp\!\!\!\!\perp}T|Z.$
In addition, dCDH replace the condition (27) in Assumption 26 with the condition (41), and Assumption 27 with the condition (42), respectively.
$D_{11}\succeq D_{10},$ (41)
$D_{01}\sim D_{00},$ (42)
where $D_{ZT}$ denotes the treatment conditional on $Z$ and $T$, $\succeq$ denotes stochastic dominance between two random variables, and $\sim$ denotes equality in distribution.
Note that the condition (42) is stronger than Assumption 27 because this assumption requires that the distribution of the treatment should be stable during the two periods in an unexposed group.
In our DID-IV designs, the identifying assumptions required for the Wald-DID estimand to capture the ACRT are unchanged under non-binary, ordered treatment settings. In addition, Assumptions 17-18 are indirectly testable using the pre-exposed period data in practice.
E.5.2 Multiple groups
In multiple groups settings, our DID-IV design is not directly compatible to Fuzzy DID design because dCDH exploit the group structure in the data ($G\in\{0,1,\dots,\bar{g}\}$) for the Wald-DID estimand to identify the following target parameter.
$\Delta^{\ast}=E[Y(1)-Y(0)|\cup_{g=0}^{\bar{g}}S_{g},T=1],$
where $S_{g}=\{D_{0}(g)\neq D_{1}(g),G=g\}$ represents the switchers in group $G=g$ [^32] Following $D_{t}(Z)$, we define $D_{t}(G)=\mathbf{1}\{V\geq v_{Gt}\}$ to be the treatment status under time $T=t$ in group $G$.
[…]
In some applications, a researcher might prefer our approach to dCDH’s approach for three reasons. First, our approach yields a policy-relevant parameter by construction if the instrument represents a policy change of interest to the researcher. Second, in multiple period settings, our approach avoids the need to estimate the super groups in each consecutive period $(t-1,t)$, which simplifies both estimation and inference, making our method easier to implement in practice.
Finally, in our approach, the monotonicity assumption can be easy to justify if the instrument is based on a policy shock. In dCDH’s approach, this assumption might be harder to defend in practice because the group variable $G$ typically represents the geographic feature in the data such as municipality, state, or country. [^34] Note that, in multiple groups settings, dCDH also maintain the monotonicity assumption on each group pair $(g,g^{\prime})$ through Assumption 28.
Remark E.4.3 (Multiple groups settings with an ordered treatment).
[…]
E.5.3 Alternative estimand
In this subsection, we compare the Wald-DID estimand to the alternative estimand proposed by dCDH that does not rely on Assumption 29 and 30.353535In this subsection, we do not consider Wald-CIC estimand proposed by dCDH, as we focus on the mean treatment effect in this article.
As an alternative to the Wald-DID estimand, dCDH propose the time-corrected Wald (TC-Wald) estimand defined below.
$W_{TC}\equiv\frac{E[Y|Z=1,T=1]-E[Y+(1-D)\delta_{0}+D\delta_{1}|Z=1,T=0]}{E[D|Z=1,T=1]-E[D|Z=1,T=0]},$
where $\delta_{d}=E[Y|Z=0,T=1,D_{1}(0)=d]-E[Y|Z=0,T=0,D_{0}(0)=d]$.
In order for the TC-Wald estimand to identify the SLATET, dCDH consider the following identifying assumption instead of imposing Assumption 29 and 30.
Assumption 35 (Conditional common trends).
For all $d\in\mathcal{S}(D),$
$E[Y(d)|Z=1,T=1,D_{0}(0)=d]-E[Y(d)|Z=1,T=1,D_{0}(0)=d]$
$=$ $E[Y(d)|Z=0,T=1,D_{0}(0)=d]-E[Y(d)|Z=0,T=1,D_{0}(0)=d].$
Note that with a binary treatment, this assumption exploits the observed outcome path in an unexposed groups similar to Assumption 18 in our DID-IV designs (see Table 4).
While the target parameter differs between the two designs, we see that using the Wald-DID estimand under DID-IV designs would be more appealing than using the TC-Wald estimand under Fuzzy DID designs for two reasons. First, when using the Wald-DID estimand in DID-IV designs, researchers can easily assess the plausibility of parallel trends assumptions in practice, employing the pre-trends tests described in Section 5.3. Although the TC-Wald estimand avoids the stable treatment effect assumption under Fuzzy DID designs, it presents challenges in testing the plausibility of Assumption 35 in pre-exposed period data: it is conditional on $D_{0}(0)=d$ and imposes restrictions on the mean evolution of $Y(d)$ for all $d\in\mathcal{S}(D)$.
[…]