### v6

Abstract
Many studies exploit variation in the timing of policy adoption across units as an instrument for treatment. This paper formalizes the underlying identification strategy as an instrumented difference-in-differences (DID-IV). In this design, a Wald-DID estimand, which scales the DID estimand of the outcome by the DID estimand of the treatment, captures the local average treatment effect on the treated (LATET). We extend the canonical DID-IV design to multiple period settings with the staggered adoption of the instrument across units. Moreover, we propose a credible estimation method in this design that is robust to treatment effect heterogeneity. We illustrate the empirical relevance of our findings, estimating returns to schooling in the United Kingdom. In this application, the two-way fixed effects instrumental variable regression, the conventional approach to implement DID-IV designs, yields a negative estimate. By contrast, our estimation method indicates a substantial gain from schooling.
Abstract
In this online appendix, we first present additional identification results. Second, we derive the influence function of our STS estimator in repeated cross section settings. Third, we decompose the TWFEIV estimand in the setting of Oreopoulos (2006). Finally, we investigate the connections between DID-IV and Fuzzy DID designs considered in de Chaisemartin and D’Haultfœuille (2018).
Keywords: difference-in-differences, instrumental variable, local average treatment effect, returns to education
1  Introduction
To identify the effect of a treatment on an outcome, many studies exploit variation in the timing of policy adoption across units as an instrument for treatment. For example, to estimate returns to schooling in Indonesia, Duflo (2001) exploits variation in the timing of the introduction of a school construction program across regions as an instrument for educational attainment. Similarly, to estimate the causal link between parents’ and children’s educational attainment, Black et al. (2005) exploit variation in the timing of the implementation of school reforms across municipalities as an instrument for parents’ educational attainment. Notably, the identification strategy underlying these studies is similar to difference-in-differences (DID) designs, in that variation in the timing of a policy shock is used to identify treatment effects. At the same time, it differs from DID designs in that this variation is used to construct an instrument rather than a treatment.
In this paper, we formalize this identification strategy as an instrumented difference-in-differences (DID-IV). We define the target parameter and the identifying assumptions in this design, and develop a credible estimation method that is robust to treatment effect heterogeneity. We illustrate the empirical relevance of our findings with the setting of Oreopoulos (2006), estimating returns to schooling in the United Kingdom. In this application, we show that the choice of estimation method matters in practice.
First, we consider a simple setting with two periods and two groups: some units are not exposed to the instrument in either period (the unexposed group), while others become exposed in the second period (the exposed group). [^1] Note that this two-period/two-group ($2\times 2$) setting has already been considered in de Chaisemartin (2010) and Hudson et al. (2017), and our $2\times 2$ DID-IV design builds on these studies. In this paper, we revisit $2\times 2$ DID-IV designs for two reasons. First, we aim to complement de Chaisemartin (2010) and Hudson et al. (2017). Specifically, we introduce the instrument path into $2\times 2$ DID-IV designs and uncover an additional identifying assumption that is not made explicit in the previous literature. Second, we compare $2\times 2$ DID-IV designs to the Fuzzy DID designs considered in de Chaisemartin and D’Haultfœuille (2018). While both de Chaisemartin (2010) and de Chaisemartin and D’Haultfœuille (2018) study the same setting, de Chaisemartin and D’Haultfœuille (2018) formalize $2\times 2$ DID-IV designs differently, calling them Fuzzy DID designs. In this setting, our DID-IV design relies on a monotonicity assumption and parallel trends assumptions for both the treatment and the outcome between the exposed and unexposed groups. Our target parameter is the local average treatment effect on the treated (LATET), which measures the treatment effect for units that belong to the exposed group and are induced to receive the treatment by the instrument in the second period. We show that, in this design, a Wald-DID estimand—defined as the ratio of the DID estimand of the outcome to that of the treatment—identifies the LATET.
de Chaisemartin and D’Haultfœuille (2018) (hereafter, “dCDH”) formalize $2\times 2$ DID-IV designs differently, calling them Fuzzy DID designs. The main difference between this paper and dCDH lies in the definition of the target parameter. While we focus on the LATET, dCDH focus on the switcher local average treatment effect on the treated (SLATET); this parameter measures the treatment effects, for those who belong to an exposed group and start receiving the treatment in the second period. Because the target parameters differ, the identifying assumptions adopted in dCDH also differ from those used in this paper.
Motivated by these differences, we next examine the detailed connections between DID-IV and Fuzzy DID and discuss their implications for treatment adoption behavior and the interpretation of the target parameter. We first show that the identifying assumptions under Fuzzy DID impose stronger restrictions on treatment adoption behavior than those under DID-IV. Under these restrictions, we then demonstrate that dCDH’s target parameter, the SLATET, can be decomposed into a weighted average of two distinct causal parameters. One parameter captures the treatment effects among the subpopulation of the compliers in the sense of Imbens and Angrist (1994), while the other captures the treatment effects for time compliers—units whose treatment status is affected by time but not by the instrument. This decomposition result has an important implication: even when the instrument is directly linked to the policy change of interest, the SLATET may fail to be a policy-relevant parameter (Heckman and Vytlacil (2001)).
Next, we extend the canonical DID-IV design to multiple period settings with the staggered adoption of the instrument across units. In most DID-IV applications, researchers exploit variation in the timing of policy adoption across units in more than two periods, instrumenting for the treatment with the natural variation. The instrument is constructed, for example, from the staggered adoption of school reforms across municipalities or countries (e.g. Oreopoulos (2006), Lundborg et al. (2014), and Meghir et al. (2018)), the phased-in introduction of Head Start across states (e.g. Johnson and Jackson (2019)), or the gradual adoption of broadband internet programs across municipalities (e.g. Akerman et al. (2015), Bhuller et al. (2013)). We refer to this identification strategy as a staggered DID-IV design, and establish the corresponding target parameter and identifying assumptions. Specifically, we first partition units into mutually exclusive and exhaustive cohorts based on the initial exposure date of the instrument. We then define our target parameter as the cohort specific local average treatment effect on the treated (CLATT). This parameter is a natural generalization of the LATET in $2\times 2$ DID-IV designs and measures the treatment effects for units that belong to cohort $e$ and are the compliers at a given relative period $l$ following the initial exposure to the instrument. Finally, we introduce two Wald–DID estimands that use either never-exposed cohorts or not-yet-exposed cohorts as control groups. For each estimand, we extend the identification assumptions from the $2\times 2$ DID–IV setting to multiple time periods, and show that, under these assumptions, the corresponding Wald–DID estimand identifies the CLATT at each relative period following the initial exposure to the instrument.
We extend our DID-IV framework along two dimensions. First, we show that it can be applied to settings with a non-binary, ordered treatment. Second, we consider extensions to repeated cross sections.
Finally, we propose a regression-based method to consistently estimate our target parameter in staggered DID-IV designs under heterogeneous treatment effects. In practice, when researchers implicitly rely on a staggered DID-IV design, they typically implement it using two-way fixed effects instrumental variable (TWFEIV) regressions (e.g. Johnson and Jackson (2019), Lundborg et al. (2014), Black et al. (2005), Akerman et al. (2015), and Bhuller et al. (2013)). In a companion paper (Miyaji (2024)), however, we show that in more than two periods, the TWFEIV estimand generally fails to summarize treatment effects if the effect of the instrument on the treatment or on the outcome is not stable over time. Our proposed method avoids this issue and is robust to treatment effect heterogeneity. The estimation procedure consists of two steps: we subset the data that contain only two cohorts and two periods and then, in each data set, we run the TWFEIV regression. We call this a stacked two stage least squares (STS) regression and ensure its validity. Following Callaway and Sant’Anna (2021), we propose a weighting scheme to summarize the treatment effects, where the weight reflects the share of compliers in a given relative period $l$ in cohort $e$. We also discuss the procedure of pre-trends tests to assess the validity of the parallel trends assumptions in DID-IV designs.
We illustrate our findings with the setting of Oreopoulos (2006), who estimates returns to schooling in the UK, exploiting variation in the timing of the implementation of school reforms between Britain and Northern Ireland as an instrument for education attainment. In this application, we first assess the plausibility of the DID-IV identification strategy implicitly adopted by Oreopoulos (2006). We then estimate the TWFEIV regression in the author’s setting. We find that the TWFEIV estimate is strictly negative and not significantly different from zero. Finally, we use our estimation method to reassess returns to schooling in the UK. We find that our STS estimates are all positive in each relative period after school reform, and our weighting scheme yields a more plausible estimate than the TWFEIV estimate. Specifically, our weighted estimate indicates roughly a $20\%$ gain from schooling in the UK.
The rest of the paper is organized as follows. The following subsection discusses the related literature. Section 2 establishes DID-IV designs in two periods and two groups settings. Section 3 formalizes the target parameter and identifying assumptions in staggered DID-IV designs. Section 4 contains extensions. Section 5 presents our estimation method. Section 6 presents our empirical application. Section 7 concludes. All proofs are given in the Appendix.
Related literature
Our paper is related to the recent DID-IV literature (de Chaisemartin (2010); Hudson et al. (2017); de Chaisemartin and D’Haultfœuille (2018); de Chaisemartin et al. (2024); Chen et al. (2025); Helmers et al. (2025)) and contributes to this literature in three ways. [^2] While Ye et al. (2023) also consider the “instrumented difference-in-differences”, their target parameter is the average treatment effect (ATE), and they impose strong assumptions for the Wald-DID estimand to identify this parameter. We therefore view Ye et al. (2023) as distinct from the recent DID-IV literature.
First, this paper investigates the detailed connections between DID-IV and the Fuzzy DID considered in de Chaisemartin and D’Haultfœuille (2018). In econometrics, a pioneering work formalizing $2\times 2$ DID-IV designs is de Chaisemartin (2010), who shows that, under parallel trends assumptions for both the treatment and the outcome, together with a monotonicity assumption, the Wald-DID estimand identifies the local average treatment effect on the treated (LATET). [^3] Blundell and Dias (2009) also consider a DID–IV setting with a binary treatment and a binary instrument in Subsection F.2 of Section III. There are two differences between de Chaisemartin (2010) and Blundell and Dias (2009). First, while de Chaisemartin (2010) focus on the LATET as the target parameter, Blundell and Dias (2009) focus on the switcher local average treatment effect (SLATET), as in dCDH. Second, while de Chaisemartin (2010) impose a parallel trends assumption on unexposed outcomes, Blundell and Dias (2009) impose a parallel trends assumption on untreated outcomes. Following de Chaisemartin (2010), Hudson et al. (2017) also study $2\times 2$ DID-IV designs with a non-binary, ordered treatment. [^4] Recently, Chen et al. (2025) extend our DID-IV framework by allowing for covariates. Helmers et al. (2025) consider an extension to the continuous treatment case. Building on de Chaisemartin (2010), however, de Chaisemartin and D’Haultfœuille (2018) formalize $2\times 2$ DID-IV designs differently, calling them Fuzzy DID designs.
In this paper, we first formalize $2\times 2$ DID-IV designs, complementing de Chaisemartin (2010) and Hudson et al. (2017). Specifically, while the framework we consider here is mainly based on de Chaisemartin (2010) and Hudson et al. (2017), we explicitly introduce the instrument path into $2\times 2$ DID-IV designs and uncover an additional identifying assumption not given in the previous literature.
Given the identification results and the terminology developed in this paper, we then compare DID-IV to Fuzzy DID. Specifically, we clarify the differences between DID-IV and Fuzzy DID and discuss their implications for treatment adoption behavior and the interpretation of the target parameter.
Second, this paper extends $2\times 2$ DID-IV designs to multiple period settings with the staggered adoption of the instrument across units, which we refer to as staggered DID-IV designs. In practice, empirical researchers often leverage variation in policy adoption timing across units as an instrument for treatment in more than two periods (e.g., Black et al. (2005), Lundborg et al. (2014), and Johnson and Jackson (2019)). However, no previous study has extended $2\times 2$ DID-IV designs to such important settings. In this paper, we formalize the underlying identification strategy as staggered DID-IV designs, and establish the target parameter and the identifying assumptions. Our staggered DID-IV designs allow practitioners to estimate the local average treatment effect even when the treatment adoption is endogenous over time.
Finally, this paper provides a credible estimation method in staggered DID-IV designs under heterogeneous treatment effects. When empirical researchers implicitly rely on the staggered DID-IV designs in practice, they commonly implement this design via TWFEIV regressions (e.g. Johnson and Jackson (2019), Lundborg et al. (2014), Black et al. (2005), Akerman et al. (2015), and Bhuller et al. (2013)). In a companion paper (Miyaji (2024)), however, we show that in more than two periods, the TWFEIV estimand generally fails to summarize the treatment effects if the effect of the instrument on the outcome or on the treatment evolves over time. [^5] See also de Chaisemartin and D’Haultfœuille (2020), who decompose the numerator and denominator of the TWFEIV estimand separately, and point out the issue of interpreting this estimand causally in Section 3.4 of their Web Appendix. Our proposed estimation method would serve as an alternative to the TWFEIV estimator and make DID-IV designs more credible in a given application.
In work related to this paper, de Chaisemartin et al. (2024) also study DID-IV designs building on de Chaisemartin (2010) and Hudson et al. (2017). Our paper is distinct from theirs in three ways. First, they consider continuous treatment with continuous but static instrument, while we consider binary or ordered treatment with binary but dynamic instrument. Second, they consider multiple period settings with potentially non-staggered instrument, while we consider multiple periods settings with staggered instrument. Third, they view Fuzzy DID as a special case of DID-IV, while we formally discuss the relationship between the two designs.
Our paper is also related to the recent DID literature in two ways. First, this paper provides an alternative identification strategy to DID designs, addressing settings in which the parallel trends assumption is unlikely to hold or no credible control group exists. When empirical researchers rely on DID designs and run two-way fixed effects regressions, they often worry that the parallel trends assumption is implausible in practice. To address this concern, they often turn to TWFEIV regressions, exploiting the timing variation of a policy shock as an instrument for treatment (e.g., Miller and Segal (2019)). In this paper, we formalize the underlying identification strategy as instrumented difference-in-differences. Our DID-IV design can be viewed as a natural extension of DID designs, in that the timing variation is used to construct an instrument rather than a treatment.
Second, in line with the DID literature, this paper develops a reliable estimation method for DID-IV designs in the presence of heterogeneous treatment effects. Recently, several studies have pointed out the issue of implementing DID designs via two-way fixed effects regressions, or its dynamic specifications under heterogeneous treatment effects (Athey and Imbens (2022); Borusyak et al. (2024); Callaway and Sant’Anna (2021); de Chaisemartin and D’Haultfœuille (2020); Goodman-Bacon (2021); Imai and Kim (2021); Sun and Abraham (2021); Callaway et al. (2024)). Some of these studies propose credible estimation methods that deliver a sensible estimand and are robust to treatment heterogeneity. In the same spirit as the recent DID literature, this paper proposes an alternative to the TWFEIV regression and illustrates its usefulness through an empirical application.
2  DID-IV in two time periods
In this section, we formalize an instrumented difference-in-differences (DID-IV) in two-period/two-group settings. Specifically, we establish the target parameter and the identifying assumptions in this design. At the end of this section, we also discuss the connections between DID-IV and Fuzzy DID proposed by de Chaisemartin and D’Haultfœuille (2018).
2.1  Set up
We introduce the notation we use throughout Section 2. We consider a panel data setting with two periods and $N$ units. For any random variable $R$, we denote $\mathcal{S}(R)$ to be its support. For each $i\in\{1,\dots,N\}$ and $t\in\{0,1\}$, let $Y_{i,t}$ denote the outcome, and $D_{i,t}\in\{0,1\}$ denote the treatment status: $D_{i,t}=1$ if unit $i$ receives the treatment in period $t$ and $D_{i,t}=0$ if unit $i$ does not receive the treatment. Let $Z_{i,t}\in\{0,1\}$ denote the instrument status: $Z_{i,t}=1$ if unit $i$ is exposed to the instrument in period $t$ and $Z_{i,t}=0$ if unit $i$ is not exposed to the instrument in period $t$. Throughout Section 2, we assume that $\{Y_{i,0},Y_{i,1},D_{i,0},D_{i,1},Z_{i,0},Z_{i,1}\}_{i=1}^{N}$ are independent and identically distributed (i.i.d).
We introduce the path of the treatment and the instrument. Let $D_{i}=(D_{i,0},D_{i,1})$ be the treatment path and $Z_{i}=(Z_{i,0},Z_{i,1})$ the instrument path. We assume no one is exposed to the instrument in period $t=0$: $Z_{i,0}=0$ for all $i$; we refer to this as a sharp assignment of the instrument. We denote $E_{i}\in\{0,1\}$ as the group variable: $E_{i}=1$ if unit $i$ is exposed to the instrument in period $t=1$ (exposed group) and $E_{i}=0$ if unit $i$ is not exposed to the instrument in period $t=1$ (unexposed group). In contrast to the sharp assignment of the instrument, we allow the general adoption process of the treatment; that is, we assume the treatment path can take four values with non-zero probability: $\{(0,0),(0,1),(1,0),(1,1)\}\in\mathcal{S}(D)$.
In practice, researchers are interested in the effect of a treatment $D_{i,t}$ on an outcome $Y_{i,t}$, and the instrument $Z_{i,1}$ typically represents a policy shock that encourages people to adopt the treatment in period $t=1$. For instance, Duflo (2001) estimates returns to schooling in Indonesia, exploiting variation arising from a new school construction program across regions and cohorts as an instrument for education attainment.
Next, we introduce the potential outcomes framework. Let $Y_{i,t}(d,z)$ denote the potential outcome in period $t$ when unit $i$ receives the treatment path $d\in\mathcal{S}(D)$ and the instrument path $z\in\mathcal{S}(Z)$. Similarly, let $D_{i,t}(z)$ denote the potential treatment status in period $t$ when unit $i$ receives the instrument path $z\in\mathcal{S}(Z)$. We refer to $D_{i,t}((0,0))$ as unexposed treatment and $D_{i,t}((0,1))$ as exposed treatment. Since the treatment and the instrument take only two values, one can write the observed treatments $D_{i,t}$ and outcomes $Y_{i,t}$ as follows.
$D_{i,t}=\sum_{z\in\mathcal{S}(Z)}\mathbf{1}\{Z_{i}=z\}D_{i,t}(z),\hskip 8.53581ptY_{i,t}=\sum_{z\in\mathcal{S}(Z)}\sum_{d\in\mathcal{S}(D)}\mathbf{1}\{Z_{i}=z,D_{i}=d\}Y_{i,t}(d,z).$
We make a no carryover assumption on the potential outcomes $Y_{i,t}(d,z)$.
Assumption 1 (No carryover assumption).
$\forall z\in\mathcal{S}(Z),\forall d\in\mathcal{S}(D),Y_{i,0}(d,z)=Y_{i,0}(d_{0},z),Y_{i,1}(d,z)=Y_{i,1}(d_{1},z),$
where $d=(d_{0},d_{1})$ is the generic element of the treatment path $D_{i}$.
Assumption 1 states that potential outcomes $Y_{i,t}(d,z)$ depend only on the current treatment status $d_{t}$ and the instrument path $z$. In the DID literature, several studies adopt this assumption in non-staggered treatment settings (e.g. de Chaisemartin and D’Haultfœuille (2020); Imai and Kim (2021)).
Next, we introduce the group variable $G_{i}^{Z}\equiv(D_{i,1}((0,0)),D_{i,1}((0,1)))$ that describes the type of unit $i$ according to the response of $D_{i,1}$ on the instrument path $z$. Following the terminology in Imbens and Angrist (1994), we define $G_{i}^{Z}=(0,0)\equiv NT^{Z}$ as the never-takers, $G_{i}^{Z}=(1,1)\equiv AT^{Z}$ as the always-takers, $G_{i}^{Z}=(0,1)\equiv CM^{Z}$ as the compliers, and $G_{i}^{Z}=(1,0)\equiv DF^{Z}$ as the defiers.
Henceforth, we keep Assumption 1. In the next subsection, we define the target parameter in $2\times 2$ DID-IV designs.
2.2  Target parameter in $2\times 2$ DID-IV designs
In $2\times 2$ DID-IV designs, our target parameter is the local average treatment effect on the treated (LATET) in period $t=1$ defined below.
Definition.
The local average treatment effect on the treated (LATET) in period $t=1$ is
$LATET$ $\equiv E[Y_{i,1}(1)-Y_{i,1}(0)|E_{i}=1,D_{i,1}((0,1))>D_{i,1}((0,0))]$
$=E[Y_{i,1}(1)-Y_{i,1}(0)|E_{i}=1,CM^{Z}].$
This parameter measures the treatment effects in period $1$ for units that belong to the exposed group ($E_{i}=1$) and are induced to receive the treatment by the instrument in that period. In the DID-IV literature, de Chaisemartin (2010) considers the same target parameter, while Hudson et al. (2017) define the local average treatment effect (LATE) in period t=1t=1—unconditional on $E_{i}$—as their target parameter. The LATET has been also proposed in heterogeneous effects IV models with binary instrument (e.g. Słoczyński (2020); Słoczyński et al. (2022)).
We focus on the LATET for two reasons. First, this parameter would be particularly of interest if the instrument reflects a policy change of interest to the researcher (Heckman and Vytlacil (2001), J. and Vytlacil,Edward (2005)). Second, the LATET is a natural extension of the target parameter in DID designs, the so-called average treatment effects on the treated (ATT); both causal parameters measure the treatment effects among the units affected by a policy shock (the treatment in DID designs) and belonging to an exposed group (treatment group).
2.3  Identification assumptions in $2\times 2$ DID-IV designs
This subsection formalizes the identification assumptions in $2\times 2$ DID-IV designs. In two periods and two groups settings, a popular estimand is the ratio between the DID estimand of the outcome and the DID estimand of the treatment (Duflo (2001), Field (2007)):
$w_{DID}$ $=\frac{E[Y_{i,1}-Y_{i,0}|E_{i}=1]-E[Y_{i,1}-Y_{i,0}|E_{i}=0]}{E[D_{i,1}-D_{i,0}|E_{i}=1]-E[D_{i,1}-D_{i,0}|E_{i}=0]}.$
Following the terminology in de Chaisemartin and D’Haultfœuille (2018), we call this the Wald-DID estimand.
We consider the following identifying assumptions for the Wald-DID estimand to capture the LATET.
Assumption 2 (Exclusion restriction for potential outcomes).
$\forall z\in\mathcal{S}(Z),\forall d\in\mathcal{S}(D),\forall t\in\{0,1\},Y_{i,t}(d,z)=Y_{i,t}(d).$
This assumption requires that the instrument path does not directly affect potential outcomes other than through treatment. This assumption is common in the IV literature; see e.g., Imbens and Angrist (1994) and Abadie (2003). In the DID-IV literature, de Chaisemartin (2010) and Hudson et al. (2017) impose a similar assumption without introducing the instrument path.
Given Assumptions 1-2, the observed outcome $Y_{i,t}$ can be written as
$Y_{i,t}=D_{i,t}Y_{i,t}(1)+(1-D_{i,t})Y_{i,t}(0).$
Assumptions 1-2 also allow us to introduce the notions of exposed and unexposed outcomes. For any $z\in\mathcal{S}(Z)$, let $Y_{i,t}(D_{i,t}(z))$ denote the outcome if the instrument path is $z$:
$Y_{i,t}(D_{i,t}(z))\equiv D_{i,t}(z)Y_{i,t}(1)+(1-D_{i,t}(z))Y_{i,t}(0).$
Hereafter, we refer to $Y_{i,t}(D_{i,t}((0,0)))$ as the unexposed outcomes and $Y_{i,t}(D_{i,t}((0,1)))$ as the exposed outcomes. [^6] Exposed and unexposed outcomes are not new concepts in econometrics. For example, in IV designs, the numerator of the Wald estimand compares expected outcomes across instrument values: E[Y|Z=1]E[Y|Z=0]E[Y|Z=1]-E[Y|Z=0]. This difference can be written as E[Y(D(1))|Z=1]E[Y(D(0))|Z=0]E[Y(D(1))|Z=1]-E[Y(D(0))|Z=0], which corresponds to a comparison between exposed and unexposed outcomes.
Next, we make the following monotonicity assumption as in Imbens and Angrist (1994).
Assumption 3 (Monotonicity assumption at period $t=1$).
$Pr(D_{i,1}((0,1))\geq D_{i,1}((0,0)))=1\hskip 8.53581pt\text{or}\hskip 5.69054ptPr(D_{i,1}((0,1))\leq D_{i,1}((0,0)))=1.$
This assumption requires that the instrument path affects the treatment choice at period $t=1$ in a monotone (uniform) way. It implies that the group variable $G_{i}^{Z}$ can take three values with non-zero probability. In the DID-IV literature, de Chaisemartin (2010) and Hudson et al. (2017) make the same assumption. Hereafter, we consider the type of monotonicity assumption that rules out the existence of the defiers $DF^{Z}$.
Assumption 4 (No anticipation in the first stage).
$D_{i,0}((0,1))=D_{i,0}((0,0))\hskip 5.69054pta.s.\hskip 8.53581pt\text{for all units $i$ with $E_{i}=1$}.$
Assumption 4 requires that the potential treatment choice before the exposure to instrument is equal to the baseline treatment choice $D_{i,0}((0,0))$ in an exposed group. [^7] In the DID literature, recent studies impose the no anticipation assumption on untreated potential outcomes in several ways. Callaway and Sant’Anna (2021) and Sun and Abraham (2021) assume the average version of the no anticipation assumption, whereas Athey and Imbens (2022) assume it for all units $i$. Roth et al. (2023) take the intermediate approach: they adopt the no anticipation assumption for the treated units. The no anticipation assumption on potential treatment choices in Assumption 4 is in line with that of Roth et al. (2023). [^8] This assumption is not provided in the previous DID-IV literature: de Chaisemartin (2010) and Hudson et al. (2017) implicitly impose this assumption by writing observed treatment choice in period $t=0$ as $D_{i,0}(0)$. This assumption restricts the anticipatory behavior and would be plausible if the instrument path is ex ante not known for all the units in an exposed group.
Next, we impose the parallel trends assumptions in the treatment and the outcome. de Chaisemartin (2010) and Hudson et al. (2017) make the similar assumptions.
Assumption 5 (Parallel Trends Assumption in the treatment).
$E[D_{i,1}((0,0))-D_{i,0}((0,0))|E_{i}=0]=E[D_{i,1}((0,0))-D_{i,0}((0,0))|E_{i}=1].$
Assumption 5 is a parallel trends assumption in the treatment. This assumption requires that the expectation of the treatment between exposed and unexposed groups would have followed the same path if the assignment of the instrument had not occurred. For instance, in Duflo (2001), this assumption requires that the evolution of mean education attainment would have been the same between exposed and unexposed groups if the policy shock had not occurred during the two periods.
Assumption 6 (Parallel Trends Assumption in the outcome).
$E[Y_{i,1}(D_{i,1}((0,0)))-Y_{i,0}(D_{i,0}((0,0)))|E_{i}=0]$
$=$ $E[Y_{i,1}(D_{i,1}((0,0)))-Y_{i,0}(D_{i,0}((0,0)))|E_{i}=1].$
Assumption 6 is a parallel trends assumption in the outcome, requiring that the evolution of the unexposed outcome is, on average, the same between exposed and unexposed groups.999de Chaisemartin et al. (2024) note that this assumption imposes restrictions on treatment effect heterogeneity when the standard parallel trends assumption holds between exposed and unexposed groups. For instance, in Duflo (2001), this assumption requires that the expectation of log annual earnings would have followed the same path from period 0 to period $1$ between exposed and unexposed groups in the absence of a policy shock.
When empirical researchers exploit variation arising from a policy shock as an instrument for treatment and use the Wald-DID estimand, they often refer to Assumptions 5 and 6. For instance, Duflo (2001) estimates returns to schooling in Indonesia, relying on “the identification assumption that the evolution of wages and education across cohorts would not have varied systematically from one region to another in the absence of the program”.
Finally, we assume a relevance condition. This assumption guarantees that the Wald-DID estimand $w_{DID}$ is well defined.
Assumption 7 (Relevance condition).
$E[D_{i,1}-D_{i,0}|E_{i}=1]-E[D_{i,1}-D_{i,0}|E_{i}=0]>0.$
The theorem below shows that if Assumptions 1-7 hold, the Wald-DID estimand captures the LATET in period $1$.
Theorem 1.
If Assumptions 1-7 hold, the Wald-DID estimand $w_{DID}$ is equal to the LATET in period $t=1$; that is,
$w_{DID}=E[Y_{i,1}(1)-Y_{i,1}(0)|E_{i}=1,CM^{Z}].$
holds.
Proof.
See Appendix. ∎
2.4  Comparing DID-IV with Fuzzy DID
de Chaisemartin and D’Haultfœuille (2018) (henceforth, “dCDH”) also investigate the identifying assumptions for the Wald-DID estimand to capture the causal effects under the same setting considered here.101010dCDH also consider the DID-IV setting as in this paper. This follows from two observations. First, the group variable $G$ is included in their treatment participation equation $D=\mathbf{1}\{V\geq v_{GT}\}$ (see Assumption 3 in dCDH), implying that $G$ plays the role of the instrument $Z$ in our framework. Second, dCDH also assume the sharp assignment of the instrument: in their treatment participation equation, they impose $v_{10}=v_{00}$ (see page $1019$ of dCDH). However, dCDH formalize $2\times 2$ DID-IV designs differently, calling them Fuzzy DID designs. Moreover, dCDH point out that the Wald-DID estimand requires the stable treatment effect assumption to identify their target parameter.
In this section, we clarify the differences between this paper and dCDH, and provide the implications of these differences for treatment adoption behavior, the interpretation of the target parameter, and the use of the Wald-DID estimand. The detailed discussions and proofs are provided in Online Appendix D.
Since dCDH formalize Fuzzy DID designs under repeated cross section data, we first introduce notation for the repeated cross section setting. Let $Y$ and $D$ denote the outcome and the treatment, respectively. Let $Z\in\{0,1\}$ denote the group indicator, where $Z=1$ indicates the exposed group. Let $T\in\{0,1\}$ denote the time indicator. Let $Y(0),Y(1)$ and $D(0),D(1)$ denote the potential outcomes and the potential treatment choices, respectively. [^11] Here, we impose the no carryover assumption and the exclusion restriction on potential outcomes, and the no anticipation assumption on potential treatment choices, as in dCDH. Let $D_{t}(0)$ and $D_{t}(1)$ denote the potential treatment choices in period $T=t$ when the group indicator takes value $Z=0$ and $Z=1$, respectively.121212dCDH introduce the treatment participation equation $D=\mathbf{1}\{V\geq v_{ZT}\}$ and define the potential treatment choice in period $t$ as $D(t)=\mathbf{1}\{V\geq v_{Zt}\}$. In this paper, we instead adopt the notation $D_{t}(Z)=\mathbf{1}\{V\geq v_{Zt}\}$ to facilitate interpretation. In dCDH, the group variable $G$ plays the role of the instrument $Z$ in this paper. Accordingly, we relabel $G$ as $Z$ in the following discussion.
The main difference between dCDH and this paper lies in the definition of the target parameter: we focus on the LATET, while dCDH focus on the switcher local average treatment effect on the treated (SLATET) defined below.
Definition.
The switcher local average treatment effect on the treated (SLATET) is
$SLATET$ $\equiv E[Y(1)-Y(0)|Z=1,T=1,D_{0}(0)<D_{1}(1)]$
$=E[Y(1)-Y(0)|Z=1,T=1,SW],$
where we use the restriction $v_{10}=v_{00}$ in dCDH, which implies $D_{0}(1)=D_{0}(0)$.
This parameter measures the treatment effects for units who belong to the exposed group ($Z=1$) and switch into treatment at time $T=1$ (switchers, SW).
Because the target parameters differ, the identifying assumptions underlying Fuzzy DID designs also differ from those in DID-IV designs, though there are similarities as well. [^13] Specifically, dCDH also impose Assumptions 1-4 and 7 in this paper. The main differences are as follows. First, dCDH assume the stable treatment rate assumption in an unexposed group (see Assumption 22 in dCDH), whereas we assume the parallel trends assumption in the treatment (Assumption 5). Second, dCDH assume the treatment participation equation (see Assumption 33 in dCDH), which imposes the monotonicity with respect to time $T$ in addition to the monotonicity with respect to instrument $Z$ (which corresponds to Assumption 3 in this paper). Finally, while we impose a parallel trends (PT) assumption on unexposed outcomes, dCDH impose a PT assumption on untreated outcomes (see Assumption $4$ in dCDH).
In Online Appendix D, we first examine how these differences imply different restrictions on treatment adoption behavior across the two designs. We begin by describing the heterogeneity in treatment adoption behavior under the setting considered here. Specifically, we introduce the group variable $G^{T}=(D_{0}(0),D_{1}(0))$, which characterizes treatment adoption over time when the instrument is $Z=0$ in the second period. We define $G^{T}=(0,0)\equiv NT^{T}$ as time never-takers, $G^{T}=(1,1)\equiv AT^{T}$ as time always-takers, $G^{T}=(0,1)\equiv CM^{T}$ as time compliers, and $G^{T}=(1,0)\equiv DF^{T}$ as time defiers. Using the group variables $G^{Z}=(D_{1}(0),D_{1}(1))$ (which was introduced in Section 2.1 for the panel data case) and $G^{T}$, we then partition units into eight types within each group, as summarized in Tables 1-2.
Next, we investigate which types are excluded by the identifying assumptions under DID-IV and Fuzzy DID, respectively. Specifically, Tables 12 and Tables 34 summarize all latent treatment adoption types under DID-IV and Fuzzy DID respectively, where the types painted in gray color are excluded by the identifying assumptions of each design. Here, in Table 3, we exclude the $DF^{Z}$ and the $DF^{T}$ by the monotonicity assumptions with respect to instrument $Z$ and time $T$, which are implied by the treatment participation equation (see Lemma 1 in Online Appendix D.2.3).
Table 1: Exposed group ($z=1$)
observed counterfactual
[…]
$D_{0}(0)=1,D_{1}(1)=1$ $AT^{Z}\land AT^{T}$ $CM^{Z}\land DF^{T}$
$D_{0}(0)=1,D_{1}(1)=0$ $DF^{Z}\land AT^{T}$ $NT^{Z}\land DF^{T}$
$D_{0}(0)=0,D_{1}(1)=1$ $AT^{Z}\land CM^{T}$ $CM^{Z}\land NT^{T}$
$D_{0}(0)=0,D_{1}(1)=0$ $DF^{Z}\land CM^{T}$ $NT^{Z}\land NT^{T}$
Table 2: Unexposed group ($z=0$)
observed counterfactual
[…]
$D_{0}(0)=1,D_{1}(0)=1$ $AT^{Z}\land AT^{T}$ $DF^{Z}\land AT^{T}$
$D_{0}(0)=1,D_{1}(0)=0$ $CM^{Z}\land DF^{T}$ $NT^{Z}\land DF^{T}$
$D_{0}(0)=0,D_{1}(0)=1$ $AT^{Z}\land CM^{T}$ $DF^{Z}\land CM^{T}$
$D_{0}(0)=0,D_{1}(0)=0$ $CM^{Z}\land NT^{T}$ $NT^{Z}\land NT^{T}$
Notes: These tables represent mutually exclusive and exhaustive types under DID-IV designs. The types painted in gray color are excluded by the monotonicity assumption with respect to instrument $Z$.
Table 3: Exposed group ($z=1$)
observed counterfactual
[…]
$D_{0}(0)=1,D_{1}(1)=1$ $AT^{Z}\land AT^{T}$ $CM^{Z}\land DF^{T}$
$D_{0}(0)=1,D_{1}(1)=0$ $DF^{Z}\land AT^{T}$ $NT^{Z}\land DF^{T}$
$D_{0}(0)=0,D_{1}(1)=1$ $AT^{Z}\land CM^{T}$ $CM^{Z}\land NT^{T}$
$D_{0}(0)=0,D_{1}(1)=0$ $DF^{Z}\land CM^{T}$ $NT^{Z}\land NT^{T}$
Table 4: Unexposed group ($z=0$)
observed counterfactual
[…]
$D_{0}(0)=1,D_{1}(0)=1$ $AT^{Z}\land AT^{T}$ $DF^{Z}\land AT^{T}$
$D_{0}(0)=1,D_{1}(0)=0$ $CM^{Z}\land DF^{T}$ $NT^{Z}\land DF^{T}$
$D_{0}(0)=0,D_{1}(0)=1$ $AT^{Z}\land CM^{T}$ $DF^{Z}\land CM^{T}$
$D_{0}(0)=0,D_{1}(0)=0$ $CM^{Z}\land NT^{T}$ $NT^{Z}\land NT^{T}$
Notes: These tables represent mutually exclusive and exhaustive types under Fuzzy DID designs. The types painted in gray color are excluded by the identifying assumptions in dCDH.
By comparing Tables 12 with Tables 34, we obtain two implications. First, the restrictions imposed by Fuzzy DID designs are stronger than those imposed by DID-IV designs. Second, while the restrictions under DID-IV designs are symmetric across exposed and unexposed groups, those under Fuzzy DID designs are asymmetric.
Building on Tables 34, we next clarify the difference in target parameter between the two designs. Specifically, we show that dCDH’s target parameter, the SLATET, can be expressed as a weighted average of two distinct causal parameters. One parameter captures the treatment effects for units of type $CM^{Z}\land NT^{T}$, a subpopulation of the compliers $CM^{Z}$, while the other captures the treatment effects among the type $AT^{Z}\land CM^{T}$ (case (i)).
Importantly, this decomposition depends on the direction of the two monotonicity assumptions. If $DF^{Z}$ and $CM^{T}$ are excluded by the two monotonicity conditions in dCDH, the SLATET identifies the treatment effects for units of type $CM^{Z}\land NT^{T}$ (case (ii)). If $CM^{Z}$ and $DF^{T}$ are excluded, the SLATET identifies the treatment effects for units of type $AT^{Z}\land CM^{T}$ (case (iii)), as formally established in Theorem 6 in Appendix D.3.
This decomposition result has several important implications. First, empirical researchers should specify the direction of the two monotonicity conditions ex ante if they wish to understand which latent type’s causal effect is identified by the SLATET. Second, the SLATET may fail to be the policy-relevant parameter (Heckman and Vytlacil (2001)), as it is generally contaminated by treatment effect for units whose treatment status is affected by time but not affected by the instrument (see cases (i) and (iii)). Finally, even when the two monotonicity conditions ensure that the SLATET is policy relevant, it identifies the treatment effects for a narrower population than the LATET (see case (ii)).
Finally, we explain why the role of the Wald–DID estimand differs between Fuzzy DID and DID-IV designs. In DID-IV, we view the Wald–DID estimand as a natural estimand for identifying the LATET. By contrast, dCDH point out that, under Fuzzy DID, the Wald-DID estimand requires the stable treatment effect assumption to identify the SLATET.141414dCDH also show that the Wald–DID estimand additionally requires homogeneity of the SLATET between the exposed and unexposed groups in order to identify the SLATET in the exposed group when the stable treatment rate condition in the unexposed group is not satisfied. In earlier work, Blundell and Dias (2009) make a similar argument in Subsection F.2 of Section III. In Remark 4 of Section D.3 in Online Appendix D, however, we show that this homogeneity condition cannot be straightforwardly interpreted as requiring homogeneous treatment effects between the two groups. We demonstrate that this difference arises because Fuzzy DID and DID-IV rely on different types of the PT assumption. Given this, we also discuss which PT assumption is more suitable for DID-IV settings.
Specifically, we first show that the Wald-DID estimand requires the stable treatment effect assumption under Fuzzy DID because dCDH impose the PT assumption on untreated outcomes. Recall that in DID-IV settings, units are allowed to adopt the treatment without instrument during the two periods. As a result, the PT assumption on untreated outcomes is insufficient to capture the average time trends of the outcome even in the unexposed group. We show that this fact leads dCDH to impose the stable treatment effect assumption for the Wald-DID estimand to identify the SLATET. By contrast, this issue does not arise in DID-IV designs because we impose the PT assumption on unexposed outcomes, which is sufficient to capture the average time trends of the outcome for the unexposed group.
Next, we argue that the PT assumption on unexposed outcomes is more suitable for DID-IV settings for three reasons. First, the PT assumption on unexposed outcomes is more consistent with the source of identifying variation in DID-IV settings than the PT assumption on untreated potential outcomes. This is because the latter relies on variation in treatment, while the former exploits variation in instrument. Second, in most applications of DID-IV methods, we cannot impute the untreated potential outcomes in general (e.g., Duflo (2001), Black et al. (2005)). For instance, in Duflo (2001), the PT assumption on untreated outcomes would require the data to include units with zero educational attainment, which is unrealistic in practice. Finally, in DID-IV settings, while the PT assumption on unexposed outcomes is indirectly testable (see Section 5.3 in this paper), the PT assumption on untreated outcomes is difficult to assess using pre-exposed period data, as some units may already adopt the treatment before period 0.
We conclude this section by providing guidance for empirical researchers choosing between DID-IV and Fuzzy DID designs in practice. First, when the goal is to identify a policy-relevant parameter, DID-IV designs are more attractive. Under Fuzzy DID designs, the target parameter may not be policy relevant in general. Second, when researchers are interested in identifying the SLATET, they should carefully assess the plausibility of the restrictions on treatment adoption behavior imposed by Fuzzy DID designs. If these restrictions appear questionable, researchers may instead target the LATET and adopt DID-IV designs, under which the only restriction on treatment adoption behavior is monotonicity with respect to the instrument. Finally, if researchers wish to assess the plausibility of the PT assumption, DID-IV designs are more appealing because the PT assumption imposed under Fuzzy DID is generally difficult to test using the pre-exposed data.
3  DID-IV in multiple time periods
We now extend the $2\times 2$ DID-IV design to multiple period settings with the staggered adoption of the instrument across units (Black et al. (2005); Bhuller et al. (2013) Lundborg et al. (2014); and Meghir et al. (2018)). We call it a staggered DID-IV design, and establish the target parameter and identifying assumptions.
3.1  Set up
We introduce the notation we use throughout Section 3 to Section 5. We consider a panel data setting with $T$ periods and $N$ units. For each $i\in\{1,\dots N\}$ and $t\in\{1,\dots,T\}$, let $Y_{i,t}$ denote the outcome, $D_{i,t}\in\{0,1\}$ denote the treatment status, and $Z_{i,t}\in\{0,1\}$ denote the instrument status. Let $D_{i}=(D_{i,1},\dots,D_{i,T})$ and $Z_{i}=(Z_{i,1},\dots,Z_{i,T})$ denote the path of the treatment and the instrument for unit $i$, respectively. Throughout Section 3 to Section 5, we assume that $\{Y_{i,t},D_{i,t},Z_{i,t}\}_{t=1}^{T}$ are i.i.d.
We make the following assumption about the assignment process of the instrument.
Assumption 8 (Staggered adoption for $Z_{i,t}$).
$Z_{i,1}=0$ for all $i$. For $s<t$, $Z_{i,s}\leq Z_{i,t}$ where $s,t\in\{1,\dots T\}$.
Assumption 8 requires that no one is exposed to the instrument in time $t=1$ and once units start getting exposed to the instrument, units remain exposed to that instrument. In the DID literature, several studies make a similar assumption for the treatment and call it the “staggered treatment adoption” (e.g. Athey and Imbens (2022); Callaway and Sant’Anna (2021); and Sun and Abraham (2021)).
Given Assumption 8, one can uniquely characterize one’s instrument path by the initial exposure date of the instrument, denoted as $E_{i}=\min\{t:Z_{i,t}=1\}$. If unit $i$ is not exposed to the instrument for all time periods, we define $E_{i}=\infty$. Based on the initial exposure period $E_{i}$, one can uniquely partition units into mutually exclusive and exhaustive cohorts $e$ for $e\in\{2,3,\dots,T,\infty\}$. Let $E_{i,e}=\mathbf{1}\{E_{i}=e\}$ denote the binary indicator that takes one if unit $i$ belongs to cohort $e$. Let $\bar{e}=\max_{i=1,\dots,n}E_{i}$ denote the largest cohort value in the dataset. Let $\mathcal{E}=\mathcal{S}(E_{i})\setminus\{\bar{e}\}\subseteq\{2,3,\dots,T\}$ denote the support of $E_{i}$ excluding $\bar{e}$.
Similar to the two periods setting in Subsection 2.1, we allow the general adoption process for the treatment: the treatment can potentially turn on/off repeatedly over time. de Chaisemartin and D’Haultfœuille (2020) and Imai and Kim (2021) also consider the same setting in the recent DID literature.
Next, we introduce the potential outcomes framework in multiple time periods. Let $Y_{i,t}(d,z)$ denote the potential outcome in period $t$ when unit $i$ receives the treatment path $d\in\mathcal{S}(D)$ and the instrument path $z\in\mathcal{S}(Z)$. Similarly, let $D_{i,t}(z)$ denote the potential treatment status in period $t$ when unit $i$ receives the instrument path $z\in\mathcal{S}(Z)$.
Assumption 8 allows us to rewrite $D_{i,t}(z)$ by the initial exposure date $E_{i}=e$. Let $D_{i,t}^{e}$ denote the potential treatment status in period $t$ if unit $i$ is first exposed to the instrument in period $e$. Let $D_{i,t}^{\infty}$ denote the potential treatment status in period $t$ if unit $i$ is never exposed to the instrument. We call $D_{i,t}^{\infty}$ the “never exposed treatment”. Since the adoption date of the instrument uniquely pins down one’s instrument path, we can write the observed treatment status $D_{i,t}$ for unit $i$ in period $t$ as
$D_{i,t}=D_{i,t}^{\infty}+\sum_{2\leq e\leq T}(D_{i,t}^{e}-D_{i,t}^{\infty})\cdot\mathbf{1}\{E_{i}=e\}.$
We define the effect of an instrument on treatment for unit $i$ in period $t$ as the difference between the observed treatment status to the never exposed treatment status: $D_{i,t}-D_{i,t}^{\infty}$. We refer to $D_{i,t}-D_{i,t}^{\infty}$ as the individual exposed effect in the first stage. [^15] In the DID literature, Callaway and Sant’Anna (2021) and Sun and Abraham (2021) define the effect of a treatment on an outcome in the same fashion.
Next, we introduce the group variable that describes the type of unit $i$ in period $t$, based on the reaction of potential treatment choices in period $t$ to the instrument path $z$. Let $G_{i,e,t}\equiv(D_{i,t}^{\infty},D_{i,t}^{e})$ ($t\geq e$) be the group variable in period $t$ for unit $i$ and the initial exposure date $e$. Following the terminology in section 2.1, we define $G_{i,e,t}=(0,0)\equiv NT_{e,t}$ as the never-takers, $G_{i,e,t}=(1,1)\equiv AT_{e,t}$ as the always-takers, $G_{i,e,t}=(0,1)\equiv CM_{e,t}$ as the compliers and $G_{i,e,t}=(1,0)\equiv DF_{e,t}$ as the defiers in period $t$ and the initial exposure date $e$.
Finally, we make a no carryover assumption on potential outcomes $Y_{i,t}(d,z)$.
Assumption 9 (No carryover assumption in multiple time periods).
$\forall z\in\mathcal{S}(Z),\forall d\in\mathcal{S}(D),\forall t\in\{1,\dots,T\},Y_{i,t}(d,z)=Y_{i,t}(d_{t},z).$
Henceforth, we keep Assumption 8 and Assumption 9. In the next section, we define the target parameter in staggered DID-IV designs.
3.2  Target parameter in staggered DID-IV designs
In staggered DID-IV designs, our target parameter is the cohort specific local average treatment effect on the treated (CLATT) defined below.
Definition.
The cohort specific local average treatment effect on the treated (CLATT) at a given relative period $l$ from the initial adoption of the instrument is
$CLATT_{e,e+l}$ $=E[Y_{i,e+l}(1)-Y_{i,e+l}(0)|E_{i}=e,D_{i,e+l}^{e}>D_{i,e+l}^{\infty}]$
$=E[Y_{i,e+l}(1)-Y_{i,e+l}(0)|E_{i}=e,CM_{e,e+l}].$
Each CLATT is a natural generalization of the LATET in Subsection 2.2 and suitable for the setting of the staggered instrument adoption. This parameter measures the treatment effects at a given relative period $l$ from the initial exposure date $E_{i}=e$, for those who belong to cohort $e$, and are the compliers $CM_{e,e+l}$ who are induced to treatment by instrument in period $e+l$. Each CLATT can potentially vary across cohorts and over time because it depends on cohort $e$, relative period $l$, and the compliers $CM_{e,e+l}$.
3.3  Identification assumptions in staggered DID-IV designs
In this subsection, we establish the identification assumptions in staggered DID-IV designs.
In staggered DID-IV designs, we first consider the following estimand to identify each $CLATT_{e,e+l}$:
$w^{DID}_{e,l}=\frac{E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=e]-(E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=\infty])}{E[D_{i,e+l}-D_{i,e-1}|E_{i}=e]-(E[D_{i,e+l}-D_{i,e-1}|E_{i}=\infty])},$
for $e\in\mathcal{E}$ and $l\in\{0,\dots,T-e\}$. Note that this estimand is the Wald-DID estimand, where the pre-exposed period is e1e-1 and the control group is the never exposed cohort. Here, we assume that the largest cohort value in the dataset is $\bar{e}=\infty$.
We consider the following identification assumptions for each $w^{DID}_{e,l}$ to capture the $CLATT_{e,e+l}$.
Assumption 10 (Exclusion restriction in multiple time periods).
$\forall z\in\mathcal{S}(Z),\forall d\in\mathcal{S}(D),\forall t\in\{1,\dots,T\},Y_{i,t}(d,z)=Y_{i,t}(d).$
Assumption 10 extends the exclusion restriction in two time periods (Assumption 2) to multiple period settings. This assumption requires that the instrument path does not directly affect the potential outcome for all time periods and its effects are only through treatment.
Given Assumption 9 and Assumption 10, we can write the potential outcome $Y_{i,t}(d,z)$ as $Y_{i,t}(d_{t})=D_{i,t}Y_{i,t}(1)+(1-D_{i,t})Y_{i,t}(0)$. Following Subsection 2.3, we introduce the potential outcomes in period $t$ if unit $i$ is assigned to the instrument path $z\in\mathcal{S}(Z)$.
$Y_{i,t}(D_{i,t}(z))\equiv D_{i,t}(z)Y_{i,t}(1)+(1-D_{i,t}(z))Y_{i,t}(0).$
Since the initial exposure date $E_{i}$ completely characterizes the instrument path, we can write the potential outcomes for cohort $e$ and cohort $\infty$ as $Y_{i,t}(D_{i,t}^{e})$ and $Y_{i,t}(D_{i,t}^{\infty})$, respectively. The potential outcome $Y_{i,t}(D_{i,t}^{e})$ represents the outcome status in period $t$ if unit $i$ is first exposed to the instrument in period $e$ and the potential outcome $Y_{i,t}(D_{i,t}^{\infty})$ represents the outcome status in period $t$ if unit $i$ is never exposed. We refer to $Y_{i,t}(D_{i,t}^{\infty})$ as the “never exposed outcome”.
Assumption 11 (Monotonicity assumption in multiple time periods).
$Pr(D_{i,e+l}^{e}\geq D_{i,e+l}^{\infty})=1\hskip 5.69054pt\text{or}\hskip 5.69054ptPr(D_{i,e+l}^{e}\leq D_{i,e+l}^{\infty})=1\hskip 5.69054pt\text{for all}\hskip 5.69054pte\in\mathcal{E}\hskip 5.69054pt\text{and all}\hskip 5.69054ptl\geq 0.$
This assumption requires that the instrument path affects the treatment adoption behavior in a monotone way for all relative periods after the initial exposure date $E_{i}=e$. Recall that we define $D_{i,t}-D_{i,t}^{\infty}$ to be the effect of an instrument on treatment for unit $i$ in period $t$. Assumption 11 requires that the individual exposed effect in the first stage should be non-negative or non-positive during the periods after the initial exposure to the instrument for all $i$. This assumption implies that the group variable $G_{i,e,t}\equiv(D_{i,t}^{\infty},D_{i,t}^{e})$ can take three values with non-zero probability for all $e\in\mathcal{E}$ and all $t\geq e$. Hereafter, we consider the type of the monotonicity assumption that rules out the existence of the defiers $DF_{e,t}$ for all $t\geq e$ in any cohort $e\in\mathcal{E}$.
Assumption 12 (No anticipation in the first stage).
$D_{i,e+l}^{e}=D_{i,e+l}^{\infty}\hskip 8.53581pt\text{for all}\hskip 5.69054pte\in\mathcal{E}\hskip 5.69054pt\text{and all}\hskip 5.69054ptl<0.$
Assumption 12 requires that potential treatment choices in any $l$ period before the initial exposure to the instrument is equal to the never exposed treatment. This assumption is a natural generalization of Assumption 4 to multiple period settings and restricts the anticipatory behavior before the initial exposure to the instrument.
Assumption 13 (Parallel trends assumption in the treatment based on a never exposed cohort).
$\text{For each}\hskip 5.69054pte\in\mathcal{E}\hskip 5.69054pt\text{and}\hskip 5.69054ptt\in\{2,\dots,T\}\hskip 5.69054pt\text{such that}\hskip 5.69054ptt\geq e,$
$E[D_{i,t}^{\infty}-D_{i,t-1}^{\infty}|E_{i}=e]=E[D_{i,t}^{\infty}-D_{i,t-1}^{\infty}|E_{i}=\infty].$
Assumption 13 is a parallel trends assumption in the treatment based on the never exposed cohort. It requires that, in the absence of exposure to the instrument, the average evolution of the treatment would have followed the same path between cohort $e$ and the never exposed cohort $\infty$. Assumption 13 is analogous to that of Callaway and Sant’Anna (2021) and Sun and Abraham (2021) in DID designs; these studies impose the same type of parallel trends assumption on untreated potential outcomes.
Assumption 14 (Parallel trends assumption in the outcome based on a never exposed cohort).
$\text{For each}\hskip 5.69054pte\in\mathcal{E}\hskip 5.69054pt\text{and}\hskip 5.69054ptt\in\{2,\dots,T\}\hskip 5.69054pt\text{such that}\hskip 5.69054ptt\geq e,$
$E[Y_{i,t}(D_{i,t}^{\infty})-Y_{i,t-1}(D_{i,t-1}^{\infty})|E_{i}=e]=E[Y_{i,t}(D_{i,t}^{\infty})-Y_{i,t-1}(D_{i,t-1}^{\infty})|E_{i}=\infty].$
Assumption 14 is a parallel trends assumption in the outcome based on the never exposed cohort. It requires that, in the absence of exposure to the instrument, the expected evolution of the outcome under no exposure to the instrument would have been the same, on average, between cohort $e$ and the never exposed cohort $\infty$.
Assumption 15 (Relevance condition based on a never exposed cohort).
$\text{For each}\hskip 5.69054pte\in\mathcal{E}\hskip 5.69054pt\text{and}\hskip 5.69054ptl\in\{0,\dots,T-e\},$
$E[D_{i,e+l}-D_{i,e-1}|E_{i}=e]-(E[D_{i,e+l}-D_{i,e-1}|E_{i}=\infty])>0.$
Assumption 15 is a relevance condition in multiple period settings, ensuring that the Wald-DID estimand $w^{DID}_{e,l}$ is well defined.
The following theorem shows that under Assumptions 815, each Wald-DID estimand $w^{DID}_{e,l}$ identifies the corresponding $CLATT_{e,e+l}$.
Theorem 2.
If Assumptions 815 hold, the Wald-DID estimand $w^{DID}_{e,l}$ identifies the corresponding $CLATT_{e,e+l}$:
$w^{DID}_{e,l}=E[Y_{i,e+l}(1)-Y_{i,e+l}(0)|E_{i}=e,CM_{e,e+l}],$
for each $e\in\mathcal{E}$ and $l\in\{0,\dots,T-e\}$.
Proof.
See Appendix. ∎
In some applications, however, there may be no never exposed cohort, or its sample size may be too small to serve as a reliable control group. Moreover, the parallel trends assumptions based on the never exposed cohort may be viewed as less credible in certain empirical settings. In such cases, it is natural to instead use the not-yet-exposed cohorts as the control group. Accordingly, we consider the following alternative Wald-DID estimand:
$w^{DID,ny}_{e,l}=\frac{E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=e]-(E[Y_{i,e+l}-Y_{i,e-1}|Z_{i,e+l}=0,E_{i,e}=0])}{E[D_{i,e+l}-D_{i,e-1}|E_{i}=e]-(E[D_{i,e+l}-D_{i,e-1}|Z_{i,e+l}=0,E_{i,e}=0])},$
for $e\in\mathcal{E}$ and $l\in\{0,\dots,T-e\}$ such that $l<\bar{e}-e$. In this Wald-DID estimand, the control group is the not-yet-exposed cohorts, that is, the cohorts not exposed to the instrument by time $t=e+l$.
If we adopt the above estimand $w^{DID,ny}_{e,l}$, we can replace Assumptions 1315 with Assumptions 1618 below.
Assumption 16 (Parallel trends assumption in the treatment based on not-yet-exposed cohorts).
$\text{For each}\hskip 5.69054pte\in\mathcal{E}\hskip 5.69054pt\text{and}\hskip 5.69054pt\text{each}\hskip 5.69054ptt\in\{2,\dots,T\}\hskip 5.69054pt\text{such that}\hskip 5.69054pte\leq t<\bar{e},$
$E[D_{i,t}^{\infty}-D_{i,t-1}^{\infty}|E_{i}=e]=E[D_{i,t}^{\infty}-D_{i,t-1}^{\infty}|Z_{i,t}=0,E_{i,e}=0].$
Assumption 17 (Parallel trends assumption in the outcome based on not-yet-exposed cohorts).
$\text{For each}\hskip 5.69054pte\in\mathcal{E}\hskip 5.69054pt\text{and}\hskip 5.69054pt\text{each}\hskip 5.69054ptt\in\{2,\dots,T\}\hskip 5.69054pt\text{such that}\hskip 5.69054pte\leq t<\bar{e},$
$E[Y_{i,t}(D_{i,t}^{\infty})-Y_{i,t-1}(D_{i,t-1}^{\infty})|E_{i}=e]=E[Y_{i,t}(D_{i,t}^{\infty})-Y_{i,t-1}(D_{i,t-1}^{\infty})|Z_{i,t}=0,E_{i,e}=0].$
Assumption 18 (Relevance condition based on not-yet-exposed cohorts).
$\text{For each}\hskip 5.69054pte\in\mathcal{E}\hskip 5.69054pt\text{and}\hskip 5.69054ptl\in\{0,\dots,T-e\}\hskip 5.69054pt\text{such that}\hskip 5.69054ptl<\bar{e}-e,$
$E[D_{i,e+l}-D_{i,e-1}|E_{i}=e]-(E[D_{i,e+l}-D_{i,e-1}|Z_{i,e+l}=0,E_{i,e}=0])>0.$
The following theorem shows that under Assumptions 812 and 1618, each Wald-DID estimand $w^{DID,ny}_{e,l}$ identifies the corresponding $CLATT_{e,e+l}$.
Theorem 3.
If Assumptions 812 and 1618 hold, the Wald-DID estimand $w^{DID,ny}_{e,l}$ identifies the corresponding $CLATT_{e,e+l}$:
$w^{DID,ny}_{e,l}=E[Y_{i,e+l}(1)-Y_{i,e+l}(0)|E_{i}=e,CM_{e,e+l}],$
for each $e\in\mathcal{E}$ and $l\in\{0,\dots,T-e\}$ such that $l<\bar{e}-e$.
Proof.
See Appendix. ∎
4  Extensions
This section contains extensions to non-binary, ordered treatments and repeated cross sections. For more details and proofs, see online Appendix Section A.
Non-binary, ordered treatment
Up to now, we have considered only the case of a binary treatment. However the same idea can be applied when treatment takes a finite number of ordered values: $D_{i,t}\in\{0,1,\dots,J\}$. When only two periods exist and treatment is non-binary, our target parameter is the average causal response on the treated (ACRT) defined below.
Definition.
The average causal response on the treated (ACRT) is
$ACRT\equiv\sum_{j=1}^{J}w_{j}\cdot E[Y_{i,1}(j)-Y_{i,1}(j-1)|D_{i,1}((0,1))\geq j>D_{i,1}((0,0)),E_{i}=1].$
where the weight $w_{j}$ is:
$w_{j}=\frac{Pr(D_{i,1}((0,1))\geq j>D_{i,1}((0,0))|E_{i}=1)}{\sum_{j=1}^{J}Pr(D_{i,1}((0,1))\geq j>D_{i,1}((0,0))|E_{i}=1)}.$
The ACRT is a weighted average of the effect of one unit increase in the treatment on the outcome, for those who belong to an exposed group and are induced to increase the treatment in period $t=1$ by instrument. The ACRT is similar to the average causal response (ACR) considered in Angrist and Imbens (1995), with the difference that each weight $w_{j}$ and the associated causal parameter in the ACRT are conditional on $E_{i}=1$.
Online Appendix Section A.1 shows that if we have a non-binary, ordered treatment, the Wald-DID estimand captures the ACRT under the same assumptions in Theorem 1.
With a non-binary, ordered treatment in staggered DID-IV designs, our target parameter is the cohort specific average causal response on the treated (CACRT), a natural generalization of the ACRT. The estimand and the associated identifying assumptions are the same as in Subsection 3.3.
Definition.
The cohort specific average causal response on the treated (CACRT) at a given relative period $l$ from the initial adoption of the instrument is
$CACRT_{e,e+l}\equiv\sum_{j=1}^{J}w^{e}_{e+l,j}\cdot E[Y_{i,e+l}(j)-Y_{i,e+l}(j-1)|E_{i}=e,D_{i,e+l}^{e}\geq j>D_{i,e+l}^{\infty}]$
where the weight $w^{e}_{e+l,j}$ is:
$w^{e}_{e+l,j}=\frac{Pr(D_{i,e+l}^{e}\geq j>D_{i,e+l}^{\infty}|E_{i}=e)}{\sum_{j=1}^{J}Pr(D_{i,e+l}^{e}\geq j>D_{i,e+l}^{\infty}|E_{i}=e)}.$
Repeated cross sections
In some applications, researchers have only access to repeated cross section data. [^16] It also includes the case that researchers use the cross section data, and exploit a policy shock across cohorts as an instrument for treatment as in Duflo (2001). Online Appendix Section A.2 presents the identification assumptions in DID-IV designs under repeated cross section settings.
5  Estimation and inference
In this section, we propose a credible estimation method in staggered DID-IV designs that is robust to treatment effect heterogeneity. First, we propose a simple regression-based method for estimating each CLATT. Following Callaway and Sant’Anna (2021), we then propose a weighting scheme to construct the summary causal parameters from each CLATT. We also discuss the pre-trends tests for checking the plausibility of parallel trends assumptions in DID-IV designs.
Remark 1.
In practice, when researchers implicitly rely on a staggered DID-IV design, they commonly implement this design via two-way fixed instrumental variable (TWFEIV) regressions (e.g., Johnson and Jackson (2019), Lundborg et al. (2014), Black et al. (2005), Akerman et al. (2015), and Bhuller et al. (2013)):
$Y_{i,t}=\phi_{i.}+\lambda_{t.}+\beta_{IV}D_{i,t}+v_{i,t},$
$D_{i,t}=\gamma_{i.}+\zeta_{t.}+\pi Z_{i,t}+\eta_{i,t}.$
In companion paper (Miyaji (2024)), however, we show that in more than two periods, the TWFEIV estimand potentially fails to summarize the treatment effects in the presence of heterogeneous treatment effects. Specifically, we show that the TWFEIV estimand is a weighted average of all possible CLATTs under staggered DID-IV designs, but some weights can be negative if the effect of the instrument on the treatment or the outcome is not stable over time.
5.1  Stacked two stage least squares regression
We propose a regression-based method to consistently estimate our target parameter in staggered DID-IV designs. Our estimation method consists of two steps.
Step 1.
We create the data sets for each $CLATT_{e,e+l}$. Each data set includes the units of time t=e1t=e-1 and $t=e+l$, who are either in cohort $e\in\mathcal{E}$ or in the set of some unexposed cohorts, $U$ ($e\notin U$). If we adopt the parallel trends assumption based on a never-exposed cohort (Assumptions 1314), we can set $U=\{\infty\}$. In this case, we can estimate $CLATT_{e,e+l}$ for all $e\in\mathcal{E}$ and $l\in\{0,\dots,T-e\}$. If we adopt the parallel trends assumption based on not-yet-exposed cohorts (Assumptions 1617), we can set $U=\{e^{\prime}\in\mathcal{S}(E_{i}):e^{\prime}>e+l\}$. In this case, we can estimate $CLATT_{e,e+l}$ for all $e\in\mathcal{E}$ and $l\in\{0,\dots,T-e\}$ such that $l<\bar{e}-e$.
Step 2.
In each data set, we run the following IV regression and obtain the IV estimator $\hat{\beta}^{e,l}_{IV}$:
$Y_{i,t}=\beta^{e,l}_{0}+\beta^{e,l}_{1}\mathbf{1}\{E_{i}=e\}+\beta^{e,l}_{2}\mathbf{1}\{T_{i}=e+l\}+\beta^{e,l}_{IV}D_{i,t}+\epsilon^{e,l}_{i,t}.$
The first stage regression is:
$D_{i,t}=\delta^{e,l}_{0}+\delta^{e,l}_{1}\mathbf{1}\{E_{i}=e\}+\delta^{e,l}_{2}\mathbf{1}\{T_{i}=e+l\}+\delta^{e,l}_{3}\mathbf{1}\{E_{i}=e\}\mathbf{1}\{T_{i}=e+l\}+\eta^{e,l}_{i,t}.$
where the group indicator $\mathbf{1}\{E_{i}=e\}$ and the post-period indicator $\mathbf{1}\{T_{i}=e+l\}$ are the included instruments and the interaction of the two is the excluded instrument.
Formally, our proposed estimator $\widehat{CLATT}_{e,e+l}\equiv\hat{\beta}^{e,l}_{IV}$ takes the following form:
$\widehat{CLATT}_{e,e+l}\equiv\hat{\beta}^{e,l}_{IV}=\frac{\hat{\alpha}^{e,l}}{\hat{\pi}^{e,l}},$
where $\hat{\alpha}^{e,l}$ and $\hat{\pi}^{e,l}$ are:
$\hat{\alpha}^{e,l}$ $=\frac{E_{N}[(Y_{i,e+l}-Y_{i,e-1})\cdot\mathbf{1}\{E_{i}=e\}]}{E_{N}[\mathbf{1}\{E_{i}=e\}]}-\frac{E_{N}[(Y_{i,e+l}-Y_{i,e-1})\cdot\mathbf{1}\{E_{i}\in U\}]}{E_{N}[\mathbf{1}\{E_{i}\in U\}]}$
$\equiv\hat{\alpha}_{e,l}^{1}-\hat{\alpha}_{e,l}^{2}.$
$\hat{\pi}^{e,l}$ $=\frac{E_{N}[(D_{i,e+l}-D_{i,e-1})\cdot\mathbf{1}\{E_{i}=e\}]}{E_{N}[\mathbf{1}\{E_{i}=e\}]}-\frac{E_{N}[(D_{i,e+l}-D_{i,e-1})\cdot\mathbf{1}\{E_{i}\in U\}]}{E_{N}[\mathbf{1}\{E_{i}\in U\}]}$
$\equiv\hat{\pi}_{e,l}^{1}-\hat{\pi}_{e,l}^{2}.$
Here, $E_{N}[\cdot]$ is the sample analog of the conditional expectation. From the estimation procedure, we call this a stacked two-stage least squares (STS) estimator. [^17] Our STS estimator is related to the DID estimators in staggered DID designs proposed by Callaway and Sant’Anna (2021) and Sun and Abraham (2021). The DID estimator of the treatment and the outcome in our STS estimator corresponds to that of Sun and Abraham (2021), and coincides with that of Callaway and Sant’Anna (2021) for the case when no covariates exist and never treated units are used as a control group. Their proposed methods avoid the issue of TWFE estimators in staggered DID designs, while our STS estimator avoids the issue of TWFEIV estimators in staggered DID-IV designs.
Theorem 4 below guarantees the validity of our STS estimator.
Theorem 4.
  • (i)
    Suppose Assumptions 815 hold. Then, the STS estimator $\widehat{CLATT}_{e,e+l}$ is consistent and asymptotically normal.
    $\sqrt{n}(\widehat{CLATT}_{e,e+l}-CLATT_{e,e+l})\xrightarrow{d}\mathcal{N}(0,V(\psi_{i,e,l})).$
  • (ii)
    Suppose that Assumptions 812 and 1618 hold. Then, the STS estimator $\widehat{CLATT}_{e,e+l}$ is consistent and asymptotically normal.
    $\sqrt{n}(\widehat{CLATT}_{e,e+l}-CLATT_{e,e+l})\xrightarrow{d}\mathcal{N}(0,V(\psi_{i,e,l})).$
From Theorem 4, we can also construct the standard error of the STS estimator, using the sample analogue of the asymptotic variance $V(\psi_{i,e,l})$.
Remark 2.
Our estimation procedure is the same if we have a non-binary, ordered treatment or use repeated cross section data. Online Appendix Section B presents the influence function of the STS estimator in repeated cross section settings.
5.2  Weighting scheme
In this subsection, we explain how one can construct the summary causal parameters from each CLATT in staggered DID-IV designs, based on the weighting scheme proposed by Callaway and Sant’Anna (2021).
We consider the following weighting scheme as in Callaway and Sant’Anna (2021):
$\theta^{IV}=\sum_{e}\sum_{t=1}^{T}w(e,t)\cdot CLATT_{e,t},$
where $w(e,t)$ are some reasonable weighting functions assigned to each $CLATT_{e,t}$.
To propose the weighting functions for a variety of summary causal parameters in staggered DID-IV designs, we define the average effect of the instrument on the treatment at a given relative period $l$ from the initial exposure to the instrument in cohort $e$, called the cohort specific average exposed effect on the treated in the first stage ($CAET_{e,e+l}^{1}$).
Definition.
The cohort specific average exposed effect on the treated in the first stage ($CAET_{e,e+l}^{1}$) at a given relative period $l$ from the initial adoption of the instrument is
$CAET_{e,e+l}^{1}=E[D_{i,e+l}-D_{i,e+l}^{\infty}|E_{i}=e].$
If treatment is binary, each $CAET_{e,l}^{1}$ is equal to the share of the compliers in cohort $e$ in period $e+l$:
$CAET^{1}_{e,e+l}$ $=E[D_{i,e+l}^{e}-D_{i,e+l}^{\infty}|E_{i}=e]$
$=Pr(CM_{e,e+l}|E_{i}=e).$
In staggered DID designs, Callaway and Sant’Anna (2021) propose various aggregated measures along with different dimensions of treatment effect heterogeneity. We can employ their framework directly, but in staggered DID-IV designs, we should more carefully specify the weighting functions assigned to each summary measure.
For instance, to aggregate dynamic treatment effects in cohort $e$ over time in staggered DID designs, Callaway and Sant’Anna (2021) consider the following summary measure:
$\theta_{sel}(\tilde{e})=\frac{1}{T-\tilde{e}+1}\sum_{t=\tilde{e}}^{T}CATT(\tilde{e},t),$
where $CATT(\tilde{e},t)$ is the cohort specific average treatment effect on the treated at time $t$ in cohort $\tilde{e}$ (see, e.g., Callaway and Sant’Anna (2021), Sun and Abraham (2021)).
Table 5: Weights for a variety of summary causal parameters
Target Parameter $w(e,t)$
$\theta_{es(l)}^{IV}$ $\mathbf{1}\{e+l\leq T\}\mathbf{1}\{t=e+l\}P(E=e|E+l\leq T)\displaystyle\frac{CAET_{e,e+l}^{1}}{\sum_{e\in\mathcal{S}(E)}CAET_{e,e+l}^{1}}$
$\theta_{es(l,l^{\prime})}^{bal,IV}$ $\mathbf{1}\{e+l^{\prime}\leq T\}\mathbf{1}\{t=e+l\}P(E=e|E+l^{\prime}\leq T)\displaystyle\frac{CAET_{e,e+l}^{1}}{\sum_{e\in\mathcal{S}(E)}CAET_{e,e+l}^{1}}$
$\theta_{sel(\tilde{e})}^{IV}$ $\mathbf{1}\{t\geq e\}\mathbf{1}\{e=\tilde{e}\}\displaystyle\frac{CAET_{\tilde{e},t}^{1}}{\sum_{t=\tilde{e}}^{T}CAET_{\tilde{e},t}^{1}}$
$\theta_{c(\tilde{t})}^{IV}$ $\mathbf{1}\{t\geq e\}\mathbf{1}\{t=\tilde{t}\}P(E=e|E\leq t)\displaystyle\frac{CAET_{e,t}^{1}}{\sum_{e\in\mathcal{S}(E)}CAET_{e,t}^{1}}$
$\theta_{c(\tilde{t})}^{cumm,IV}$ $\mathbf{1}\{t\geq e\}\mathbf{1}\{t\leq\tilde{t}\}P(E=e|E\leq t)\displaystyle\frac{CAET_{e,t}^{1}}{\sum_{e\in\mathcal{S}(E)}CAET_{e,t}^{1}}$
$\theta_{W}^{o,IV}$ $\mathbf{1}\{t\geq e\}P(E=e|E\leq T)/\sum_{e\in\mathcal{S}(E)}\sum_{t=1}^{T}\mathbf{1}\{t\geq e\}P(E=e|E\leq T)$
$\theta_{sel}^{o,IV}$ $\mathbf{1}\{t\geq e\}P(E=e|E\leq T)\displaystyle\frac{CAET_{e,t}^{1}}{\sum_{t=e}^{T}CAET_{e,t}^{1}}$
Notes: This table represents the specific expressions for the weights on each $CLATT(e,t)$ or each $CACRT(e,t)$ in a variety of summary causal parameters. Each target parameter without the superscript IV is defined in Callaway and Sant’Anna (2021). The superscript IV is added to associate each parameter with staggered DID-IV designs.
[…]
$\theta_{sel}(\tilde{e})^{IV}=\sum_{t=\tilde{e}}^{T}\frac{CAET^{1}_{\tilde{e},t}}{\sum_{t=\tilde{e}}^{T}CAET^{1}_{\tilde{e},t}}CLATT_{\tilde{e},t}.$
This parameter summarizes each $CLATT_{\tilde{e},t}$ in cohort $\tilde{e}$ across all post-exposed periods and the weight assigned to each $CLATT_{\tilde{e},t}$ reflects the relative share of the compliers $CM_{\tilde{e},t}$ in period $t$ during the periods after the initial exposure to the instrument in cohort $\tilde{e}$. This weighting scheme would be reasonable in that it is designed to be larger in the period when the proportion of the compliers is relatively higher in cohort $\tilde{e}$.
By similar arguments, we can specify the weighting functions for various summary measures, which correspond with those in Callaway and Sant’Anna (2021). Table 5 summarizes the specific expressions for the weights assigned to each $CLATT_{\tilde{e},t}$ in each summary causal parameter. Note that our proposed weighting functions are the same under non-binary, ordered treatment settings, in which our target parameter is the $CACRT_{\tilde{e},t}$.
Estimation and inference
We can construct the consistent estimator for the summary causal parameter $\theta^{IV}$ as follows.
$\hat{\theta}^{IV}=\sum_{e}\sum_{t=1}^{T}\hat{w}(e,t)\cdot\widehat{CLATT}_{e,t},$
where $\hat{w}(e,t)$ is the sample analog of each w(e,t){w}(e,t) and $\widehat{CLATT}_{e,t}$ is the consistent estimator for each $CLATT_{e,t}$, obtained from the STS regression in Subsection 5.1. Each $\hat{w}(e,t)$ is the regular asymptotically linear estimator:
$\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\zeta_{i,e,t}^{w}+o_{p}(1),$
where $\zeta_{i,e,t}^{w}$ represents the influence function, and satisfies $E[\zeta_{i,e,t}^{w}]=0$ and $E[\zeta_{i,e,t}^{w}{\zeta_{i,e,t}^{w}}^{\top}]<\infty$.
Corollary 1 below represents the asymptotic distribution of the plug-in estimator $\hat{\theta}_{IV}$ and ensures its validity.
Corollary 1.
If the assumptions of Theorem 4 hold,
$\sqrt{n}(\hat{\theta}_{IV}-\theta_{IV})\xrightarrow{d}\mathcal{N}(0,V(l_{i}^{\theta_{IV}})),$
where the influence function $l_{i}^{\theta_{IV}}$ takes the following form:
$l_{i}^{\theta_{IV}}=\sum_{e}\sum_{t=1}^{T}\left(w(e,t)\cdot\psi_{i,e,t}+\zeta_{i,e,t}^{w}\cdot{CLATT}_{e,t}\right).$
5.3  Pre-trends test
In most DID applications, researchers assess the plausibility of the parallel trends assumption by testing for pre-treatment differences in the outcome between treatment and control groups (pre-trends test). In this section, we describe the procedure of the pre-trends test in DID-IV designs to check the validity of the parallel trends assumption in the treatment and the outcome.
Suppose that data are available for period t=1t=-1. Then, one can assess the plausibility of Assumption 5 between period t=1t=-1 and $t=0$ by testing the following null hypothesis
$E[D_{i,0}-D_{i,-1}|E_{i}=0]=E[D_{i,0}-D_{i,-1}|E_{i}=1]$
$\iff$ $E[D_{i,0}((0,0))-D_{i,-1}((0,0))|E_{i}=0]=E[D_{i,0}((0,0))-D_{i,-1}((0,0))|E_{i}=1].$ (1)
Similarly, one can assess the plausibility of Assumption 6 between period t=1t=-1 and $t=0$ by testing the following null hypothesis
$E[Y_{i,0}-Y_{i,-1}|E_{i}=0]=E[Y_{i,0}-Y_{i,-1}|E_{i}=1]$
$\iff$ $E[Y_{i,0}(D_{i,0}((0,0)))-Y_{i,-1}(D_{i,-1}((0,0)))|E_{i}=0]$
$=$ $E[Y_{i,0}(D_{i,0}((0,0)))-Y_{i,-1}(D_{i,-1}((0,0)))|E_{i}=1].$ (2)
These tests can be generalized to multiple pre-exposed periods settings by using the pre-trends tests recently developed in the context of DID designs (Callaway and Sant’Anna (2021), Sun and Abraham (2021), Borusyak et al. (2024)); that is, one can apply these tools to the first stage and the reduced form ($D_{i,t}$ or $Y_{i,t}$ is outcome, $Z_{i,t}$ is treatment) respectively to confirm the plausibility of Assumptions 1314 or Assumptions 1617 in pre-exposed periods.
6  Application
We illustrate the empirical relevance of our findings in the setting of Oreopoulos (2006), estimating returns to schooling in the United Kingdom. We first assess the identifying assumptions in staggered DID-IV designs implicitly imposed by Oreopoulos (2006). We then estimate the TWFEIV regression in the author’s setting. Finally, we estimate the target parameter and its summary measure by employing our proposed method and weighting scheme.
6.1  Setting
Oreopoulos (2006) estimates returns to schooling using a major education reform in the UK that increased the years of compulsory schooling from 14 to 15. Specifically, Oreopoulos (2006) exploits variation resulting from the different timing of implementation of school reforms between Britain (England, Scotland, and Wales) and Northern Ireland as an instrument for education attainment: the school-leaving age increased in Britain in 1947, but was not implemented until 1957 in Northern Ireland. [^18] In the first part of his analysis, Oreopoulos (2006) adopts regression discontinuity designs (RDD) and analyzes the data sets in Britain and Northern Ireland separately. Due to the imprecision of their standard errors, Oreopoulos (2006) then moves to “a difference-in-differences and instrumental-variables analysis by combining the two sets of U.K. data”. The data are a sample of individuals in Britain and Northern Ireland, who were aged $14$ between $1936$ and $1965$, constructed from combining the series of U.K. General Household Surveys between $1984$ and $2006$; see Oreopoulos (2006), Oreopoulos (2008) for details.
Oreopoulos (2006) (more precisely, Oreopoulos (2008)) runs the following two-way fixed effects instrumental variable regression with the education reform as an excluded instrument for education attainment.
$Y_{i,t}=\mu_{n.}+\delta_{.t}+\beta_{IV}D_{i,t}+\epsilon_{i,t},$ (3)
$D_{i,t}=\gamma_{n.}+\zeta_{.t}+\pi Z_{i,t}+\eta_{i,t}.$ (4)
Here, a cohort (a year when aged $14$) plays a role of time as it determines the exposure to the policy change. The dependent variables $Y_{i,t}$ and $D_{i,t}$ are log annual earnings and education attainment for unit $i$ and cohort $t$, respectively. Both the first stage and the reduced form regressions include a birth cohort fixed effect and a North Ireland fixed effect. The binary instrument $Z_{i,t}\in\{0,1\}$ takes one if unit $i$ in cohort $t$ is exposed to the policy change.
The staggered introduction of the school reform can be viewed as a natural experiment, but is not randomized across regions in reality; Oreopoulos (2006) notes that the reform was implemented with political support, taking into account costs and benefit. This indicates that Oreopoulos (2006) implicitly relies on a staggered DID-IV identification strategy instead of exploiting the random variation of the policy change. Indeed, Oreopoulos (2006) presents the corresponding plots of British and Northern Irish average education attainment and average log earnings by cohort to illustrate the evolution of these variables before and after the policy shock.
6.2  Assessing the identifying assumptions in staggered DID-IV designs
We first discuss the validity of the staggered DID-IV identification strategy in Oreopoulos (2006). In the author’s setting, our target parameter is the cohort specific average causal response on the treated (CACRT), as education attainment is a non-binary, ordered treatment.
Exclusion restriction.
It would be plausible, given that the policy reform did not affect log annual earnings other than by increasing education attainment. This assumption may be violated for instance if the reform affected both the quality and quantity of education.
Monotonicity assumption.
It would be automatically satisfied in the author’s setting: the policy change (instrument) increased the minimum schooling-leaving age from $14$ to $15$, ensuring that there are no defiers during the periods after the policy shock.
No anticipation in the first stage.
It would be plausible that there is no anticipation, if the treatment adoption behavior is the same as the one in the absence of the policy change before its implementation in England. This assumption may be violated if units have private knowledge about the probability of extended education and manipulate their education attainment before the policy shock.
Next, we assess the validity of the parallel trends assumptions in the treatment and the outcome, using the interacted two-way fixed effects regressions proposed by Sun and Abraham (2021) in the first stage and the reduced form, respectively. The results are shown in Figure 1.
Figure 1: The effect of the instrument in the first stage and reduced form in the setting of Oreopoulos (2006). Notes: This figure presents the results of the effect of the school reform on education attainment (Panel (a)) and on log annual earnings (Panel (b)) under the staggered DID-IV identification strategy. The unexposed group $U$ is a last-exposed cohort, North Ireland and the reference period is $t=1946$. The blue lines represent the estimates with pointwise $95\%$ confidence intervals for pre-exposed periods in both panels. These should be equal to zero under the null hypothesis that the parallel trends assumptions in the treatment and the outcome hold. The red lines represent the estimates with pointwise $95\%$ confidence intervals for post-exposed periods in both panels.
Parallel trends assumption in the treatment.
It requires that the expectation of education attainment would have followed the same path between England and North Ireland across cohorts in the absence of the school reform. Panel (a) in Figure 1 plots the results of the interacted two-way fixed effects regression in the first stage along with a $95\%$ pointwise confidence interval. The pre-exposed estimates are not significantly different from zero and indicate the validity of Assumption 13.
Parallel trends assumption in the outcome.
It requires that the expectation of log annual earnings would have followed the same evolution between England and North Ireland across cohorts in the absence of the school reform. Panel (b) in Figure 1 plots the results of the interacted two-way fixed effects regression in the reduced form along with a $95\%$ pointwise confidence interval. The pre-exposed estimates seems consistent with Assumption 14: though an upward pre-trends exists, all the estimates before the initial exposure to the policy change are not significantly different from zero.
Figure 1 also sheds light on the dynamic effects of the school reform on education attainment and log annual earnings in post-exposed periods. In Panel (a), the estimated increase in years of schooling after the reform ranges from $0.56$ to $0.87$, and all estimates are statistically significant. In Panel (b), the estimated increase in log annual earnings after the reform varies from $8\%$ to $25\%$, and $5$ out of $10$ estimates are statistically significant. [^19] Note that the post-exposed estimates in Panel (b) do not capture each CACRT in post-exposed periods: each estimate in the reduced form is not scaled by the corresponding estimate in the first stage.
6.3  Illustrating our estimation method
We start by estimating two-way fixed effects instrumental variable regression in the author’s setting. To clearly illustrate the pitfalls of TWFEIV regression, in our estimation, we slightly modify the author’s specification. Oreopoulos (2006) (more precisely Oreopoulos (2008)) includes some covariates (survey year, sex, and a quartic in age) and runs the weighted regression in the main specification (see Oreopoulos (2006), Oreopoulos (2008) for details), whereas we exclude such covariates and do not apply their weights to our regression.
The result is shown in Table 6. The TWFEIV estimate is 0.009-0.009 and not significantly different from 0. This indicates that, on the whole, the returns to schooling in the UK is nearly zero. [^20] Young (2024) revisits Oreopoulos (2006) and shows that the 2SLS estimates in this setting can be numerically unstable due to near collinearity among cohort indicators. Following Young (2024), we conduct stability checks based on variable and observation ordering, and find that our TWFEIV estimate does not suffer from such numerical instability. However, this may be a misleading conclusion. We cannot interpret that the TWFEIV estimand captures the properly weighted average of all possible CACRTs if the effect of the school reform on education attainment or log annual earnings is not stable across cohorts (Miyaji (2024)).
In Online Appendix Section C, we quantify the bias terms of the TWFEIV estimand by using Lemma 7 in Miyaji (2024), and show that the TWFEIV estimand is negatively biased in Oreopoulos (2006). Specifically, we show that all the bias terms are positive, while all the assigned weights are negative, which yields the downward bias for the TWFEIV estimand. [^21] When we run the TWFEIV regression in the author’s setting, we treat the set of already exposed units in North Ireland as controls between $1957$ and $1965$, as both regions are already exposed to the policy shock during these periods. This procedure performed by the TWFEIV regression is called the “bad comparisons” in the recent DID literature (c.f. Goodman-Bacon (2021)), and yields the bias in the TWFEIV estimand from the properly weighted average of all possible CACRTs.
Figure 2: The cohort specific average causal response on the treated in each relative period in the setting of Oreopoulos (2006). Notes: This figure shows the results for returns to schooling under the staggered DID-IV identification strategy. The unexposed group $U$ is North Ireland and the reference period is period $t=1946$. The red lines represent the stacked two stage least squares estimates with pointwise $95\%$ confidence intervals for post-exposed periods.
We now employ the proposed method to estimate each CACRT. First, we create data sets by cohort $t$ ($t=1947,\dots,1956$). Each data set only contains units of cohort $t$ and cohort $t=1946$ in England and North Ireland. We define North Ireland as an unexposed group $U$. We then run the STS regression in each data set. The standard error is calculated by using the influence function shown in equation (B) in Online Appendix Section B.
Table 6: Returns to schooling in Oreopoulos (2006)
Estimate Standard Error 95% CI
TSLS with fixed effects -0.009 0.039 [-0.085, 0.068]
$\theta_{sel}^{IV}(e)$ 0.240 0.098 [0.047, 0.433]
Notes: Sample size $82790$ observations.
Figure 2 plots the point estimates and the corresponding $95\%$ confidence intervals in each relative period after the school reform. The estimates for each $CACRT_{e,t}$ range from $14\%$ to $38\%$ with wide confidence intervals, and $4$ out of $10$ estimates are statistically significant.
Finally, we estimate the summary causal measure by aggregating each $CACRT_{e,t}$. Specifically, we estimate the weighted average of each $CACRT_{e,t}$ during post-exposed periods in England ($e=1947$).
$\theta_{sel}^{IV}(e)=\sum_{t=1947}^{1956}\frac{CAET^{1}_{e,t}}{\sum_{t=1947}^{1956}CAET^{1}_{e,t}}CACRT_{e,t}.$
Here, each weight assigned to $CACRT_{e,t}$ represents the relative amount of the effect of the policy reform on education attainment during post-exposed periods in England.
The result is shown in Table 6. The estimate is $0.24$ and it is significantly different from zero. The estimated returns to schooling are substantial, perhaps because each $CACRT_{e,t}$ captures the returns to schooling among the compliers: such units may belong to relatively low-skilled labor or low-income family that potentially have much to gain from school reform.
The estimates obtained from our proposed method and weighting scheme are significantly different from the TWFEIV estimate: our STS estimates and its summary measure are all positive, whereas the TWFEIV estimate is strictly negative. Overall, our results indicate that the economic returns of education are substantial in the UK, and the estimation method matters in staggered DID-IV designs in practice.
7  Conclusion
In this paper, we formalize an instrumented difference-in-differences (DID-IV). First, we consider a simple setting with two periods and two groups. In this setting, our DID-IV design mainly comprises a monotonicity assumption, and parallel trends assumptions in the treatment and the outcome between the two groups. We show that in $2\times 2$ DID-IV designs, the Wald-DID estimand captures the local average treatment effect on the treated (LATET). After establishing $2\times 2$ DID-IV designs, we clarify the differences between DID-IV and Fuzzy DID designs considered in de Chaisemartin and D’Haultfœuille (2018), and provide the implications of these differences for treatment adoption behavior, the interpretation of the target parameter, and the use of the Wald-DID estimand.
Next, we consider the DID-IV design in more than two periods with units being exposed to the instrument at different times. We call this a staggered DID-IV design, and formalize the target parameter and identifying assumptions. Specifically, in this design, our target parameter is the cohort specific local average treatment effects on the treated (CLATT). The identifying assumptions are the natural generalization of those in $2\times 2$ DID-IV designs.
We also provide the estimation method in staggered DID-IV designs that does not require strong restrictions on treatment effect heterogeneity. Our estimation method carefully chooses the comparison groups and does not suffer from the bias arising from the time-varying exposed effects. We also propose the weighting scheme in staggered DID-IV designs, and explain how one can conduct pre-trends tests to assess the validity of the parallel trends assumptions in DID-IV designs.
Finally, we illustrate the empirical relevance of our findings with the setting of Oreopoulos (2006), who estimates returns to schooling in the UK, exploiting the timing variation of the introduction of school reforms between British and North Ireland. In this application, the TWFEV regression, the conventional approach to implement a staggered DID-IV design, yields a negative estimate. By contrast, our STS regression and weighting scheme indicate the substantial gain from schooling. This empirical application illustrates that the estimation method matters in staggered DID-IV designs in practice.
Overall, this paper provides a new econometric framework for estimating the causal effects when the treatment adoption is potentially endogenous over time, but researchers can exploit variation in policy adoption timing as an instrument for treatment. To avoid the issue of using the TWFEIV estimator, we also provide a reliable estimation method that is free from strong restrictions on treatment effect heterogeneity. Further developing alternative estimation methods and diagnostic tools will be a promising area for future research, facilitating the credibility of DID-IV design in practice.
Appendix
Proof of Theorem 1
Proof.
First we consider the numerator of the Wald-DID estimand. For the numerator, we have
$E[Y_{i,1}-Y_{i,0}|E_{i}=1]-E[Y_{i,1}-Y_{i,0}|E_{i}=0]$
$=$ $E[Y_{i,1}(D_{i,1}((0,1)))-Y_{i,0}(D_{i,0}((0,1)))|E_{i}=1]$
- $E[Y_{i,1}(D_{i,1}((0,0)))-Y_{i,0}(D_{i,0}((0,0)))|E_{i}=0]$
$=$ $E[Y_{i,1}(D_{i,1}((0,1)))-Y_{i,1}(D_{i,1}((0,0)))|E_{i}=1]$
$+$ $E[Y_{i,1}(D_{i,1}((0,0)))-Y_{i,0}(D_{i,0}((0,1)))|E_{i}=1]$
- $E[Y_{i,1}(D_{i,1}((0,0)))-Y_{i,0}(D_{i,0}((0,0)))|E_{i}=0]$
$=$ $E[Y_{i,1}(D_{i,1}((0,1)))-Y_{i,1}(D_{i,1}((0,0)))|E_{i}=1]$
$+$ $E[Y_{i,1}(D_{i,1}((0,0)))-Y_{i,0}(D_{i,0}((0,0)))|E_{i}=1]$
- $E[Y_{i,1}(D_{i,1}((0,0)))-Y_{i,0}(D_{i,0}((0,0)))|E_{i}=0]$
$=$ $E[(D_{i,1}((0,1))-D_{i,1}((0,0)))\cdot(Y_{i,1}(1)-Y_{i,1}(0))|E_{i}=1]$
$=$ $LATET\cdot Pr(CM^{Z}|E_{i}=1).$ (5)
The first equality follows from Assumption 1 and Assumption 2. The second equality follows from the simple manipulation. The third equality follows from Assumption 4. The fourth equality follows from Assumption 6 and simple calculation. The fifth equality follows from the Law of Iterated Expectations and Assumption 3.
For the denominator, we have
$E[D_{i,1}-D_{i,0}\mid E_{i}=1]-E[D_{i,1}-D_{i,0}\mid E_{i}=0]$
$=$ $E[D_{i,1}((0,1))-D_{i,0}((0,1))\mid E_{i}=1]-E[D_{i,1}((0,0))-D_{i,0}((0,0))\mid E_{i}=0]$
$=$ $E[D_{i,1}((0,1))-D_{i,1}((0,0))\mid E_{i}=1]$
$+\Big\{E[D_{i,1}((0,0))-D_{i,0}((0,1))\mid E_{i}=1]-E[D_{i,1}((0,0))-D_{i,0}((0,0))\mid E_{i}=0]\Big\}$
$=$ $E[D_{i,1}((0,1))-D_{i,1}((0,0))\mid E_{i}=1]$
$+\Big\{E[D_{i,1}((0,0))-D_{i,0}((0,0))\mid E_{i}=1]-E[D_{i,1}((0,0))-D_{i,0}((0,0))\mid E_{i}=0]\Big\}$
$=$ $\Pr(CM^{Z}\mid E_{i}=1).$ (6)
Here, the second equality follows from the simple manipulation. The third equality follows from Assumption 4 and the forth equality follows from Assumption 3 and Assumption 5.
Combining the result (5) with (6), we obtain the desirable result. ∎
Proof of Theorem 2
Proof.
Fix $e\in\mathcal{E}$ and $l\in\{0,\dots,T-e\}$. Note that Assumption 15 guarantees that the Wald-DID estimand $w_{e,l}^{DID}$ is well defined.
First, we consider the numerator of $w_{e,l}^{DID}$. From Assumptions 815, we have:
$E[Y_{i,e+l}-Y_{i,e-1}\mid E_{i}=e]-E[Y_{i,e+l}-Y_{i,e-1}\mid E_{i}=\infty]$
$=$ $E[Y_{i,e+l}(D^{e}_{i,e+l})-Y_{i,e-1}(D^{e}_{i,e-1})\mid E_{i}=e]$
- $E[Y_{i,e+l}(D^{\infty}_{i,e+l})-Y_{i,e-1}(D^{\infty}_{i,e-1})\mid E_{i}=\infty]$
$=$ $E[Y_{i,e+l}(D^{e}_{i,e+l})-Y_{i,e-1}(D^{\infty}_{i,e-1})\mid E_{i}=e]$
- $E[Y_{i,e+l}(D^{\infty}_{i,e+l})-Y_{i,e-1}(D^{\infty}_{i,e-1})\mid E_{i}=\infty]$
$=$ $E[Y_{i,e+l}(D^{e}_{i,e+l})-Y_{i,e+l}(D^{\infty}_{i,e+l})\mid E_{i}=e]$
$+$ $E[Y_{i,e+l}(D^{\infty}_{i,e+l})-Y_{i,e-1}(D^{\infty}_{i,e-1})\mid E_{i}=e]$
- $E[Y_{i,e+l}(D^{\infty}_{i,e+l})-Y_{i,e-1}(D^{\infty}_{i,e-1})\mid E_{i}=\infty]$
$=$ $E[Y_{i,e+l}(D^{e}_{i,e+l})-Y_{i,e+l}(D^{\infty}_{i,e+l})\mid E_{i}=e]$
$=$ $E[(D^{e}_{i,e+l}-D^{\infty}_{i,e+l})\cdot(Y_{i,e+l}(1)-Y_{i,e+l}(0))\mid E_{i}=e]$
$=$ $CLATT_{e,e+l}\Pr(CM_{e,e+l}\mid E_{i}=e).$ (7)
The first equality follows from Assumptions 810. The second equality follows from Assumption 12. The third equality follows from Assumption 14. The final equality follows from Assumption 11.
Next, we consider the denominator of $w_{e,l}^{DID}$. From Assumptions 815, we have:
$E[D_{i,e+l}-D_{i,e-1}\mid E_{i}=e]-E[D_{i,e+l}-D_{i,e-1}\mid E_{i}=\infty]$
$=$ $E[D^{e}_{i,e+l}-D^{e}_{i,e-1}\mid E_{i}=e]-E[D^{\infty}_{i,e+l}-D^{\infty}_{i,e-1}\mid E_{i}=\infty]$
$=$ $E[D^{e}_{i,e+l}-D^{\infty}_{i,e-1}\mid E_{i}=e]-E[D^{\infty}_{i,e+l}-D^{\infty}_{i,e-1}\mid E_{i}=\infty]$
$=$ $E[D^{e}_{i,e+l}-D^{\infty}_{i,e+l}\mid E_{i}=e]+E[D^{\infty}_{i,e+l}-D^{\infty}_{i,e-1}\mid E_{i}=e]$
- $E[D^{\infty}_{i,e+l}-D^{\infty}_{i,e-1}\mid E_{i}=\infty]$
$=$ $E[D^{e}_{i,e+l}-D^{\infty}_{i,e+l}\mid E_{i}=e]$
$=$ $\Pr(CM_{e,e+l}\mid E_{i}=e).$ (8)
The first equality follows from Assumption 8. The second equality follows from Assumption 12. The third equality follows from Assumption 13. The final equality follows from Assumption 11.
Combining (7) with (8), we obtain the desirable result. ∎
Proof of Theorem 3
Proof.
Fix $e\in\mathcal{E}$ and $l\in\{0,\dots,T-e\}$ such that $l<\bar{e}-e$. Note that Assumption 18 guarantees that the Wald-DID estimand $w_{e,l}^{DID,ny}$ is well defined.
First, we consider the numerator of $w_{e,l}^{DID,ny}$. From Assumptions 812 and Assumption 17, we have:
$E[Y_{i,e+l}-Y_{i,e-1}\mid E_{i}=e]-E[Y_{i,e+l}-Y_{i,e-1}\mid Z_{i,e+l}=0,E_{i,e}=0]$
$=$ $E[Y_{i,e+l}(D^{e}_{i,e+l})-Y_{i,e-1}(D^{\infty}_{i,e-1})\mid E_{i}=e]$
- $E[Y_{i,e+l}(D^{\infty}_{i,e+l})-Y_{i,e-1}(D^{\infty}_{i,e-1})\mid Z_{i,e+l}=0,E_{i,e}=0]$
$=$ $E[Y_{i,e+l}(D^{e}_{i,e+l})-Y_{i,e+l}(D^{\infty}_{i,e+l})\mid E_{i}=e]$
$+$ $E[Y_{i,e+l}(D^{\infty}_{i,e+l})-Y_{i,e-1}(D^{\infty}_{i,e-1})\mid E_{i}=e]$
- $E[Y_{i,e+l}(D^{\infty}_{i,e+l})-Y_{i,e-1}(D^{\infty}_{i,e-1})\mid Z_{i,e+l}=0,E_{i,e}=0]$
$=$ $E[Y_{i,e+l}(D^{e}_{i,e+l})-Y_{i,e+l}(D^{\infty}_{i,e+l})\mid E_{i}=e]$
$=$ $E[(D^{e}_{i,e+l}-D^{\infty}_{i,e+l})\cdot(Y_{i,e+l}(1)-Y_{i,e+l}(0))\mid E_{i}=e]$
$=$ $CLATT_{e,e+l}\Pr(CM_{e,e+l}\mid E_{i}=e).$ (9)
The first equality follows from Assumptions 810 and Assumption 12. The third equality follows from Assumption 17. The final equality follows from Assumption 11.
Next, we consider the denominator of $w_{e,l}^{DID}$. From Assumptions 812 and Assumption 16, we have:
$E[D_{i,e+l}-D_{i,e-1}\mid E_{i}=e]-E[D_{i,e+l}-D_{i,e-1}\mid Z_{i,e+l}=0,E_{i,e}=0]$
$=$ $E[D^{e}_{i,e+l}-D^{\infty}_{i,e-1}\mid E_{i}=e]-E[D^{\infty}_{i,e+l}-D^{\infty}_{i,e-1}\mid Z_{i,e+l}=0,E_{i,e}=0]$
$=$ $E[D^{e}_{i,e+l}-D^{\infty}_{i,e+l}\mid E_{i}=e]+E[D^{\infty}_{i,e+l}-D^{\infty}_{i,e-1}\mid E_{i}=e]$
- $E[D^{\infty}_{i,e+l}-D^{\infty}_{i,e-1}\mid Z_{i,e+l}=0,E_{i,e}=0]$
$=$ $E[D^{e}_{i,e+l}-D^{\infty}_{i,e+l}\mid E_{i}=e]$
$=$ $\Pr(CM_{e,e+l}\mid E_{i}=e).$ (10)
The first equality follows from Assumption 8 and Assumption 12. The third equality follows from Assumption 13. The final equality follows from Assumption 11.
Combining (9) with (10), we obtain the desirable result. ∎
Proof of Theorem 4
Proof.
First, we show that $\widehat{CLATT}_{e,e+l}$ is consistent for the corresponding $CLATT_{e,e+l}$. Consider case (i), where a never-exposed cohort is used as the control group. By the Law of Large Numbers (LLN) and the Continuous Mapping Theorem (CMT), $\widehat{CLATT}_{e,e+l}\xrightarrow[p]{}w^{DID}_{e,l}$ as defined in Section 3.3. Then, by Theorem 2, we have $w^{DID}_{e,l}=CLATT_{e,e+l}$. The same argument, together with Theorem 3, establishes consistency under case (ii), where not-yet-exposed cohorts are used as the control group.
Next, we prove that our STS estimator is asymptotically normal, deriving its influence function. This derivation does not depend on whether we are in case (i) or case (ii). We define $\theta_{e,l}$ to be
$\theta_{e,l}$ $\equiv\frac{E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=e]-E[Y_{i,e+l}-Y_{i,e-1}|E_{i}\in U]}{E[D_{i,e+l}-D_{i,e-1}|E_{i}=e]-E[D_{i,e+l}-D_{i,e-1}|E_{i}\in U]}(=CLATT_{e,e+l})$
$\equiv\frac{\alpha_{e,l}^{1}-\alpha_{e,l}^{2}}{\pi_{e,l}^{1}-\pi_{e,l}^{2}}.$
The following fact, also found in de Chaisemartin and D’Haultfœuille (2018), is repeatedly used in the derivation of the influence function.
Fact 1.
If
$\sqrt{n}(\hat{A}-A)=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}a_{i}+o_{p}(1),\hskip 8.53581pt\sqrt{n}(\hat{B}-B)=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}b_{i}+o_{p}(1),$
we have
$\sqrt{n}\left(\frac{\hat{A}}{\hat{B}}-\frac{A}{B}\right)=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\frac{a_{i}-(A/B)b_{i}}{B}+o_{p}(1).$
Then, by some calculations, we obtain
$\sqrt{n}(\widehat{CLATT}_{e,e+l}-CLATT_{e,e+l})$ $=\frac{1}{\sqrt{n}}\sum_{i}^{n}\psi_{i,e,l}+o_{p}(1).$
The influence function $\psi_{i,e,l}$ is:
$\psi_{i,e,l}$
$=\frac{1}{\pi_{e,l}^{1}-\pi_{e,l}^{2}}\Bigg(\frac{\mathbf{1}\{E_{i}=e\}\left[\delta_{i,e,l}^{p}-E[\delta_{i,e,l}^{p}|E_{i}=e]\right]}{E[\mathbf{1}\{E_{i}=e\}]}-\frac{\mathbf{1}\{E_{i}\in U\}\left[\delta_{i,e,l}^{p}-(E[\delta_{i,e,l}^{p}|E_{i}\in U])\right]}{E[\mathbf{1}\{E_{i}\in U\}]}\Bigg),$ (11)
where we define $\delta_{i,e,l}^{p}=(Y_{i,e+l}-Y_{i,e-1})-\theta_{e,l}\cdot(D_{i,e+l}-D_{i,e-1})$.
Therefore, the asymptotic variance $V(\psi_{i,e,l})$ is
$V(\psi_{i,e,l})$ $=\frac{1}{(\pi_{e,l}^{1}-\pi_{e,l}^{2})^{2}}\Bigg(\frac{E\big[\mathbf{1}\{E_{i}=e\}\left[\delta_{i,e,l}^{p}-E[\delta_{i,e,l}^{p}|E_{i}=e]\right]^{2}\big]}{E[\mathbf{1}\{E_{i}=e\}]^{2}}$
$+\frac{E\big[\mathbf{1}\{E_{i}\in U\}\left[\delta_{i,e,l}^{p}-(E[\delta_{i,e,l}^{p}|E_{i}\in U]\right]^{2}\big]}{E[\mathbf{1}\{E_{i}\in U\}]^{2}}\Bigg).$
Appendix A Extensions
This section considers the non-binary, ordered treatment and repeated cross section settings in DID-IV designs. It also includes the case where we introduce the treatment path in $2\times 2$ DID-IV designs.
A.1  Non-binary, ordered treatment
In this subsection, we present the identification results in DID-IV designs with non-binary, ordered treatment. Theorem 5 below shows that if we have a non-binary, ordered treatment, the Wald-DID estimand is equal to the ACRT under Assumptions 1-7, which replace the binary treatment with a non-binary one.
Theorem 5.
If Assumptions 1-7 hold, the Wald-DID estimand $W_{DID}$ is equal to the ACRT.
Proof.
Let $\lambda_{Z,j}^{t}=\mathbf{1}(D_{t}((0,z))\geq j)$ be the indicator function for $t\in\{0,1\},z\in\{0,1\}$ and $j\in\{0,\dots,J+1\}$. Note that $\lambda_{Z,0}^{t}=1$ and $\lambda_{Z,J+1}^{t}=0$ hold for all $t$ and $z$ by construction. Here, we can rewrite the observed outcome $Y_{t}$ as follows:
$Y_{t}$ $=Z_{1}\cdot Y_{t}(D_{t}((0,1)))+(1-Z_{1})\cdot Y_{t}(D_{t}((0,0)))$
$=\left\{Z_{1}\cdot\sum_{j=0}^{J}Y_{t}(j)\cdot(\lambda_{1,j}^{t}-\lambda_{1,j+1}^{t})\right\}+\left\{(1-Z_{1})\cdot\sum_{j=0}^{J}Y_{t}(j)\cdot(\lambda_{0,j}^{t}-\lambda_{0,j+1}^{t})\right\}.$
where we use Assumption 1 and 2.
For the numerator of the Wald DID estimand, we have
$E[Y_{1}-Y_{0}|E=1]-E[Y_{1}-Y_{0}|E=0]$
$=$ $E[Y_{1}(D_{1}((0,1)))-Y_{1}(D_{1}((0,0)))|E=1].$
This follows from Assumption 4 and Assumption 6.
In terms of the $\lambda_{Z,j}^{t}$, we write the previous expression as:
$E[Y_{1}(D_{1}((0,1)))-Y_{1}(D_{1}((0,0)))|E=1]$
$=E\left[\sum_{j=0}^{J}Y_{1}(j)\cdot[\lambda_{1,j}^{1}-\lambda_{1,j+1}^{1}-\lambda_{0,j}^{1}+\lambda_{0,j+1}^{1}]|E=1\right]$
$=E\left[\sum_{j=1}^{J}(Y_{1}(j)-Y_{1}(j-1))\cdot[\lambda_{1,j}^{1}-\lambda_{0,j}^{1}]+Y_{1}(0)\cdot(\lambda_{1,0}^{1}-\lambda_{0,0}^{1})|E=1\right]$
$=E\left[\sum_{j=1}^{J}(Y_{1}(j)-Y_{1}(j-1))\cdot[\lambda_{1,j}^{1}-\lambda_{0,j}^{1}]|E=1\right].$
where the second equality follows from $\lambda_{1,0}^{1}-\lambda_{0,0}^{1}=1$.
Note that $\lambda_{1,j}^{1}\geq\lambda_{0,j}^{1}$ from Assumption 3 and that $\lambda_{1,j}^{1}$ and $\lambda_{0,j}^{1}$ can take only two values, zero or one. Therefore, we obtain:
$E\left[\sum_{j=1}^{J}(Y_{1}(j)-Y_{1}(j-1))\cdot[\lambda_{1,j}^{1}-\lambda_{0,j}^{1}]|E=1\right]$
$=\sum_{j=1}^{J}E[(Y_{1}(j)-Y_{1}(j-1)|\lambda_{1,j}^{1}-\lambda_{0,j}^{1}=1,E=1]Pr(\lambda_{1,j}^{1}-\lambda_{0,j}^{1}=1|E=1)$
$=\sum_{j=1}^{J}E[(Y_{1}(j)-Y_{1}(j-1)|D_{1}((0,1))\geq j>D_{1}((0,0)),E=1]$
$\times Pr(D_{1}((0,1))\geq j>D_{1}((0,0))|E=1).$ (12)
For the denominator of the Wald-DID estimand, by similar argument, we obtain:
$E[D_{1}-D_{0}|E=1]-E[D_{1}-D_{0}|E=0]$
$=$ $E[D_{1}((0,1))-D_{1}((0,0))|E=1].$
This follows from Assumptions 45.
Then, we have:
$E[D_{1}((0,1))-D_{1}((0,0))|E=1]$
$=E\left[\sum_{j=0}^{J}j\cdot[\lambda_{1,j}^{1}-\lambda_{1,j+1}^{1}-\lambda_{0,j}^{1}+\lambda_{0,j+1}^{1}]|E=1\right]$
$=E\left[\sum_{j=1}^{J}[\lambda_{1,j}^{1}-\lambda_{0,j}^{1}]|E=1\right]$
$=\sum_{j=1}^{J}Pr(D_{1}((0,1))\geq j>D_{1}((0,0))|E=1).$ (13)
Combining the result (13) with (12), we obtain the desired result. ∎
A.2  Repeated cross sections
Here, we present our identification results under repeated cross section settings. First, we consider two groups and two periods settings. Next, we consider multiple period settings with staggered instrument.
A.2.1 Two time periods
Let $Y_{i}$ and $D_{i}$ denote the outcome and the treatment for unit $i$. Let $Z_{i}$ denote the instrument path for unit $i$: $Z_{i}=(0,0)$ if unit $i$ is not exposed to the instrument and $Z_{i}=(0,0)$ if unit $i$ is exposed to the instrument. Let $E_{i}\in\{0,1\}$ denote the group indicator for unit $i$: $E_{i}=1$ if $Z_{i}=(0,1)$ and $E_{i}=0$ if $Z_{i}=(0,0)$. Let $T_{i}\in\{0,1\}$ denote the binary indicator for time. For all $z\in\mathcal{S}(Z)$, let $Y_{i}(0,z),Y_{i}(1,z)$, and $D_{i}(z)$ denote potential outcomes and potential treatment choices for unit $i$. Let $D_{i,t}(z)$ denote potential treatment choices under $T_{i}=t$. We assume that $\{Y_{i},D_{i},Z_{i},E_{i},T_{i}\}_{i=1}^{n}$ are i.i.d.
In two periods and two groups settings, our target parameter is the local average treatment effect on the treated in period $1$:
$LATET$ $\equiv E[Y_{i}(1)-Y_{i}(0)|T_{i}=1,E_{i}=1,D_{i,1}((0,1))>D_{i,1}((0,0))]$
$=E[Y_{i}(1)-Y_{i}(0)|T_{i}=1,E_{i}=1,CM^{Z}].$
We make the following identification assumptions for the Wald-DID estimand to capture the LATET. These assumptions are suitable for repeated cross section settings.
Assumption 19 (Exclusion restriction).
$\forall z\in\mathcal{S}(Z),Y_{i}(0,z)=Y_{i}(0),Y_{i}(1,z)=Y_{i}(1).$
Assumption 20 (Monotonicity).
$Pr(D_{i,1}((0,1))\geq D_{i,1}((0,0)))=1\hskip 5.69054pt\text{or}\hskip 5.69054ptPr(D_{i,1}((0,1))\leq D_{i,1}((0,0)))=1.$
Here, we consider the monotonicity assumption that rules out the defiers, as our target parameter measures the treatment effects among the compliers.
Assumption 21 (No anticipation in the first stage).
$D_{i,0}((0,1))=D_{i,0}((0,0))\hskip 5.69054pta.s.\hskip 8.53581pt\text{for all units $i$ with $E_{i}=1$}.$
Assumption 22 (Parallel trends assumption in the treatment).
$E[D_{i}((0,0))|E_{i}=1,T_{i}=1]-E[D_{i}((0,0))|E_{i}=1,T_{i}=0]$
$=$ $E[D_{i}((0,0))|E_{i}=0,T_{i}=1]-E[D_{i}((0,0))|E_{i}=0,T_{i}=0].$
Assumption 23 (Parallel trends assumption in the outcome).
$E[Y_{i}(D_{i}(0,0))|E_{i}=1,T_{i}=1]-E[Y_{i}(D_{i}(0,0))|E_{i}=1,T_{i}=0]$
$=$ $E[Y_{i}(D_{i}(0,0))|E_{i}=0,T_{i}=1]-E[Y_{i}(D_{i}(0,0))|E_{i}=0,T_{i}=0].$
Assumption 24 (Relevance condition).
$E[D_{i}|E_{i}=1,T_{i}=1]-E[D_{i}|E_{i}=1,T_{i}=0]$
- $(E[D_{i}|E_{i}=0,T_{i}=1]-E[D_{i}|E_{i}=0,T_{i}=0])>0.$
By similar arguments in the proof of Theorem 1 (Theorem 5), one can show that the Wald-DID estimand is equal to the LATET (ACRT) in period 11 under Assumptions 1924 (thus, we omit the proof for brevity).
A.2.2 Multiple time periods
Let $T_{i}\in\{1,\dots,T\}$ denote the time period when unit $i$ is observed and let $Z_{i,t}$ denote the instrument status under $T_{i}=t$. We adopt Assumption 8 in Section 3; that is, we assume the staggered assignment of the instrument across units. In addition, we adopt Assumption 19 (Exclusion restriction), which allows us to write $Y_{i}(D_{i},z)=Y_{i}(D_{i})$. Let $E_{i}=\min\{t:Z_{i,t}=1\}\in\{2,\dots,T,\infty\}$ denote the cohort which unit $i$ belongs to. Let $E_{i,e}=\mathbf{1}\{E_{i}=e\}$ denote the binary indicator that takes one if unit $i$ belongs to cohort $e$. Let $\bar{e}=\max_{i=1,\dots,n}E_{i}$ denote the largest cohort value in the dataset. Let $\mathcal{E}=\mathcal{S}(E_{i})\setminus\{\bar{e}\}\subseteq\{2,3,\dots,T\}$ denote the support of $E_{i}$ excluding $\bar{e}$.
Let $Y_{i}(D_{i}^{e})$ and $D_{i}^{e}$ denote outcomes and potential treatment choices when unit $i$ belongs to cohort $e$. Let $D_{i,t}^{e}$ denote potential treatment choices under $T_{i}=t$ when unit $i$ belongs to cohort $e$. We assume that $\{Y_{i},D_{i},Z_{i},E_{i},T_{i}\}_{i=1}^{n}$ are i.i.d.
In multiple periods and cohorts settings with a binary treatment, our target parameter is the $CLATT_{e,e+l}$ in period $T=e+l$:
$CLATT_{e,e+l}$ $\equiv E[Y_{i}(1)-Y_{i}(0)|T_{i}=e+l,E_{i}=e,D_{i,e+l}^{e}>D_{i,e+l}^{\infty}]$
$=E[Y_{i}(1)-Y_{i}(0)|T_{i}=e+l,E_{i}=e,CM_{e,e+l}].$
If we have an ordered, non-binary treatment, our target parameter is the $CACRT_{e,e+l}$ in period $T=e+l$:
$CACRT_{e,e+l}\equiv\sum_{j=1}^{J}w^{e}_{e+l,j}\cdot E[Y_{i,e+l}(j)-Y_{i,e+l}(j-1)|T_{i}=e+l,E_{i}=e,D_{i,e+l}^{e}\geq j>D_{i,e+l}^{\infty}]$
where the weight $w^{e}_{e+l,j}$ is:
$w^{e}_{e+l,j}=\frac{Pr(D_{i,e+l}^{e}\geq j>D_{i,e+l}^{\infty}|T_{i}=e+l,E_{i}=e)}{\sum_{j=1}^{J}Pr(D_{i,e+l}^{e}\geq j>D_{i,e+l}^{\infty}|T_{i}=e+l,E_{i}=e)}.$
Similar to Section 3.3, we first consider the following estimand to identify each $CLATT_{e,e+l}$:
$w^{DID,r}_{e,l}=\frac{\Delta^{l}_{Y,E=e}-\Delta^{l}_{Y,E=\infty}}{\Delta^{l}_{D,E=e}-\Delta^{l}_{D,E=\infty}},\qquad\text{for }e\in\mathcal{E}\text{ and }l\in\{0,\dots,T-e\}.$
where, for any $e\in\mathcal{S}(E_{i})$,
$\Delta^{l}_{Y,E=e}$ $=E[Y_{i}\mid E_{i}=e,T_{i}=e+l]-E[Y_{i}\mid E_{i}=e,T_{i}=e-1],$
$\Delta^{l}_{D,E=e}$ $=E[D_{i}\mid E_{i}=e,T_{i}=e+l]-E[D_{i}\mid E_{i}=e,T_{i}=e-1].$
Note that in this Wald-DID estimand, the pre-exposed period is e1e-1 and the control group is the never exposed cohort. Here, we assume that the largest cohort value in the dataset is $\bar{e}=\infty$.
When repeated cross section data are available, the staggered DID-IV designs consist of Assumptions 8, 11, 12 (in Section 3), Assumption 19, and the following parallel trends assumptions and the relevance condition.
Assumption 25 (Parallel trends assumption in the treatment based on a never exposed cohort).
$\text{For each}\hskip 5.69054pte\in\mathcal{E}\hskip 5.69054pt\text{and}\hskip 5.69054ptt\in\{2,\dots,T\}\hskip 5.69054pt\text{such that}\hskip 5.69054ptt\geq e,$
$E[D_{i}^{\infty}|E_{i}=e,T_{i}=t]-E[D_{i}^{\infty}|E_{i}=e,T_{i}=t-1]$
$E[D_{i}^{\infty}|E_{i}=\infty,T_{i}=t]-E[D_{i}^{\infty}|E_{i}=\infty,T_{i}=t-1].$
Assumption 26 (Parallel trends assumption in the outcome based on a never exposed cohort).
$\text{For each}\hskip 5.69054pte\in\mathcal{E}\hskip 5.69054pt\text{and}\hskip 5.69054ptt\in\{2,\dots,T\}\hskip 5.69054pt\text{such that}\hskip 5.69054ptt\geq e,$
$E[Y_{i}(D_{i}^{\infty})|E_{i}=e,T_{i}=t]-E[Y_{i}(D_{i}^{\infty})|E_{i}=e,T_{i}=t-1]$
$E[Y_{i}(D_{i}^{\infty})|E_{i}=\infty,T_{i}=t]-E[Y_{i}(D_{i}^{\infty})|E_{i}=\infty,T_{i}=t-1].$
Assumption 27 (Relevance condition based on a never exposed cohort).
$\text{For each}\hskip 5.69054pte\in\mathcal{E}\hskip 5.69054pt\text{and}\hskip 5.69054ptl\in\{0,\dots,T-e\},$
$\Delta^{l}_{D,E=e}-\Delta^{l}_{D,E=\infty}>0.$
By similar arguments in the proof of Theorem 2, one can show that each Wald-DID estimand $w^{DID,r}_{e,l}$ is equal to the $CLATT_{e,e+l}$ ($CACRT_{e,e+l}$ under an ordered, non-binary treatment) for each $e\in\mathcal{E}$ and $l\in\{0,\dots,T-e\}$ under Assumptions 8, 11, 12 in Section 3, Assumption 19, and Assumptions 2527 (thus, we omit the proof for brevity).
When there is no never-exposed cohort, or its sample size is too small, one can instead consider the following estimand:
$w^{DID,ny,r}_{e,l}=\frac{\Delta^{l}_{Y,E=e}-\Delta^{l}_{Y,ny}}{\Delta^{l}_{D,E=e}-\Delta^{l}_{D,ny}},$
where
$\Delta^{l}_{Y,ny}$ $=E[Y_{i}\mid Z_{i}=0,E_{i,e}=0,T_{i}=e+l]-E[Y_{i}\mid Z_{i}=0,E_{i,e}=0,T_{i}=e-1],$
$\Delta^{l}_{D,ny}$ $=E[D_{i}\mid Z_{i}=0,E_{i,e}=0,T_{i}=e+l]-E[D_{i}\mid Z_{i}=0,E_{i,e}=0,T_{i}=e-1].$
for $e\in\mathcal{E}$ and $l\in\{0,\dots,T-e\}$ such that $l<\bar{e}-e$. In this Wald-DID estimand, the control group is the not-yet-exposed cohorts.
If we adopt the above estimand $w^{DID,ny,r}_{e,l}$, we can replace Assumptions 2527 with Assumptions 2830 below.
Assumption 28 (Parallel trends assumption in the treatment based on not-yet-exposed cohorts).
$\text{For each}\hskip 5.69054pte\in\mathcal{E}\hskip 5.69054pt\text{and}\hskip 5.69054pt\text{each}\hskip 5.69054ptt\in\{2,\dots,T\}\hskip 5.69054pt\text{such that}\hskip 5.69054pte\leq t<\bar{e},$
$E[D_{i}^{\infty}|E_{i}=e,T_{i}=t]-E[D_{i}^{\infty}|E_{i}=e,T_{i}=t-1]$
$E[D_{i}^{\infty}|Z_{i}=0,E_{i,e}=0,T_{i}=t]-E[D_{i}^{\infty}|Z_{i}=0,E_{i,e}=0,T_{i}=t-1].$
Assumption 29 (Parallel trends assumption in the outcome based on not-yet-exposed cohorts).
$\text{For each}\hskip 5.69054pte\in\mathcal{E}\hskip 5.69054pt\text{and}\hskip 5.69054pt\text{each}\hskip 5.69054ptt\in\{2,\dots,T\}\hskip 5.69054pt\text{such that}\hskip 5.69054pte\leq t<\bar{e},$
$E[Y_{i}(D_{i}^{\infty})|E_{i}=e,T_{i}=t]-E[Y_{i}(D_{i}^{\infty})|E_{i}=e,T_{i}=t-1]$
$E[Y_{i}(D_{i}^{\infty})|Z_{i}=0,E_{i,e}=0,T_{i}=t]-E[Y_{i}(D_{i}^{\infty})|Z_{i}=0,E_{i,e}=0,T_{i}=t-1].$
Assumption 30 (Relevance condition based on not-yet-exposed cohorts).
$\text{For each}\hskip 5.69054pte\in\mathcal{E}\hskip 5.69054pt\text{and}\hskip 5.69054ptl\in\{0,\dots,T-e\}\hskip 5.69054pt\text{such that}\hskip 5.69054ptl<\bar{e}-e,$
$\Delta^{l}_{D,E=e}-\Delta^{l}_{D,ny}>0.$
Again, by similar arguments in the proof of Theorem 3, one can show that each Wald-DID estimand $w^{DID,ny,r}_{e,l}$ is equal to the $CLATT_{e,e+l}$ ($CACRT_{e,e+l}$ under an ordered, non-binary treatment) for each $e\in\mathcal{E}$ and $l\in\{0,\dots,T-e\}$ such that $l<\bar{e}-e$ under Assumptions 8, 11, 12 in Section 3), 19, and Assumptions 2830 (thus, we omit the proof for brevity).
Appendix B Deriving the influence function in repeated cross section settings
In this appendix, we present the influence function of our STS estimator when repeated cross section data are available. [^22] The derivation here is essentially the same as in de Chaisemartin and D’Haultfœuille (2018), who present the influence function of the Wald-DID estimator in repeated cross section settings with two periods and two groups. Let $\theta_{e,l}^{r}$ define
$\theta_{e,l}^{r}\equiv\frac{\alpha_{e,e+l}-\alpha_{e,e-1}-[\beta_{U,e+l}-\beta_{U,e-1}]}{\pi_{e,e+l}-\pi_{e,e-1}-[\gamma_{U,e+l}-\gamma_{U,e-1}]},$
where
$\alpha_{e,t}=E[Y_{i}|E_{i}=e,T_{i}=t],\hskip 5.69054pt\beta_{U,t}=E[Y_{i}|E_{i}\in U,T_{i}=t],$
$\pi_{e,t}=E[D_{i}|E_{i}=e,T_{i}=t],\hskip 5.69054pt\gamma_{U,t}=E[D_{i}|E_{i}\in U,T_{i}=t]\hskip 5.69054pt(t=e+l,e-1).$
In repeated cross section settings, our STS estimator is the sample analog of $\theta_{e,l}^{r}$, which we denote $\hat{\theta}_{e,l}^{r}$. Then, if we set $U=\{\infty\}$ ($U=\{e^{\prime}\in\mathcal{S}(E_{i}):e^{\prime}>e+l\}$), one can show that $\hat{\theta}_{e,l}^{r}$ is consistent for $CLATT_{e,e+l}$ under Assumptions 8, 11, 12, 19 and Assumptions 2527 (Assumptions 2830), and have the following influence function:
$\psi_{i,e,l}^{r}=\frac{1}{\pi_{e,e+l}-\pi_{e,e-1}-[\gamma_{U,e+l}-\gamma_{U,e-1}]}$
$\times\Bigg(\frac{\mathbf{1}_{e,e+l}\cdot\left[\delta^{r}_{i,e,l}-E[\delta^{r}_{i,e,l}|E_{i}=e,T_{i}=e+l]\right]}{E[\mathbf{1}_{e,e+l}]}-\frac{\mathbf{1}_{e,e-1}\cdot\left[\delta^{r}_{i,e,l}-E[\delta^{r}_{i,e,l}|E_{i}=e,T_{i}=e-1]\right]}{E[\mathbf{1}_{e,e-1}]}$
$-\frac{\mathbf{1}_{U,e+l}\cdot\left[\delta^{r}_{i,e,l}-E[\delta^{r}_{i,e,l}|E_{i}\in U,T_{i}=e+l]\right]}{E[\mathbf{1}_{U,e+l}]}+\frac{\mathbf{1}_{U,e-1}\cdot\left[\delta^{r}_{i,e,l}-E[\delta^{r}_{i,e,l}|E_{i}\in U,T_{i}=e-1]\right]}{E[\mathbf{1}_{U,e-1}]}\Bigg),$ (14)
where $\mathbf{1}_{A}$ is the indicator function and takes one if $A$ is true, and $\delta_{i,e,l}^{r}=Y_{i}-\theta_{e,l}^{r}\cdot D_{i}$.
Appendix C Decomposing the TWFEIV estimand in Oreopoulos (2006)
In this appendix, we quantify the bias terms in the TWFEIV estimand, arising from the bad comparisons in Oreopoulos (2006). Let $e=1947$ and $U=1957$ denote England and Northern Ireland respectively, and let $N_{e,t}$ denotes the sample size for England in cohort $t$. Let $R_{e,t}\equiv\frac{1}{N_{e,t}}\sum_{i}^{N_{e,t}}R_{e(i),t}$ denote the sample mean of random variable $R_{i,t}$ in region $e$ ($e\in\{1947,1957\}$) and cohort $t$. Here, $R_{e(i),t}$ indicates that the unit $i$ belongs to cohort $e$.
First, we decompose the TWFEIV estimator $\hat{\beta}_{IV}$ as follows [^23] For the detailed calculation steps, see the proof of Lemma 7 in Miyaji (2024). In our decomposition, the reference period ($t=1$) in the proof is $t=1946$.:
$\hat{\beta}_{IV}$ $=\frac{\sum_{t}N_{e,t}\hat{Z}_{e,t}\left[Y_{e,t}-Y_{e,1946}-(Y_{U,t}-Y_{U,1946})\right]}{\sum_{t}N_{e,t}\hat{Z}_{e,t}\left[D_{e,t}-D_{e,1946}-(D_{U,t}-D_{U,1946})\right]}$
$=\frac{\sum_{t}N_{e,t}\hat{Z}_{e,t}\left[D_{e,t}-D_{e,1946}-(D_{U,t}-D_{U,1946})\right]\cdot\widehat{WDID}_{e,t}}{\sum_{t}N_{e,t}\hat{Z}_{e,t}\left[D_{e,t}-D_{e,1946}-(D_{U,t}-D_{U,1946})\right]},$
where we define:
$\widehat{WDID}_{e,t}\equiv\frac{\left[Y_{e,t}-Y_{e,1946}-(Y_{U,t}-Y_{U,1946})\right]}{\left[D_{e,t}-D_{e,1946}-(D_{U,t}-D_{U,1946})\right]},$
and $\hat{Z}_{e,t}$ is the residuals from regressing $Z_{i,t}$ on the cohort and Northern Ireland fixed effects. [^24] We can write $\hat{Z}_{i,t}=\hat{Z}_{e,t}$ because $Z_{i,t}$ only varies across cohorts and regions.
Here, to clarify the interpretation of the TWFEIV estimand, we assume the parallel trends assumption in the treatment and the outcome over the entire sample period from $t=1936$ to $t=1965$. Then, by using the similar arguments in the proof of Theorem 5 for multiple time periods repeatedly, we obtain the following decomposition result for the TWFEIV estimand in Oreopoulos (2006):
$\beta_{IV}=$ $\sum\limits_{U-1\geq t\geq e}w_{e,t}^{1}\cdot CACRT_{e,t}+\sum\limits_{t\geq U}w_{e,t}^{2}\cdot\Delta_{e,t},$ (15)
where we define:
$CAET^{1}_{e,t}=E[D_{i}-D_{i}^{\infty}|T_{i}=t,E_{i}=e],$
$\Delta_{e,t}=\displaystyle\frac{CAET^{1}_{e,t}\cdot CACRT_{e,t}-CAET^{1}_{U,t}\cdot CACRT_{U,t}}{CAET^{1}_{e,t}-CAET^{1}_{U,t}},$
and the weights $w^{1}_{e,t}$ and $w^{2}_{e,t}$ are:
$w_{e,t}^{1}$ (16)
$=\frac{E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot CAET^{1}_{e,t}}{\sum\limits_{U-1\geq t\geq e}E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot CAET^{1}_{e,t}+\sum\limits_{t\geq U}E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot(CAET^{1}_{e,t}-CAET^{1}_{U,t})},$ (17)
$w_{e,t}^{2}$ (18)
$=\frac{E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot(CAET^{1}_{e,t}-CAET^{1}_{U,t})}{\sum\limits_{U-1\geq t\geq e}E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot CAET^{1}_{e,t}+\sum\limits_{t\geq U}E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot(CAET^{1}_{e,t}-CAET^{1}_{U,t})},$ (19)
where $n_{e,t}$ and $E[\hat{Z}_{i,t}|E_{i}=e]$ represent the population share and residuals for England in cohort $t$, respectively.
In Oreopoulos (2006), we can only identify $CACRT_{e,t}$ between $1947$ and $1956$ because the policy change occurred in England in $1947$ and in Northern Ireland in $1957$. This implies that each $\Delta_{e,t}$ in equation (15) is the bias term arising from the bad comparisons between $1957$ and $1965$ performed by the TWFEIV regression.
Figure 3 plots the weight and the corresponding estimate for each $CACRT_{e,t}$ and $\Delta_{e,t}$ (bias term). The consistent estimators for each $w^{1}_{e,t}$, $w^{2}_{e,t}$, $CACRT_{e,t}$ and $\Delta_{e,t}$ are constructed from the sample analogue, exploiting $\widehat{WDID}_{e,t}$ and $(D_{e,t}-D_{e,1946})-(D_{U,t}-D_{U,1946})$.
Figure 3 shows that all the CACRTs are positive and positively weighted, whereas all the bias terms are positive and negatively weighted. This indicates that the TWFEIV estimand is negatively biased in the setting of Oreopoulos (2006).
Figure 3: Decomposition result for the TWFEIV estimand in Oreopoulos (2006). Notes: This figure plots the estimated weights and the corresponding estimates for each $CACRT_{e,t}$ and $\Delta_{e,t}$ (bias term). The closed squares are the weights and the corresponding estimates for each $CACRT_{e,t}$. The closed triangles are the weights and the corresponding estimates for each $\Delta_{e,t}$ (bias term).
Appendix D Comparing DID-IV with Fuzzy DID
In this section, we compare the DID-IV with the Fuzzy DID proposed in de Chaisemartin and D’Haultfœuille (2018) (henceforth, “dCDH”). Although both papers consider the same empirical setting and the same Wald-DID estimand, they differ in their target parameters and associated identifying assumptions. This section clarifies the implications of these differences for treatment adoption behavior, the interpretation of the target parameter, and the use of the Wald-DID estimand.
D.1  Similarities and differences between this paper and dCDH
In this subsection, we first clarify the similarities and differences between this paper and dCDH with respect to the empirical setting, the target parameter, the estimand, and the identifying assumptions. This subsection provides the foundation for the subsequent comparison.
Empirical setting
In this paper, we mainly consider a two-period, two-group setting in which one group becomes exposed to the instrument in the second period (the exposed group), while the other group is never exposed to the instrument in either period (the unexposed group). The key features of this setting are that (i) the instrument is assigned only in the second period (sharp assignment of the instrument), and (ii) the group indicator captures exposure to the instrument. We refer to this setup as the “DID-IV” setting.
Under this definition, dCDH also consider the DID-IV setting. In particular, the group variable $G$ in dCDH plays the role of the instrument $Z$ in this paper. This follows from the fact that $G$ is included in their treatment participation equation:
$D=\mathbf{1}\{V\geq v_{GT}\}$
(see Assumption 3 in dCDH), which corresponds to condition (ii) of the DID-IV setting. Condition (i) is also assumed in dCDH: they impose $v_{10}=v_{00}$ throughout their proofs (see page 1019 of dCDH), implying that the treatment participation thresholds at time $T=0$ are identical across groups $G=0$ and $G=1$.
Here, it is useful to clarify the distinction between the empirical setting and the design in dCDH. In dCDH, the Fuzzy DID design is described as a situation in which the treatment rate in one group increases more than that in another group (see Assumption 1 in dCDH). However, this characterization corresponds to an identifying assumption, rather than a primitive feature of the empirical setting. Independently of Assumption $1$, dCDH also define the Fuzzy DID setting as $D\neq G\times T$. This definition corresponds to the DID-IV setting once combined with the restriction $v_{10}=v_{00}$.
Based on the above discussion, in what follows, we replace $G$ with $Z$ and refer to the groups $Z=1$ and $Z=0$ as the exposed and unexposed groups, respectively.
The target parameter
In this paper, we consider the LATET as our target parameter. By contrast, dCDH focus on the switcher local average treatment effect on the treated (SLATET) defined below.
Definition.
The switcher local average treatment effect on the treated (SLATET) is
$SLATET$ $\equiv E[Y(1)-Y(0)|Z=1,T=1,D_{0}(0)<D_{1}(1)]$
$=E[Y(1)-Y(0)|Z=1,T=1,SW],$
where we use the restriction $v_{10}=v_{00}$ in dCDH, which implies $D_{0}(1)=D_{0}(0)$.
Here, to ease interpretation, we slightly modify the notation in dCDH. Specifically, we define $D_{t}(Z)=\mathbf{1}\{V\geq v_{Zt}\}$ as the treatment status at time $T=t$ for group $Z$, rather than using the notation $D(t)$ in dCDH.
This parameter measures the treatment effects for units who belong to the exposed group ($Z=1$) and switch into treatment at time $T=1$ (switchers). In subsection D.3, we clarify the differences between the SLATET and the LATET considered in this paper.
The estimand
Both papers consider the same Wald-DID estimand. However, the stance regarding the use of the Wald-DID estimand differs between this paper and dCDH. Specifically, this paper views the Wald-DID estimand as the natural estimand for identifying the LATET. By contrast, dCDH point out that the Wald-DID estimand identifies the SLATET only if the treatment effect is stable over time, and propose alternative estimands that avoid this restriction. We return to this point in subsection D.4.
The identifying assumptions
dCDH impose the following identifying assumptions for the Wald-DID estimand to identify the SLATET. For detailed discussions of each assumption, see dCDH.
Assumption 31 (Fuzzy design).
$E[D|Z=1,T=1]>E[D|Z=1,T=0],$ (20)
$E[D|Z=1,T=1]-E[D|Z=1,T=0]>E[D|Z=0,T=1]-E[D|Z=0,T=0].$ (21)
Assumption 32 (Stable percentage of treated units in an unexposed group).
$0<E[D|Z=0,T=1]=E[D|Z=0,T=0]<1.$
Assumption 33 (Treatment participation equation).
$D=\mathbf{1}\{V\geq v_{ZT}\},\hskip 2.84526pt\text{with}\hskip 5.69054ptV\mathop{\perp\!\!\!\!\perp}T|Z.$
Assumption 34 (Parallel trends assumption in the untreated outcome).
$E[Y(0)|Z=1,T=1]-E[Y(0)|Z=1,T=0]$
$=$ $E[Y(0)|Z=0,T=1]-E[Y(0)|Z=0,T=0].$
Assumption 35 (Stable treatment effect over time).
$E[Y(1)-Y(0)|Z,T=1,D_{0}(Z)=1]=E[Y(1)-Y(0)|Z,T=0,D_{0}(Z)=1].$
Again, in Assumption 35, we adopt the notation $D_{t}(Z)=\mathbf{1}\{V\geq v_{Zt}\}$, rather than $D(t)$ used in dCDH, to facilitate interpretation.
We now state the detailed connections between the identifying assumptions in dCDH and those in this paper. Because dCDH consider a repeated cross-section setting, when comparing the identifying assumptions across the two papers, we refer to Assumptions 19-24.
First, both dCDH and this paper impose condition (21) in Assumption 31, which corresponds to the standard relevance condition. Next, both dCDH and this paper impose Assumption 20. Specifically, Assumption 33 (treatment participation equation) in dCDH implies Assumption 20 in this paper (Vytlacil (2002)).
Moreover, dCDH implicitly impose Assumptions 19 and 21 when formalizing potential outcomes and treatment participation. Specifically, dCDH do not introduce instrument paths in their framework and define potential outcomes as $Y(D)$, which implies Assumption 19. In addition, as shown in Assumption 33, dCDH assume that the treatment participation equation depends only on the current instrument status. Combined with sharp assignment of the instrument ($v_{10}=v_{00}$), this implies Assumption 21.
The differences between dCDH’s assumptions and ours are as follows. First, dCDH impose the parallel trends assumption in untreated outcomes (Assumption 34), whereas we impose the parallel trends assumption in unexposed outcomes. Second, dCDH assume stable treatment rates in the unexposed group (Assumption 32), whereas we assume the parallel trends in treatment between exposed and unexposed groups. Third, dCDH impose an independence assumption between unobserved heterogeneity $V$ and time $T$ conditional on $Z$ in Assumption 33, whereas we do not impose such a restriction. Fourth, the treatment participation equation in Assumption 33 is stronger than Assumption 20 in this paper. In particular, Assumption 33 imposes the monotonicity with respect to time $T$, in addition to the monotonicity with respect to instrument $Z$ (see Lemma 1 in Subsection D.2). Finally, dCDH impose condition (20) in Assumption 31, which requires that the treatment rate in the exposed group increases between the two periods, whereas we do not impose this condition.
In the next subsection, we examine how these differences imply different restrictions on treatment adoption behavior across the two designs.
Remark 3.
When Assumption 32 is violated, dCDH consider the following assumption for the Wald-DID estimand to capture the SLATET.
Assumption 36 (Homogeneous treatment effect between exposed and unexposed groups).
$SLATET=SLATET^{\prime},$
where we define252525dCDH denote SLATET and SLATET’ as $\Delta$ and $\Delta^{\prime}$ respectively. When we define the SLATET, we use $D_{0}(0)=D_{0}(1)$, which follows from the sharp assignment of the instrument: $v_{10}=v_{00}$.:
$SLATET$ $=E[Y(1)-Y(0)|Z=1,T=1,D_{0}(0)<D_{1}(1)],$
$SLATET^{\prime}$ $=E[Y(1)-Y(0)|Z=0,T=1,D_{0}(0)\neq D_{1}(0)].$
This assumption requires that the treatment effects among the switchers should be the same between exposed and unexposed groups, which dCDH call the homogeneous treatment effect assumption. In Appendix D.3, we present the decomposition result for the SLATET, and show that we cannot straightforwardly interpret this assumption as requiring the homogeneous treatment effect between the two groups. In the following discussion, we therefore treat Assumptions 31-35 as dCDH’s main identifying assumptions.
D.2  Restrictions on treatment adoption behavior
In this subsection, we compare the restrictions on treatment adoption behavior between the two designs. Since both this paper and dCDH impose Assumption 21 (see the previous section), we adopt the notation $(D_{0}(0),D_{1}(0),D_{1}(1))$ for potential treatment choices in the following. Note that while we do not use $D_{0}(1)$ in DID-IV designs due to the sharp assignment of the instrument, we can define $D_{0}(1)=\mathbf{1}\{V\geq v_{10}\}$ in dCDH’s notation. However, since dCDH assume $v_{10}=v_{00}$, we have $D_{0}(1)=D_{0}(0)$ almost surely. We use this fact repeatedly in the following discussion.
D.2.1 Preparation
We first describe the heterogeneity in treatment adoption behavior under the DID-IV setting. This allows us to clarify which latent adoption types are ruled out by the identifying assumptions in Fuzzy DID, and how these restrictions differ from those imposed in DID-IV. To do so, we introduce the following notation. Let $G^{T}=(D_{0}(0),D_{1}(0))$ denote the group variable that represents the treatment adoption process across the two periods when the instrument is $Z=0$ in the second period. That is, $G^{T}$ captures whether an individual would start, stop, or never adopt treatment over time in the absence of the instrument. Analogous to $G^{Z}=(D_{1}(0),D_{1}(1))$ (which was introduced in Section 2.1 for the panel data case), we define $G^{T}=(0,0)\equiv NT^{T}$ as time never-takers, $G^{T}=(1,1)\equiv AT^{T}$ as time always-takers, $G^{T}=(0,1)\equiv CM^{T}$ as time compliers, and $G^{T}=(1,0)\equiv DF^{T}$ as time defiers.
Using the group variables $G^{Z}$ and $G^{T}$, we can partition units into eight types within each group. Specifically, we obtain the following tables for the exposed and unexposed groups, respectively.
Table 7: Exposed group ($z=1$)
observed counterfactual
[…]
$D_{0}(0)=1,D_{1}(1)=1$ $AT^{Z}\land AT^{T}$ $CM^{Z}\land DF^{T}$
$D_{0}(0)=1,D_{1}(1)=0$ $DF^{Z}\land AT^{T}$ $NT^{Z}\land DF^{T}$
$D_{0}(0)=0,D_{1}(1)=1$ $AT^{Z}\land CM^{T}$ $CM^{Z}\land NT^{T}$
$D_{0}(0)=0,D_{1}(1)=0$ $DF^{Z}\land CM^{T}$ $NT^{Z}\land NT^{T}$
Table 8: Unexposed group ($z=0$)
observed counterfactual
[…]
$D_{0}(0)=1,D_{1}(0)=1$ $AT^{Z}\land AT^{T}$ $DF^{Z}\land AT^{T}$
$D_{0}(0)=1,D_{1}(0)=0$ $CM^{Z}\land DF^{T}$ $NT^{Z}\land DF^{T}$
$D_{0}(0)=0,D_{1}(0)=1$ $AT^{Z}\land CM^{T}$ $DF^{Z}\land CM^{T}$
$D_{0}(0)=0,D_{1}(0)=0$ $CM^{Z}\land NT^{T}$ $NT^{Z}\land NT^{T}$
Notes: These tables represent mutually exclusive and exhaustive types under the DID-IV setting.
In the following, we examine which types are excluded by the identifying assumptions under DID-IV and Fuzzy DID, respectively.
D.2.2 DID-IV
In DID-IV designs, the only restriction on treatment adoption behavior is the monotonicity assumption (Assumption 20), which rules out either $DF^{Z}$ or $CM^{Z}$. Therefore, if we rule out $DF^{Z}$, we obtain the following table.
Table 9: Exposed group ($z=1$)
observed counterfactual
[…]
$D_{0}(0)=1,D_{1}(1)=1$ $AT^{Z}\land AT^{T}$ $CM^{Z}\land DF^{T}$
$D_{0}(0)=1,D_{1}(1)=0$ $DF^{Z}\land AT^{T}$ $NT^{Z}\land DF^{T}$
$D_{0}(0)=0,D_{1}(1)=1$ $AT^{Z}\land CM^{T}$ $CM^{Z}\land NT^{T}$
$D_{0}(0)=0,D_{1}(1)=0$ $DF^{Z}\land CM^{T}$ $NT^{Z}\land NT^{T}$
Table 10: Unexposed group ($z=0$)
observed counterfactual
[…]
$D_{0}(0)=1,D_{1}(0)=1$ $AT^{Z}\land AT^{T}$ $DF^{Z}\land AT^{T}$
$D_{0}(0)=1,D_{1}(0)=0$ $CM^{Z}\land DF^{T}$ $NT^{Z}\land DF^{T}$
$D_{0}(0)=0,D_{1}(0)=1$ $AT^{Z}\land CM^{T}$ $DF^{Z}\land CM^{T}$
$D_{0}(0)=0,D_{1}(0)=0$ $CM^{Z}\land NT^{T}$ $NT^{Z}\land NT^{T}$
Notes: These tables represent mutually exclusive and exhaustive types under DID-IV designs. The types painted in gray color are excluded by Assumption 20.
D.2.3 Fuzzy DID
Next, we consider the treatment adoption behavior under Fuzzy DID. We first prove some lemmas implied by dCDH’s identifying assumptions. Based on these lemmas, we then construct tables that summarize all latent types in the exposed and unexposed groups under Fuzzy DID.
We begin with Lemma 1 below.
Lemma 1.
Assumption 33 implies the following two monotonicity conditions:
Assumption 37 (IV monotonicity at time $T=1$).
$Pr(D_{1}(1)\geq D_{1}(0))=1\hskip 5.69054pt\text{or}\hskip 5.69054ptPr(D_{1}(1)\leq D_{1}(0))=1.$
Assumption 38 (Time monotonicity for each $z\in\{0,1\}$).
$Pr(D_{0}(z)\geq D_{1}(z))=1\hskip 5.69054pt\text{or}\hskip 5.69054ptPr(D_{0}(z)\leq D_{1}(z))=1$
for each $z\in\{0,1\}$.
Proof.
We show that Assumption 33 implies Assumption 37. Note that we should have $v_{11}\geq v_{01}$ or $v_{11}\leq v_{01}$. Without loss of generality, suppose $v_{11}\geq v_{01}$. Then, we obtain $D_{1}(1)=\mathbf{1}\{V\geq v_{11}\}\leq D_{1}(0)=\mathbf{1}\{V\geq v_{01}\}$ almost surely.
By comparing $v_{z0}$ and $v_{z1}$ for each $z\in\{0,1\}$, we can also show that Assumption 33 implies Assumption 38. ∎
Next, we prove Lemma 2 below.
Lemma 2.
Suppose Assumptions 32 and 33 hold. Then, there are no time compliers $CM^{T}$ and time defiers $DF^{T}$ in an unexposed group $Z=0$.
Proof.
Note that Assumption 33 also implies the time monotonicity given $Z=0$:
$Pr(D_{0}(0)\geq D_{1}(0)|Z=0)=1\hskip 5.69054pt\text{or}\hskip 5.69054ptPr(D_{0}(0)\leq D_{1}(0)|Z=0)=1.$
This implies that at most one of the two switching types ($CM^{T}$ or $DF^{T}$) can exist in an unexposed group $Z=0$. Without loss of generality, suppose that $DF^{T}$ is ruled out in the unexposed group.
Then, Assumption 32 implies:
$E[D|T=1,Z=0]=E[D|T=0,Z=0]$
$\iff$ $Pr(D_{1}(0)>D_{0}(0)|Z=0)=0$
$\iff$ $Pr(CM^{T}|Z=0)=0$ (22)
The first equivalence follows from $V\mathop{\perp\!\!\!\!\perp}T|Z$ in Assumption 33. Equation (22) shows that there are no time compliers $CM^{T}$ in an unexposed group $Z=0$. ∎
Next, we prove Lemma 3 below. This lemma shows that in an exposed group $Z=1$, there are no units switching from treatment to non-treatment between period 0 and period $1$. This lemma also implies that the condition (20) in Assumption 31 and Assumption 33 ensures the existence of the switchers in an exposed group (that is, the SLATET is well defined).
Lemma 3.
The condition (20) in Assumption 31 and Assumption 33 imply $Pr(D_{0}(0)\leq D_{1}(1)|Z=1)$=1 and $Pr(D_{0}(0)<D_{1}(1)|Z=1)>0$.
Proof.
Note that Assumption 33 also implies the time monotonicity given $Z=1$:
$Pr(D_{0}(1)\geq D_{1}(1)|Z=1)=1\hskip 5.69054pt\text{or}\hskip 5.69054ptPr(D_{0}(1)\leq D_{1}(1)|Z=1)=1.$
In what follows, we replace $D_{0}(1)$ with $D_{0}(0)$.
From the condition (20) in Assumption 31, we have
$E[D|Z=1,T=1]>E[D|Z=1,T=0]$
$\iff$ $E[D_{1}(1)-D_{0}(0)|Z=1]>0$
$\iff$ $Pr(D_{1}(1)>D_{0}(0)|Z=1)-Pr(D_{0}(0)>D_{1}(1)|Z=1)>0,$ (23)
where the first equivalence follows from $V\mathop{\perp\!\!\!\!\perp}T|Z$ in Assumption 33 and $D_{0}(1)=D_{0}(0)$. Suppose that $Pr(D_{0}(0)\geq D_{1}(1)|Z=1)=1$ holds. Then, we have $Pr(D_{1}(1)>D_{0}(0)|Z=1)=0$. This contradicts with the condition (23). Therefore, we should have $Pr(D_{0}(0)\leq D_{1}(1)|Z=1)=1$. From the condition (23), we also have $Pr(D_{0}(0)<D_{1}(1)|Z=1)>0$. ∎
Note that this lemma also implies that we cannot impose Assumptions 37 and 38 that rule out both $CM^{Z}$ and $CM^{T}$, which contradicts with $Pr(D_{0}(0)<D_{1}(1)|Z=1)>0$.
Finally, we prove Lemma 4 below.
Lemma 4.
The condition (20) in Assumption 31 and Assumption 33 imply $Pr(D_{0}(0)\leq D_{1}(1)|Z=0)$=1. Equivalently, the type $(DF^{Z}\land AT^{T})$ is ruled out in an unexposed group.
Proof.
From Lemma 3, we have $Pr(D_{0}(0)\leq D_{1}(1)|Z=1)$=1. Under the threshold representation $D_{0}(0)=\mathbf{1}\{V\geq v_{00}\}$ and $D_{1}(1)=\mathbf{1}\{V\geq v_{11}\}$, this implies $v_{11}\leq v_{00}$. Hence, $D_{0}(0)\leq D_{1}(1)$ holds almost surely, and therefore $Pr(D_{0}(0)\leq D_{1}(1)\mid Z=0)=1$ as well. ∎
We now clarify how dCDH’s identifying assumptions restrict the treatment adoption behavior across units in exposed and unexposed groups, respectively. Tables 11-12 show the heterogeneity in treatment adoption behavior under Fuzzy DID. Here, $DF^{Z}$ and $DF^{T}$ are excluded by Assumptions 37 and 38 (under $z=0$), respectively. Since these restrictions depend on the direction of the two monotonicity conditions, we paint these types by blue. $DF^{Z}\land AT^{T}$ and $NT^{Z}\land DF^{T}$ in an exposed group are excluded by Lemma 3 and $DF^{T}$, $CM^{T}$ and $DF^{Z}\land AT^{T}$ in an unexposed group are excluded by Lemma 2 and Lemma 4. Since these restrictions do not depend on the direction of the two monotonicity conditions, we paint these types by gray.
These tables show that four types are excluded in an exposed group and five types are excluded in an unexposed group under Fuzzy DID. Note that if we exclude $DF^{Z}$ and $CM^{T}$ by Assumption 37 and Assumption 38 (under $z=0$), the same number of types are excluded in exposed and unexposed groups, respectively. If we exclude $CM^{Z}$ and $DF^{T}$ by Assumptions 37 and 38 (under $z=0$), four types are excluded in an exposed group and six types are excluded in an unexposed group.
Table 11: Exposed group ($z=1$)
observed counterfactual
[…]
$D_{0}(0)=1,D_{1}(1)=1$ $AT^{Z}\land AT^{T}$ $CM^{Z}\land DF^{T}$
$D_{0}(0)=1,D_{1}(1)=0$ $DF^{Z}\land AT^{T}$ $NT^{Z}\land DF^{T}$
$D_{0}(0)=0,D_{1}(1)=1$ $AT^{Z}\land CM^{T}$ $CM^{Z}\land NT^{T}$
$D_{0}(0)=0,D_{1}(1)=0$ $DF^{Z}\land CM^{T}$ $NT^{Z}\land NT^{T}$
Table 12: Unexposed group ($z=0$)
observed counterfactual
[…]
$D_{0}(0)=1,D_{1}(0)=1$ $AT^{Z}\land AT^{T}$ $DF^{Z}\land AT^{T}$
$D_{0}(0)=1,D_{1}(0)=0$ $CM^{Z}\land DF^{T}$ $NT^{Z}\land DF^{T}$
$D_{0}(0)=0,D_{1}(0)=1$ $AT^{Z}\land CM^{T}$ $DF^{Z}\land CM^{T}$
$D_{0}(0)=0,D_{1}(0)=0$ $CM^{Z}\land NT^{T}$ $NT^{Z}\land NT^{T}$
Notes: These tables represent mutually exclusive and exhaustive types under Fuzzy DID designs. The types painted in gray color are excluded by Lemmas 2 and 3. The types painted in blue color are excluded by the two monotonicity assumptions 37 and 38 (under $z=0$).
D.2.4 Comparison
By comparing Tables 9-10 with Tables 11-12, we obtain two implications. First, the restrictions imposed under Fuzzy DID are stronger than those under DID-IV. Second, while the restrictions under DID-IV are symmetric between exposed and unexposed groups, they are asymmetric under Fuzzy DID.
Note that these differences are driven by differences in the identifying assumptions between the two designs. Assumption 33, which is stronger than Assumption 20, imposes time monotonicity and thus rules out one of the two time-switching types ($CM^{T}$ or $DF^{T}$). The condition (20) in Assumption 31, which DID-IV does not impose, rules out the type $NT^{Z}\land DF^{T}$ in the exposed group and the type $DF^{Z}\land AT^{T}$ in the unexposed group. Finally, Assumption 32, combined with Assumption 20, eliminates both $DF^{T}$ and $CM^{T}$ in the unexposed group under Fuzzy DID.
D.3  The difference in target parameter
In this subsection, building on Tables 11-12, we next clarify the difference between dCDH’s target parameter (SLATET) and ours (LATET). Specifically, we decompose the SLATET under the identifying assumptions in Fuzzy DID and discuss its implications.
We prove Theorem 6 below.
Theorem 6.
Suppose Assumptions 31-35 hold. Then, we can decompose the SLATET as follows, which depends on the direction of the two monotonicity assumptions 37 and 38 (under $z=0$).
  • (i)
    Under the monotonicity assumptions that exclude $DF^{Z}$ and $DF^{T}$, we have:
    $SLATET\equiv E[Y(1)-Y(0)|Z=1,T=1,D_{0}(0)<D_{1}(1)]$
    $=E[Y(1)-Y(0)|Z=1,T=1,AT^{Z}\land CM^{T}]Pr(AT^{Z}\land CM^{T}|Z=1,SW)$
    $+E[Y(1)-Y(0)|Z=1,T=1,CM^{Z}\land NT^{T}]Pr(CM^{Z}\land NT^{T}|Z=1,SW).$ (24)
  • (ii)
    Under the monotonicity assumptions that exclude $DF^{Z}$ and $CM^{T}$, we have:
    $SLATET$ $\equiv E[Y(1)-Y(0)|Z=1,T=1,D_{0}(0)<D_{1}(1)]$
    $=E[Y(1)-Y(0)|Z=1,T=1,CM^{Z}\land NT^{T}].$
  • (iii)
    Under the monotonicity assumptions that exclude $CM^{Z}$ and $DF^{T}$, we have:
    $SLATET$ $\equiv E[Y(1)-Y(0)|Z=1,T=1,D_{0}(0)<D_{1}(1)]$
    $=E[Y(1)-Y(0)|Z=1,T=1,AT^{Z}\land CM^{T}].$
Proof.
Regarding the case (i), the proof directly follows from Table 11. We can also construct the table in an exposed group for the cases (ii) and (iii), which ensures the above interpretation of the SLATET. ∎
Theorem 6 has several important implications. First, the interpretation of the SLATET depends on the direction of the two monotonicity conditions 37 and 38 (under $z=0$). This implies that, if empirical researchers wish to understand which latent type’s causal effect is identified by the SLATET, they need to specify the direction of these monotonicity conditions ex ante.
Second, the SLATET may fail to be a policy-relevant treatment parameter (Heckman and Vytlacil (2001)), even when the instrument represents the policy change of interest to the researcher. Specifically, under case (i), the SLATET is contaminated by the treatment effects among the type $AT^{Z}\land CM^{T}$, who are affected by time but not affected by the instrument. Under case (iii), the SLATET captures only the treatment effects among the type $AT^{Z}\land CM^{T}$.
Third, the SLATET may be defined on a strictly smaller population than the LATET. In particular, under case (ii), the SLATET identifies the causal effect among the type $CM^{Z}\land NT^{T}$, which is a strict subpopulation of $CM^{Z}$. This implies that even when empirical researchers choose the direction of the two monotonicity conditions so that the SLATET is policy relevant, it will measure the treatment effects for a narrower population than the LATET.
Remark 4.
Using the decomposition result for the SLATET, we can also clarify the interpretation of Assumption 36, which dCDH introduce as an alternative to Assumption 32.
First, we describe the heterogeneity in treatment adoption behavior in the unexposed group without Assumption 32. If we do not impose Assumption 32, we obtain Table 13 for the unexposed group. Here, we consider the case in which Assumptions 37 and 38 exclude $DF^{Z}$ and $DF^{T}$ in the unexposed group, respectively.
Table 13: Unexposed group ($z=0$)
observed counterfactual
[…]
$D_{0}(0)=1,D_{1}(0)=1$ $AT^{Z}\land AT^{T}$ $DF^{Z}\land AT^{T}$
$D_{0}(0)=1,D_{1}(0)=0$ $CM^{Z}\land DF^{T}$ $NT^{Z}\land DF^{T}$
$D_{0}(0)=0,D_{1}(0)=1$ $AT^{Z}\land CM^{T}$ $DF^{Z}\land CM^{T}$
$D_{0}(0)=0,D_{1}(0)=0$ $CM^{Z}\land NT^{T}$ $NT^{Z}\land NT^{T}$
Notes: This table represents mutually exclusive and exhaustive types in an unexposed group under Fuzzy DID without Assumption 32. The types painted in gray color are excluded by Lemma 4. The types painted in blue color are excluded by the two monotonicity assumptions 37 and 38 (under $z=0$).
Then, from Table 13, we can interpret the SLATET’ (see Assumption 36) as the treatment effects among the type $AT^{Z}\land CM^{T}$ in an unexposed group:
$SLATET^{\prime}$ $\equiv E[Y(1)-Y(0)|Z=0,T=1,D_{0}(0)\neq D_{1}(0)]$
$=E[Y(1)-Y(0)|Z=0,T=1,AT^{Z}\land CM^{T}].$ (25)
Combining (25) with (24), we see that Assumption 36 ($SLATET=SLATET^{\prime}$) does not require homogeneous treatment effects for the same latent type across exposed and unexposed groups. Rather, it equates a weighted average of treatment effects among types $AT^{Z}\land CM^{T}$ and $CM^{Z}\land NT^{T}$ in the exposed group with the treatment effect for type $AT^{Z}\land CM^{T}$ in the unexposed group.
D.4  The use of the Wald-DID estimand
In this section, we explain why the role of the Wald–DID estimand differs between Fuzzy DID and DID-IV designs. In DID-IV, we adopt the Wald–DID estimand as a natural estimand for identifying the LATET. By contrast, dCDH point out that, under Fuzzy DID, the Wald-DID estimand requires the stable treatment effect assumption (Assumption 35) to identify the SLATET. In the following discussion, we show that this difference arises because Fuzzy DID and DID-IV rely on different types of the parallel trends assumption.
D.4.1 Why the Wald-DID estimand requires Assumption 35 under Fuzzy DID
We begin by explaining why the Wald-DID estimand requires Assumption 35 under Fuzzy DID. We show that this requirement does not stem from the Wald-DID estimand itself, but rather from imposing the parallel trends assumption on untreated outcomes.
To make this point precise, we first clarify Assumption 35 using latent treatment adoption types. This assumption is in fact conditional on $D_{0}(Z)=1$ in addition to $Z$ and $T$, but it is not immediately clear from dCDH why such a strong restriction is imposed only on this subpopulation.
From Tables 11-12, we can reinterpret Assumption 35 as follows [^26] Under the identifying assumptions in Fuzzy DID, the type $DF^{Z}\land AT^{T}$ is eliminated in both groups. Therefore, we write the type $AT^{Z}\land AT^{T}$ as $AT^{T}$.:
$E[Y(1)-Y(0)|Z=0,T=1,D_{0}(0)=1]$ $=E[Y(1)-Y(0)|Z=0,T=0,D_{0}(0)=1]$
$\iff E[Y(1)-Y(0)|Z=0,T=1,AT^{T}]$ $=E[Y(1)-Y(0)|Z=0,T=0,AT^{T}],$ (26)
$E[Y(1)-Y(0)|Z=1,T=1,D_{0}(1)=1]$ $=E[Y(1)-Y(0)|Z=1,T=0,D_{0}(1)=1]$
$\iff E[Y(1)-Y(0)|Z=1,T=1,AT^{T}]$ $=E[Y(1)-Y(0)|Z=1,T=0,AT^{T}].$ (27)
Here, we use $D_{0}(0)=D_{0}(1)$ from the sharp assignment of the instrument $v_{10}=v_{00}$. Equations (26)-(27) show that Assumption 35 actually requires the stable treatment effect only for the time always-takers $AT^{T}$ within each group.
We now show that Assumption 35 is not required by the Wald-DID estimand per se, but arises because dCDH impose the parallel trends assumption on untreated outcomes. The key point is that imposing the parallel trends assumption in untreated outcomes under the DID-IV setting forces us to impute the counterfactual untreated potential outcome for the time always-takers $AT^{T}$, which in turn requires imposing Assumption 35.
To see this formally, we focus on the numerator of the Wald-DID estimand. Under Fuzzy DID, we can decompose the conditional expectations E[Y|Z=0,T=1]E[Y|Z=0,T=0]E[Y|Z=0,T=1]-E[Y|Z=0,T=0] and E[Y|Z=1,T=1]E[Y|Z=1,T=0]E[Y|Z=1,T=1]-E[Y|Z=1,T=0] as follows.
$E[Y|Z=0,T=1]-E[Y|Z=0,T=0]$
$=$ $(E[Y(1)-Y(0)|Z=0,T=1,AT^{T}]-E[Y(1)-Y(0)|Z=0,T=0,AT^{T}])$
$\times$ $Pr(AT^{T}|Z=0)+\{E[Y(0)|Z=0,T=1]-E[Y(0)|Z=0,T=0]\}.$
$E[Y|Z=1,T=1]-E[Y|Z=1,T=0]$
$=$ $(E[Y(1)-Y(0)|Z=1,T=1,AT^{T}]-E[Y(1)-Y(0)|Z=1,T=0,AT^{T}])$
$\times$ $Pr(AT^{T}|Z=1)+\{E[Y(0)|Z=1,T=1]-E[Y(0)|Z=1,T=0]\}$
$+$ $E[Y(1)-Y(0)|SW,Z=1,T=1]Pr(SW|Z=1).$
These decompositions clarify why Assumption 35 becomes necessary. Imposing parallel trends on untreated outcomes requires constructing the average time trends of the untreated potential outcome in both groups (the second term in each conditional expectation). Since these trends are not observed for time always-takers $AT^{T}$, this construction necessarily introduces the first term in the above decompositions, which captures changes in the average treatment effect among $AT^{T}$. For the Wald-DID estimand to identify the SLATET, one must therefore impose stable treatment effects for $AT^{T}$, as in Assumption 35.272727Here, we do not require the stable treatment assumption for the time compliers $CM^{T}$ and the type $CM^{Z}\land NT^{T}$ in an exposed group: by adding and subtracting the expectation of the untreated potential outcome path for these types, we have the third term E[Y(1)Y(0)|SW,Z=1,T=1]Pr(SW|Z=1)E[Y(1)-Y(0)|SW,Z=1,T=1]Pr(SW|Z=1), which is necessary for identifying the SLATET.
Importantly, this restriction is not intrinsic to the Wald-DID estimand itself, but is a consequence of imposing parallel trends on untreated potential outcomes.
D.4.2 Why the issue does not arise under DID-IV
In contrast to Fuzzy DID, the issue described above does not arise under DID-IV. The key difference is that, in DID-IV designs, the parallel trends assumption is imposed on unexposed outcomes rather than untreated potential outcomes. As a result, the Wald-DID estimand does not require imputing the counterfactual untreated potential outcomes for time always-takers, and hence does not require Assumption 35.
D.5  Which parallel trends assumption is suitable for the DID-IV setting
In Subsection D.4, we showed that differences in the parallel trends (PT) assumption lead to different stances toward the use of the Wald-DID estimand across Fuzzy DID and DID-IV designs. In this subsection, we argue that the PT assumption on unexposed outcomes is more suitable for DID-IV settings in terms of (i) its alignment with the source of identifying variation, (ii) empirical relevance, and (iii) testability.
Which PT assumption aligns with the source of identifying variation
The PT assumption on unexposed outcomes is more consistent with the source of identifying variation in DID-IV settings than the PT assumption on untreated potential outcomes. This is because the latter relies on variation in treatment, while the former exploits variation in instrument.
Note that in the canonical DID designs, we rely on the PT assumption in untreated outcomes because the units are partitioned into treatment and control groups based on the exposure to the treatment in the second period. In the DID-IV settings, however, the units are partitioned into exposed and unexposed groups based on the exposure to the instrument in the second period. Therefore, imposing the PT assumption on unexposed outcomes would be more natural than imposing the PT assumption on untreated potential outcomes in DID-IV settings.
Empirical relevance
In DID-IV settings, the parallel trends (PT) assumption on untreated potential outcomes is often implausible in practice, while the PT assumption on unexposed outcomes is more appealing. More precisely, because imposing parallel trends on untreated potential outcomes is often not attractive, empirical researchers exploit variation in exposure to an instrument and rely on parallel trends in unexposed outcomes in practice.
A canonical example is Duflo (2001), who study the returns to schooling in Indonesia. The outcome is log annual earnings and the treatment is educational attainment. In this context, imposing PT on untreated potential outcomes would require that the average time trends of log earnings be the same across individuals in the absence of schooling. Given the strong selection into education, such an assumption is difficult to justify empirically. To address this concern, Duflo (2001) exploit the staggered rollout of a school construction program as an instrument for schooling and instead rely on the identifying assumption that, in the absence of the program, the evolution of wages and education across cohorts would not have differed systematically across regions. This identifying restriction corresponds precisely to the parallel trends assumptions on treatment and on unexposed outcomes in DID-IV designs.
Testability
In DID-IV settings, the PT assumption on unexposed outcomes is indirectly testable using pre-exposed period data as we discussed in Subsection 5.3. By contrast, the PT assumption on untreated potential outcomes is not testable in general. This is because some units may already adopt the treatment before period 0, which prevents a comparison of pre-trends in untreated outcomes between exposed and unexposed groups. From this perspective, the PT assumption on unexposed outcomes is more attractive to empirical researchers.
Although dCDH propose a placebo test in their appendix, its scope for assessing the plausibility of the parallel trends assumption in untreated potential outcomes is somewhat limited. Their procedure consists of two steps. [^28] Although dCDH assume conditions (28) and (29), they note that when either condition fails, the second step can no longer be used to assess Assumption 34. We therefore describe their procedure in two steps. First, we examine whether the share of treated units remain unchanged in both groups from T=1T=-1 to $T=0$:
$E[D|Z=1,T=0]-E[D|Z=1,T=-1]=0,$ (28)
$E[D|Z=0,T=0]-E[D|Z=0,T=-1]=0.$ (29)
Second, we test the following null hypothesis:
$E[Y|Z=1,T=0]-E[Y|Z=1,T=-1]$
$=$ $E[Y|Z=0,T=0]-E[Y|Z=0,T=-1].$ (30)
This testing strategy is informative under the maintained conditions (28) and (29), but its applicability may be limited in practice. First, these first-step conditions are strong, and the procedure does not provide guidance when they are violated. Second, the second step does not directly assess the evolution of the mean untreated outcome between exposed and unexposed groups in pre-exposed periods. Indeed, dCDH interpret (30) as a necessary condition for Assumptions 34 and 35 to hold under (28) and (29). Finally, while the discussion focuses on the case in which data are available for a single pre-exposed period, it is not clear how the procedure should be extended to settings with multiple pre-exposed periods.
Notably, dCDH’s procedure corresponds to our pre-trend test in Subsection 5.3 once the first-step conditions (28)–(29) are replaced by the weaker condition:
$E[D|Z=1,T=0]-E[D|Z=1,T=-1]$ (31)
$=$ $E[D|Z=0,T=0]-E[D|Z=0,T=-1].$ (32)
Under repeated cross-section settings, conditions (30)–(31) coincide with conditions (1)–(2) in Subsection 5.3. This observation suggests that the underlying logic of dCDH’s placebo test is more aligned with the assessment of parallel trends assumptions in DID-IV designs.
D.6  Other comparisons
In this subsection, we document the relationship between DID-IV and Fuzzy DID in settings with non-binary, ordered treatments and multiple groups. We also discuss the alternative estimands proposed in dCDH that avoid Assumption 35.
D.6.1 Non-binary, ordered treatment
Again, we first document the similarities and differences between this paper and dCDH with respect to the empirical setting, the target parameter and the identifying assumptions.
Empirical setting
In the case of non-binary, ordered treatment, the condition (ii) of the DID-IV setting is satisfied in dCDH because the group variable $G$ is included in their ordered treatment equation (see Assumption 3’ in dCDH). By contrast, to the best of our knowledge, the condition (i) of the DID-IV setting is not imposed in dCDH: the threshold $v^{d}_{GT}$ in their ordered treatment equation is allowed to depend on $G$ in period $T=0$ for all $d\in\{0,\dots,\bar{d}\}$. This implies that the DID-IV setting differs from the setting considered in dCDH.
Nevertheless, based on their empirical application, it is reasonable to infer that dCDH have the DID-IV setting in mind. In particular, dCDH revisit Duflo (2001), in which the outcome $Y$ is log annual earnings, the treatment $D$ is education attainment (an ordered treatment), and the group variable $G\in\{0,1\}$ indicates the regions that are highly exposed to the policy shock in the second period. Given this, the absence of condition (i) in dCDH can be interpreted as reflecting an additional restriction on treatment adoption behavior in period $T=0$.
From the above discussion, in what follows, we replace $G$ with $Z$.
The target parameter
In this paper, we define the ACRT as our target parameter. By contrast, dCDH consider the following target parameter, which we call the switcher average causal response on the treated (SACRT).
$SACRT\equiv\sum_{j=1}^{J}w_{j}\cdot E(Y(j)-Y(j-1)|Z=1,T=1,D_{0}(1)<j<D_{1}(1)),$
where $w_{j}$ is
$w_{j}=\frac{P(D\geq j|Z=1,T=1)-P(D\geq j|Z=1,T=0)}{E[D|Z=1,T=1]-E[D|Z=1,T=0]}.$
Here, to ease the interpretation, we slightly modify the notation in dCDH. Specifically, we define $D_{t}(Z)=\sum_{j=1}^{J}\mathbf{1}\{V\geq v_{Zt}\}$ as the treatment status at time $T=t$ for group $Z$, rather than using the notation $D(t)$ in dCDH (see Assumption 39 below).
The estimand
Both papers consider the same Wald-DID estimand. This paper regards the Wald-DID estimand as the natural estimand for identifying the ACRT. By contrast, dCDH point out that this estimand identifies the SACRT only if the stable treatment effect assumption (see Assumption 40 below) is satisfied.
The identifying assumptions
In this paper, we consider the same identifying assumptions under non-binary, ordered treatment settings. By contrast, while dCDH maintain Assumption 31 and Assumption 34, they replace Assumption 33 with the following one.
Assumption 39 (Ordered treatment equation).
$D=\sum_{j=1}^{J}\mathbf{1}\{V\geq v^{j}_{ZT}\},\hskip 5.69054pt\text{with}\hskip 5.69054pt-\infty=v^{0}_{ZT}<v^{1}_{ZT}\dots<v^{J+1}_{ZT}=\infty\hskip 5.69054pt\text{and}\hskip 5.69054ptV\mathop{\perp\!\!\!\!\perp}T|Z.$
In addition, they replace Assumption 32 with the following two conditions (33)-(34).
$D_{11}\succeq D_{10},$ (33)
$D_{01}\sim D_{00}.$ (34)
Here, $D_{ZT}$ denotes the treatment conditional on $Z$ and $T$, $\succeq$ denotes stochastic dominance between two random variables, and $\sim$ denotes equality in distribution. Under non-binary, ordered treatment settings, Assumption 35 is formulated as follows:
Assumption 40 (Stable treatment effect over time with non-binary, ordered treatment).
For all $d\in\{0,1,\dots,J\}$,
$E[Y(d)-Y(0)|Z,T=1,D_{0}(Z)=d]=E[Y(d)-Y(0)|Z,T=0,D_{0}(Z)=d].$
Next, we compare the restrictions on treatment adoption behavior between this paper and dCDH. We also discuss the difference in target parameter. Finally, we discuss the different stances regarding the use of the Wald-DID estimand.
Restrictions on treatment adoption behavior
In DID-IV, we restrict the treatment adoption behavior through the monotonicity with respect to the instrument in the second period (Assumption 20). [^29] Note that dCDH do not impose this restriction, as Assumption 39 does not imply Assumption 20 in general. However, Assumption 39 implies the following restrictions: $v^{j}_{1t}\geq v^{j}_{0t}\hskip 5.69054pt\text{or}\hskip 5.69054ptv^{j}_{1t}\leq v^{j}_{0t}.$ for all $j\in\{1,\dots,J\}$ and all $t\in\{0,1\}$.
In Fuzzy DID, dCDH impose the restrictions on treatment adoption behavior through the conditions (33)-(34). Specifically, as dCDH point out in their proof (see page 1025 in dCDH), these conditions combined with Assumption 39 imply:
$v^{j}_{01}$ $=v^{j}_{00},$ (35)
$v^{j}_{11}$ $\leq v^{j}_{10}$ (36)
for all $j\in\{1,\dots,J\}$. Here, the condition (35) requires that the treatment rate should be stable over time for all units in the absence of exposure to the instrument. The condition (36) requires that the treatment should be weakly increasing over time for all units under exposure to the instrument.
The difference in target parameter
In general, the SACRT considered in dCDH differs from the ACRT defined in this paper. As in the case of the SLATET, the SACRT may not be policy relevant. This is because the type $SW^{{}^{\prime}}=\{Z=1,T=1,D_{0}(1)<j<D_{1}(1)\}$ is induced to increase the treatment by time, and not equal to the type $\{Z=1,T=1,D_{1}(0)<j<D_{1}(1)\}$ in general.
Note that our target parameter, the ACRT, is the conditional version of the ACR defined in Angrist and Imbens (1995). As a result, the ACRT is policy-relevant parameter by construction.
The use of the Wald-DID estimand
dCDH point out that the stable treatment effect assumption (Assumption 40) is also required for the Wald–DID estimand to identify the SACRT. As in the binary treatment case, however, this requirement arises from imposing the parallel trends assumption on untreated outcomes.
To see this formally, we first rewrite Assumption 40 as follows:
$E[Y(d)\mid Z,T=1,D_{0}(Z)=d]-E[Y(d)\mid Z,T=0,D_{0}(Z)=d]$
$=$ $E[Y(0)\mid Z,T=1,D_{0}(Z)=d]-E[Y(0)\mid Z,T=0,D_{0}(Z)=d].$
From the above expression, this assumption can be interpreted as requiring that, for units with $D_{0}(Z)=d$, the average time trend of the potential outcome $Y(d)$ coincides with that of the untreated potential outcome $Y(0)$, conditional on $Z$.
Recall that under Fuzzy DID with ordered treatments, Assumptions (33)–(34) imply that treatment adoption is stable over time in the unexposed group and weakly increasing over time in the exposed group. As a result, units with $D_{0}(0)=d$ in the unexposed group necessarily remain at treatment level $d$ under $T=1$, while in the exposed group such units can be partitioned into those whose treatment remains at $d$ and those whose treatment increases from $d$ to $d^{\prime}$.
For units whose treatment remains constant over time, the left-hand side of Assumption 40 is observed. The assumption therefore uses this observed quantity to impute the unobserved average time trends of untreated potential outcomes for these units. For units in the exposed group whose treatment increases from $d$ to $d^{\prime}$, the same assumption additionally allows one to isolate the causal effect associated with the change from $d$ to $d^{\prime}$, which constitutes a component of the SACRT.
In short, Assumption 40 is required not because of the Wald–DID estimand itself, but because imposing the parallel trends assumption on untreated outcomes forces one to impute the unobserved evolution of untreated potential outcomes for units with $D_{0}(Z)=d$.
D.6.2 Multiple groups
In multiple group settings, the DID-IV design considered in this paper is not directly compatible with the Fuzzy DID framework. There are two reasons. First, the group variable $G$ takes a finite number of ordered values: $G\in\{0,1,\dots,\bar{g}\}$. Second, dCDH construct the “super groups” (see dCDH) from the group structure in the data, $G\in\{0,1,\dots,\bar{g}\}$, and use the Wald-DID estimand to identify the following target parameter:
$\Delta^{\ast}=E[Y(1)-Y(0)|\cup_{g=0}^{\bar{g}}S_{g},T=1],$
where $S_{g}=\{D_{0}(g)\neq D_{1}(g),G=g\}$ denotes the set of switchers in group G=gG=g. [^30] Following $D_{t}(Z)$, we define $D_{t}(G)=\mathbf{1}\{V\geq v_{Gt}\}$ to be the treatment status under time $T=t$ in group $G$.
However, even in multiple group settings, dCDH continue to impose Assumption 33. This implies that their framework still maintains two restrictions on treatment adoption behavior: monotonicity with respect to time $T$ and monotonicity with respect to the group variable $G$, in the sense of Imbens and Angrist (1994).
Remark 5.
While dCDH and its supplemental appendix study multiple groups and ordered treatments separately, to the best of our knowledge, we are not aware of the identification results for settings with multiple groups under ordered treatments in either the main text or the supplement.
D.6.3 Alternative estimand
As an alternative to the Wald-DID estimand, dCDH propose the time-corrected Wald (TC-Wald) estimand and the changes-in-changes Wald (CIC-Wald) estimand, which do not rely on Assumptions 34 and 35. Under both estimands, however, dCDH continue to impose Assumptions 31-33 to identify the SLATET. This implies that the same discussions in Subsections D.2-D.3 can be applied to both estimands regarding the restrictions on treatment adoption behavior and the target parameter, the SLATET (or SACRT). These estimands do not require Assumption 35 because Assumption 34 is not imposed for the time-always takers $AT^{T}$.
D.7  Summary and guidance
Throughout this section, we compare the DID-IV with Fuzzy DID considered in dCDH. To sum up, we obtain the following implications. First, the restrictions on treatment adoption behavior under Fuzzy DID are generally stronger than those under DID-IV. Second, the target parameter in Fuzzy DID, the SLATET, is not policy relevant in general, and even when it is policy relevant, it is defined on a strictly smaller population than the LATET. Finally, the Wald-DID estimand requires the stable treatment effect assumption under Fuzzy DID because dCDH imposes the parallel trends assumption on untreated outcomes.
We conclude this section by summarizing practical considerations that may guide empirical researchers in choosing between DID-IV and Fuzzy DID designs under the DID-IV setting.
Policy-relevant parameter
If empirical researchers are interested in identifying a policy-relevant parameter, DID-IV designs are preferable. Under Fuzzy DID designs, the target parameter may not be policy relevant in general. If researchers instead focus on the SLATET (or the SACRT), it is recommended that they clarify why the causal effect for units whose treatment status is affected by time but not by the instrument is of substantive interest in their empirical context.
Restrictions on treatment adoption behavior
If empirical researchers are interested in identifying the SLATET (SACRT), they should carefully assess the plausibility of the restrictions on treatment adoption behavior imposed under Fuzzy DID designs. If these restrictions appear implausible, researchers may instead target the LATET (ACRT) and adopt DID-IV designs, under which the only restriction on treatment adoption behavior is monotonicity with respect to the instrument.
Testability of the parallel trends assumption
If empirical researchers wish to assess the plausibility of the parallel trends assumption, DID-IV designs are preferable. Under Fuzzy DID designs, Assumption 34 (and its counterparts for the alternative estimands) is typically difficult to assess using pre-exposure period data (see Subsection 1.1 of the Supplemental Appendix in dCDH).